### v1

# Benchmark on Drug Target Interaction Modeling from a Structure Perspective

Abstract
The prediction modeling of drug-target interactions is crucial to drug discovery and design, which has seen rapid advancements owing to deep learning technologies. Recently developed methods, such as those based on graph neural networks (GNNs) and Transformers, demonstrate exceptional performance across various datasets by effectively extracting structural information. However, the benchmarking of these novel methods often varies significantly in terms of hyperparameter settings and datasets, which limits algorithmic progress. In view of these, we conduct a comprehensive survey and benchmark for drug-target interaction modeling from a structure perspective, via integrating tens of explicit (i.e., GNN-based) and implicit (i.e., Transformer-based) structure learning algorithms. To this end, we first unify the hyperparameter setting within each class of structure learning methods. Moreover, we conduct a macroscopical comparison between these two classes of encoding strategies as well as the different featurization techniques that inform molecules’ chemical and physical properties. We then carry out the microscopical comparison between all the integrated models across the six datasets, via comprehensively benchmarking their effectiveness and efficiency. Remarkably, the summarized insights from the benchmark studies lead to the design of model combos. We demonstrate that our combos can achieve new state-of-the-art performance on various datasets associated with cost-effective memory and computation. Our code is available at https://github.com/justinwjl/GTB-DTI/tree/main.
1 Introduction
The prediction modeling of drug-target interactions (DTI) has emerged as an irreplaceable task for efficacious therapeutic interventions. The binding affinity between a drug molecule and its target protein plays a significant role in the design and repurpose of drugs, where a high affinity typically indicates the desired therapeutics, target specificity, long residence, and drug resistance delay [1, 2, 3]. The precise modeling of DTI can expedite the drug discovery process and circumvent the associated cost [4, 5]. Deep learning based frameworks have recently revolutionized this field, enabling more accurate and efficient predictions compared with laboratory experimental methods [6, 7, 8].
Within the deep learning frameworks [9, 10], drugs are commonly represented using the Simplified Molecular Input Line Entry System (SMILES) [11], and proteins are represented as sequences of amino acids. These representations are processed by separate convolutional neural networks (CNNs) [12, 13] and subsequently integrated and processed using a multi-layer perceptron (MLP) for DTI prediction. It is notorious that the reliance on sequence-based representations can result in the loss of structural information, which can potentially compromise the DTI predictive capability. From the drug perspective, molecular structure modeling helps identify the specific binding sites [14], contribute to predicting pharmacokinetic properties [15], and allow conformational flexibility [16].
To address this problem, a number of drug algorithms have been proposed to promote DTI prediction, which can be categorized into explicit and implicit structure learning. First, graph neural networks (GNNs) [17, 18] have been widely adopted to learn the molecular structures, owing to their ability to directly operate on graph-based representations of molecules. By explicitly propagating information through the graph, GNNs can learn node and edge features and thereby capture the structural and functional relationships between atoms and bonds. Second, Transformers, originally focused on natural language processing [19], have also shown promise in biomedical applications [20, 21]. They rely on self-attention mechanisms to implicitly weight the correlations between different parts of the input SMILES, allowing them to capture long-range dependencies and contextual information.
While these techniques contribute to the learning of drug structures, there is still a key knob under-explored: we lack a systematic study to benchmark their effectiveness and efficiency. Without such a standardized benchmark, it is unachievable to offer fair comparisons and subsequently summarize designing philosophy necessary to inform DTI. There have been several surveys and benchmarks on computational methods for DTI prediction [9, 8, 22, 23], which leave out the recent developments of structure learning algorithms and unavoidably fail to focus on drug structure benchmarking. Moreover, although massive efforts [24, 25, 18] have been made to explore the effectiveness of modeling structural information, they predominantly use their proprietary training hyperparameters, datasets, and evaluation metrics. Due to the various settings, one cannot reach convincing answers whether a configuration of structure encoders and/or featurization methods generally performs well. The complex of DTI classification and regression tasks and datasets complicates the benchmark comparison.
In this study, we introduce GTB-DTI, a comprehensive benchmark customized for GNN and Transformer-based methodologies for DTI prediction. i) We thoroughly examine the implementation details for each category of drug structure learning methods and integrate three widely-used datasets for classification and regression tasks, respectively. Then, we harmonize the sensitive hyperparameters across different methods using a greedy search to identify an optimal sweet spot configuration. The unified setting lays the foundation for a fair and reproducible benchmark. ii) To gain macroscopical insights into the structure encoders and featurization methods, we fix the drug encoder to be either GNN or Transformer-based approaches and benchmark these two strategies in the various settings. We also integrate tens of drug features given their importance to inform molecules’ chemistry and physical properties and evaluate them on the representative datasets. iii) To gain macroscopical insights into nuance between 31 concerned models, we conduct the benchmark studies of their effectiveness on the six datasets with the unified setting. Moreover, we assess the efficiency of each method by measuring peak GPU memory usage, running time, and convergences. iv) The comprehensive study finally provides a number of surprising observations: ❶ The CNN-encoder accompanied with integer features has the close protein embedding performance compared to the Transformer or larger language models, but they are more efficient. ❷ The explicit and implicit structure encoders for drugs exhibit unequal performances across the different datasets, which suggests their hybrid usage for generalization purpose. ❸ Inspired from these insights, we conclude with a model combos that leads us to attaining state-of-the-art (SOTA) regression results and performing similarly to SOTA in the DTI classifications. Our combos further deliver cost-effective memory usage and running time as well as faster convergence, which can serve as new baseline for the following explorations.
2 Formulations for Drug-target Interaction Modeling
In this research, we focus on the formulations of recently-emerging structure modeling approaches for drug molecules, which could be categorized into explicit methods based on graph neural networks and implicit methods based on transformer. The target proteins are learned by the sophisticated tools of convolutional/recurrent neural networks (CNNs/RNNs) or transformers, after which both the molecules’ and proteins’ embeddings are integrated to facilitate interaction prediction. We will also summarize and benchmark the various widely-adopted molecule features.
2.1 Graph Neural Networks based Methods
A drug molecule is typically represented as a graph , where and denotes the sets of atoms and chemical bonds, respectively. The classical GNN frameworks involve key processes of aggregating and updating node features, collectively referred to as message passing, which can be mathematically represented as [26, 27]:
where is the feature representation of node at layer , is the feature representation of edge between nodes and , refers to the set of neighboring nodes next to node . Functions and aim to aggregate the neighborhood representations and integrate them together with the nodes features, respectively. Additionally, and are feature mapping functions, parameterized by and , respectively. The molecule’s representation can be derived using READOUT function, which processes on the set of vertex features at the last layer.
Graph Convolutional Networks (GCN). Given a molecule with atoms, the adjacency matrix indicates its connectivity, with if atom is adjacent to atom , and otherwise. Considering the self-connection of atoms, we have . Let denote the initial atom feature matrix. GCN [28] models the message passing as follows:
where is the node feature matrix at layer , starting with . Matrix represents the learnable weights for layer , denotes a non-linear activation function, e.g., ReLU, and is a diagonal degree matrix of . A couple of pioneering works have leveraged GCN to facilitate drug-protein interaction prediction [29, 30, 31, 32]. For example, DeepGLSTM [29] uses mixture-of-depths GCNs to capture drug representations from different scales. CPI [31] considers cross-atom distance and introduces the concept of r-radius subgraphs [33], using r-radius vertices and edges to redefine the structure of graphs.
Graph Isomorphism Networks (GIN). GIN excels in learning distinct graph features by approximating the Weisfeiler-Lehman test, enabling it to distinguish a wide range of graph structures [34]. The message passing process at the -th layer is of the following form:
where is a multi-layer perceptron that parameterizes the update function, and is a learnable parameter. We benchmark several GIN-based drug-target interaction modeling methods. GraphCPI [35] and GraphDTA [18] adopt GIN-based models with batch normalization to obtain the drug representation. SubMDTA [36] uses subgraph’s generation task and contrastive learning to pretrain a molecular graph encoder with multiple GIN layers for further prediction.
Graph Attention Networks (GAT). Unlike fixed-weight aggregation, GAT [37] employs an attention mechanism to determine neighborhood importance and learn the node embeddings as:
2.2 Transformer-based Methods
Besides the graph representation, drugs could also be decorated as SMILES strings [39] and encoded similarly to natural language processing. Specifically, after tokenizing SMILES strings, Transformer model utilizes multi-head attention to model the interactions between different segments of the input and obtain the molecular representations. Positional encodings are also integrated to preserve the sequence order, enhancing the model’s ability to process sequential information effectively. We review and benchmark two typical types of attention mechanisms used for molecular representations.
Self-Attention [20, 40, 41]. Self-attention computes a weighted sum of all input values based on their relevance to each other. Considering an embedding of SMILES sequence at a specific transformer layer, where and are token length and dimension, respectively, the attention is calculated by . and are projections of the input matrix . Multi-head attention combines these projections across different subspaces for a more detailed analysis. Following by normalization and feed-forward neural networks, the SMILES embedding is updated as and the output from the last layer is treated as molecular representations. Transformer encoders like MolTrans [20] and FOTFCPI [41] are adopted to enhance sub-structure embeddings in proteins and drugs.
Cross Attention [42, 40]. Cross-attention is designed to capture the interaction between the drug and protein sequences, with the query matrix derived from one sequence and the key and value matrices from another. This mechanism is particularly useful in integrating hybrid representations such as drug graphs and SMILES [38], as well as drugs and proteins [32, 42].
2.3 Feature Processing Methods
Beyond the drugs’ structure or sequence learning with GNNs or Transformers, the extra molecular properties, such as molecular weight, solubility, and lipophilicity, are crucial for building accurate and quantitative drug-target relationship models. We summarize two typical featurization methods.
Sequence Processing Methods. Both drugs and proteins are input as strings of ASCII characters, whose features can be extracted using statistical solutions. Integer encoding [18] simply converts the string to a sequence of integers, which assigns an integer to each character. The N-gram [43] captures the statistical dependencies between characters in an input string. Specifically, a 3-gram model breaks down a sequence into , analyzing the relationship between adjacent characters.
Drug-unique Featurization Methods. The additional chemical properties and structural details of SMILES strings are often considered to gain a more comprehensive understanding. Extended-Connectivity Fingerprints (ECFP) [44, 45], involves generating unique identifiers for atoms based on their local chemical environment and iteratively updating these through a hash function to capture a broader molecular context, ultimately producing a set of fingerprints that represent the molecule’s overall structure. Another approach, RDKit, is used to convert SMILES into molecular graphs [46, 18], where nodes represent the physical and chemical properties of molecules, and bonds are represented by an adjacency matrix. For example, atomic properties such as atom type, degree, and hydrogen information (like the number of explicit hydrogens) are all crucial for constructing a graph. More detailed properties can be found in Appendix F.
Embedding Featurization Methods. Embedding methods are used to translate these discrete sequences into continuous embedding spaces. Notably, Smi2Vec [47] and Prot2Vec [48] convert discrete tokens of drug SMILES and protein sequences into vectors that encapsulate semantic and syntactic similarities, effectively grouping similar tokens together in vector space. Additionally, pretrained language models [24, 49] are increasingly utilized to leverage large-scale learned patterns, fine-tuned to analyze complex protein data representations effectively.
3 A Fair Benchmark Platform Setup
Benchmark Model and Dataset Selection. From the perspective of reproducibility, we restrict our analysis to models for which the source code has been publicly released. To enhance the comprehensiveness, credibility, and sophistication of our benchmark, we conduct experiments on more than 30 models, including both GNN-based and Transformer-based methods. These models are derived from papers spanning the years 2018 to 2024.
We run these models on 6 frequently evaluated datasets including both binary interaction classification and continuous affinity regression. For the classification aspect, we utilize datasets including Human [50], Caenorhabditis elegans (C. elegans) [31], and DrugBank [51]. For regression, we employ the Davis [52], KIBA [53], and BindingDB [54] with dissociation constant (Kd) measures datasets, as processed in [22]. The statistical details of these models and datasets are presented in Section B and Table 4 of Appendix, respectively.
Unifying Hyperparameter Configuration. Given the critical role of hyperparameters in achieving optimal performance, we perform a detailed review of the hyperparameters associated with the selected models in Section E of the Appendix. There is significant variability in the hyperparameters across different models, making it unfair to conduct comparisons directly. To achieve equitable comparisons between varied models, we select two representative approaches from both the GNN-based and Transformer-based categories, i.e., GraphDTA [18], GraphCPI [35], MRBDTA [55], and TransformerCPI [21], to perform a greedy hyperparameter search to find their sweet spot for classification and regression tasks, respectively. For the search space of hyperparameters, we mainly focus on the influence of batch size (BS), learning rate (LR), and dropout rate (DR), as these are the common hyperparameters utilized by all models. Additionally, we standardize the hyperparameters for epochs, weight decay, and the choice of optimizer, setting a consistent 1000 epochs for GNN-based methods and 300 epochs for Transformer-based methods, with a weight decay of 0 and the Adam optimizer for all models according to Table 5. We illustrate the selected results for the metrics MSE and CI for the regression task, along with AUC-ROC and accuracy for the classification task, in Fig. 1 and the results of all metrics in Table 6. In all experiments, we employ the five-fold cross-validation method with a random split to evaluate all different methods and report the averaged results.
We observe that different models exhibit distinct preferences for hyperparameters. Taking into account various models and metrics, we recommend the configuration as the sweet point hyperparameter configuration for the GNN-based model. Similarly, for the Transformer-based model, we suggest . We strictly follow it in the following experiments.
4 A Macroscopical Benchmark on Encoder and Featurization Strategies
Encoder Exploration for Drugs and Proteins. To investigate the influence of different encoding strategies for extracting the structural information of drugs, we employ GCN [17] and vanilla Transformer [56] as the encoders for drugs. Meanwhile, integer encoding with CNN, n-gram encoding with CNN, and the vanilla Transformer are considered to capture protein’s representations, which are frequently adopted. To leverage the advantages of the pretrained protein information, we include a language model, i.e., Evolutionary Scale Modeling (ESM2) [49]. The results of various combinations of drug and protein encoders are listed in Table 1.
Obs. 1. GNN and Transformer-based drug encoders exhibit unequal performance depending on DTI tasks. When the encoder for the protein sequence is fixed, drug features extracted by the Transformer generally perform better than those by GNNs in regression tasks, but the opposite is true in classification tasks. Notably, in the classification tasks on Human dataset, the combination of GNN and Transformer used, respectively, for drugs and proteins yields excellent performance but falls short in the regression task. This disparity may be due to the smaller size of the Human dataset compared to the Davis dataset, which allows for faster convergence in classification tasks than in regression tasks under a fixed epoch.
Obs. 2. Transformer models are better in extracting features from protein. Although we only consider the simplest pretrained protein language model of ESM2, it still significantly outperforms other encoders. This improvement can likely be attributed to the robust and generalizable representations learned from extensive data by the pretrained model. In the classification task, transformers achieve the best performance, underscoring their effectiveness in extracting protein sequence features.
Obs 3. Integer encoding appears to be more effective when paired with a CNN as the protein encoder and a fixed drug encoder. Compared to this specific model configuration, the local context provided by 3-gram encoding does not significantly enhance the model’s predictive performance. This implies that the simple relationships in amino acids’ immediate neighbors, as modeled by Word2Vec, do not capture much useful information compared with simple integer encoding.
| Drug Encoder | Protein Encoder | Regression | Classification | |||||||||||
| MSE | MAE | R2 | PCC | CI | Spearman | ROC-AUC | PR-AUC | Log-AUC | Acc. | Precision | Recall | F1 | ||
| GCN | Int encoding with CNN | 0.3181 | 0.3314 | 0.5263 | 0.7315 | 0.8311 | 0.5906 | 0.9122 | 0.8761 | 0.4452 | 0.9127 | 0.9033 | 0.9301 | 0.9165 |
| GCN | 3-gram pretrained with CNN | 0.3361 | 0.3613 | 0.4994 | 0.7133 | 0.8332 | 0.5956 | 0.9016 | 0.8634 | 0.4045 | 0.9021 | 0.8944 | 0.9183 | 0.9061 |
| GCN | ESM2 | 0.2611 | 0.3112 | 0.6111 | 0.7843 | 0.8665 | 0.6480 | 0.9331 | 0.9046 | 0.5034 | 0.9334 | 0.9275 | 0.9446 | 0.9371 |
| GCN | Trans | 0.2911 | 0.3326 | 0.5664 | 0.7600 | 0.8571 | 0.6329 | 0.9435 | 0.9192 | 0.5282 | 0.9438 | 0.9395 | 0.9524 | 0.9458 |
| Trans | Int Encoding with CNN | 0.2553 | 0.2853 | 0.6197 | 0.7890 | 0.8620 | 0.6397 | 0.9096 | 0.8783 | 0.4623 | 0.9096 | 0.9142 | 0.9099 | 0.9121 |
| Trans | 3-gram pretrained with CNN | 0.2618 | 0.3032 | 0.6101 | 0.7820 | 0.8581 | 0.6338 | 0.9011 | 0.8674 | 0.4327 | 0.9012 | 0.9054 | 0.9030 | 0.9040 |
| Trans | ESM2 | 0.2609 | 0.3052 | 0.6115 | 0.7869 | 0.8617 | 0.6391 | 0.9203 | 0.8924 | 0.4902 | 0.9203 | 0.9257 | 0.9192 | 0.9224 |
| Trans | Trans | 0.2828 | 0.3222 | 0.5788 | 0.7721 | 0.8554 | 0.6297 | 0.9306 | 0.9073 | 0.5158 | 0.9304 | 0.9391 | 0.9250 | 0.9319 |
Featurization Exploration. Despite the efficacy of GNNs in learning drug structures, the featurization of nodes plays a critical role in capturing both the intrinsic properties of atoms and their contextual relevance. We conduct a detailed analysis of various methods (summarized in Section F of Appendix) for constructing graph features within the DTI context. The node feature is constructed via various characteristics, such as chemical and physical properties. We categorize each feature into five main classes, e.g., atomic properties (AP), hydrogen information (HI), electron properties (EP), stereochemistry (Ste) and structural information (Str). To better determine which types of features are more effective in capturing the structural information, we conduct an ablation study on the different featurization strategies. Here we choose GprahDTA [18] and GraphCPI [35] as our backbone models. The results of feature combinations are reported in Fig. 2.
Obs. 4. More complex featurization does not necessarily bring positive effect. Despite employing different protein representations (different colors in Fig. 2), GraphCPI and GraphDTA perform stably on the benchmark dataset Davis. Moreover, the basic feature configuration has the lowest MSE compared to other featurization strategies, suggesting that additional features may complicate the model’s ability to discern critical information for regression task. Additionally, the increased CI with more complex features, such as stereochemistry and structure, suggests that while they might not improve the prediction accuracy, they do contribute positively to the model’s ranking capability.
Obs. 5. The role of atomic and electron properties in modeling drug features may not be inherently detrimental, but their contribution appears to be context-dependent. In Fig. 2, it suggests that these properties, when not combined with other informative features, don’t significantly enhance two models’ performance. In the classification task, as reflected by the F1 scores and ROC-AUC, it is evident that stereochemistry and structural information substantially improve the model’s accuracy. Thus, while atomic and electron properties are fundamental, their full potential is unlocked when integrated with stereochemical and structural information, underscoring the importance of a multifaceted approach in node featurization.
5 A Microscopic Benchmark on DTI Models
Benchmark over Effectiveness. As shown in Table. 2 and Table. 3, we conduct experiments on 31 different models across two tasks and three datasets, respectively (see more comprehensive model comparisons in Section G of Appendix). All results are averaged by five-fold cross-validation.
Obs. 6. Molecular graphs are better than fingerprints to capture the graph features of drug. In reference to Table. 7, it is evident that GNN-based approaches utilizing the molecular graph generally yield superior performance compared with fingerprints (CPI [31], BACPI [57], GANDTI [58]). This reinforces the idea that the rich structural and atom property information inherent to molecular graphs is pivotal for representation extraction, leading to enhanced model performance.
Obs. 7. Graph structure is a crucial part in extracting drug’s features. Different GNNs have the distinct performances in both tasks when the protein representation is fixed. Specifically, GIN, with its unique ability to distinguish non-isomorphic graphs, consistently outperforms other models across different protein encoders in regression tasks. Although transformer-based methods such as MRBDTA are proficient in handling sequential information from SMILES and proteins, the depth of information they capture appears to be marginally less comprehensive than that provided by molecular graph-based approaches. This is substantiated by the superior performance of GNN-based methods, including MGraphDTA, ColdDTA, and SubMDTA, which suggest that GNN captures intricate structural details more effectively.
| Category | Models | DAVIS | KIBA | BindingDB_Kd | Avg. Reduce | ||||||
| MSE | R2 | CI | MSE | R2 | CI | MSE | R2 | CI | of MSE (%) | ||
| GNN | GraphDTA-GIN [18] | 0.2309 | 0.6562 | 0.8711 | 0.0004 | 0.5544 | 0.7996 | 0.5033 | 0.7259 | 0.8576 | 8.5625 |
| GraphCPI-GIN [35] | 0.2413 | 0.6407 | 0.8671 | 0.0005 | 0.4963 | 0.7808 | 0.5069 | 0.7239 | 0.8572 | 10.2845 | |
| MGraphDTA [59] | 0.2179 | 0.6755 | 0.8820 | 0.0003 | 0.7208 | 0.8649 | 0.4887 | 0.7338 | 0.8649 | 4.9795 | |
| SAGDTA [60] | 0.2656 | 0.6045 | 0.8675 | 0.0039 | -3.3777 | 0.8096 | 0.6590 | 0.6410 | 0.8366 | 27.6575 | |
| EmbedDTI [61] | 0.2561 | 0.6186 | 0.8624 | 0.0007 | 0.2364 | 0.6374 | 0.5095 | 0.7225 | 0.8559 | 12.3450 | |
| DeepGLSTM [29] | 0.2915 | 0.5659 | 0.8476 | 0.0003 | 0.6919 | 0.8480 | 0.5385 | 0.7067 | 0.8529 | 19.1015 | |
| CPI [31] | 0.3503 | 0.4784 | 0.8319 | 0.0003 | 0.6274 | 0.8190 | 0.6962 | 0.6208 | 0.8261 | 35.8330 | |
| BACPI [57] | 0.4036 | 0.3990 | 0.7982 | 0.0006 | 0.5320 | 0.8175 | 0.6468 | 0.6477 | 0.8297 | 36.0894 | |
| DeepNC-HGC [30] | 0.2782 | 0.5857 | 0.8551 | 0.0005 | 0.4983 | 0.7823 | 0.5611 | 0.6944 | 0.8464 | 20.0167 | |
| DeepNC-GEN [30] | 0.2543 | 0.6213 | 0.8634 | 0.0004 | 0.5355 | 0.7981 | 0.5561 | 0.6971 | 0.8496 | 17.1559 | |
| DrugBAN [62] | 0.2391 | 0.6440 | 0.8757 | 0.0004 | 0.5611 | 0.8388 | 0.4485 | 0.7557 | 0.8693 | 2.3692 | |
| GANDTI [58] | 0.3082 | 0.5410 | 0.8414 | 0.0003 | 0.6569 | 0.8342 | 0.6714 | 0.6343 | 0.8322 | 31.4522 | |
| PGraphDTA-CNN [24] | 0.3273 | 0.5126 | 0.8701 | 0.0005 | 0.4833 | 0.7473 | 0.5334 | 0.7095 | 0.8590 | 22.0042 | |
| BridgeDPI [63] | 0.3623 | 0.6477 | 0.8991 | 0.0004 | 0.5686 | 0.7849 | 0.4482 | 0.7559 | 0.8698 | 17.1661 | |
| ColdDTA [64] | 0.2346 | 0.6507 | 0.8693 | 0.0004 | 0.5948 | 0.8018 | 0.4697 | 0.7442 | 0.8644 | 4.6828 | |
| SubMDTA [36] | 0.2326 | 0.6537 | 0.8691 | 0.0003 | 0.6855 | 0.8485 | 0.4566 | 0.7513 | 0.8670 | 2.5816 | |
| IMAEN [65] | 0.2412 | 0.6409 | 0.8721 | 0.0004 | 0.5800 | 0.8061 | 0.4720 | 0.7429 | 0.8553 | 5.8716 | |
| Transformer | CSDTI [32] | 0.3029 | 0.5490 | 0.8395 | 0.0007 | 0.2448 | 0.6475 | 0.6408 | 0.6510 | 0.8369 | 28.8755 |
| TDGraphDTA [25] | 0.2217 | 0.6698 | 0.6685 | 0.0008 | 0.0533 | 0.3429 | 0.4750 | 0.7413 | 0.7887 | 3.6989 | |
| AMMVF [38] | 0.3325 | 0.5048 | 0.8307 | 0.0006 | 0.4696 | 0.7711 | 0.6597 | 0.6407 | 0.8336 | 32.3429 | |
| IIFDTI [66] | 0.2741 | 0.5918 | 0.8500 | 0.0005 | 0.2584 | 0.7952 | 0.5097 | 0.7170 | 0.8585 | 14.3568 | |
| ICAN [42] | 0.3481 | 0.4816 | 0.8211 | 0.0008 | 0.1159 | 0.8256 | 0.6582 | 0.6415 | 0.8277 | 33.3035 | |
| MolTrans [20] | 0.2588 | 0.6146 | 0.8601 | 0.0003 | 0.6378 | 0.8453 | 0.5138 | 0.7201 | 0.8570 | 13.0935 | |
| TransformerCPI [21] | 0.2869 | 0.5728 | 0.8326 | 0.0008 | 0.0728 | 0.8357 | 0.5704 | 0.6894 | 0.8426 | 21.7224 | |
| MRBDTA [55] | 0.2350 | 0.6499 | 0.8775 | 0.0006 | 0.3210 | 0.7239 | 0.4977 | 0.7289 | 0.8629 | 8.4004 | |
| FOTFCPI [41] | 0.2803 | 0.5825 | 0.8546 | 0.0004 | 0.5342 | 0.7947 | 0.5743 | 0.6872 | 0.8444 | 21.4386 | |
| Our combos | 0.2063 | 0.6927 | 0.8901 | 0.0003 | 0.7168 | 0.8677 | 0.4651 | 0.7467 | 0.8683 | 0 | |
| Category | Models | Human | C.elegans | Drugbank | Avg. Improve | ||||||
| ROC-AUC | Accuracy | F1 | ROC-AUC | Accuracy | F1 | ROC-AUC | Accuracy | F1 | of Accuracy (%) | ||
| GNN | GraphDTA-GATGCN[18] | 0.9296 | 0.9297 | 0.9319 | 0.9487 | 0.9484 | 0.9485 | 0.7712 | 0.7711 | 0.7764 | 2.5140 |
| GraphCPI-GATGCN[35] | 0.9097 | 0.9099 | 0.9129 | 0.9372 | 0.9370 | 0.9369 | 0.7561 | 0.7561 | 0.7593 | 4.3335 | |
| MGraphDTA[59] | 0.9408 | 0.9410 | 0.9429 | 0.9631 | 0.9630 | 0.9628 | 0.8146 | 0.8146 | 0.8157 | -0.1030 | |
| SAGDTA[60] | 0.9021 | 0.9018 | 0.9032 | 0.9380 | 0.9378 | 0.9375 | 0.7655 | 0.7655 | 0.7662 | 4.2494 | |
| EmbedDTI[61] | - | - | - | - | - | - | 0.7625 | 0.7625 | 0.7674 | 5.3770 | |
| DeepGLSTM[29] | 0.9180 | 0.9183 | 0.9212 | 0.9414 | 0.9412 | 0.9409 | 0.7640 | 0.7640 | 0.7674 | 3.5182 | |
| CPI[31] | 0.9110 | 0.9111 | 0.9138 | 0.9312 | 0.9310 | 0.9309 | 0.7467 | 0.7466 | 0.7515 | 4.9098 | |
| BACPI[57] | 0.9249 | 0.9251 | 0.9276 | 0.9556 | 0.9555 | 0.9551 | 0.7748 | 0.7748 | 0.7776 | 2.2746 | |
| DeepNC-HGC[30] | 0.8796 | 0.8798 | 0.8834 | 0.9418 | 0.9417 | 0.9413 | 0.7673 | 0.7673 | 0.7718 | 4.9057 | |
| DeepNC-GEN[30] | 0.9178 | 0.9180 | 0.9205 | 0.9501 | 0.9499 | 0.9498 | 0.7464 | 0.7463 | 0.7571 | 3.8865 | |
| DrugBAN[62] | 0.9302 | 0.9302 | 0.9320 | 0.9596 | 0.9594 | 0.9593 | 0.8188 | 0.8188 | 0.8179 | 0.2732 | |
| GANDTI [58] | 0.9333 | 0.9333 | 0.9351 | 0.9404 | 0.9403 | 0.9400 | 0.7427 | 0.7425 | 0.7550 | 3.8110 | |
| PGraphDTA-CNN [24] | - | - | - | - | - | - | 0.7910 | 0.7910 | 0.7948 | 1.5803 | |
| BridgeDPI [63] | 0.9456 | 0.9456 | 0.9471 | 0.9651 | 0.9651 | 0.9646 | 0.7825 | 0.7825 | 0.7853 | 0.8392 | |
| ColdDTA [64] | 0.9420 | 0.9421 | 0.9436 | 0.9623 | 0.9622 | 0.9619 | 0.8186 | 0.8186 | 0.8203 | -0.2608 | |
| SubMDTA [36] | 0.9326 | 0.9328 | 0.9351 | 0.9570 | 0.9568 | 0.9566 | 0.8045 | 0.8045 | 0.8072 | 0.8055 | |
| IMAEN [65] | 0.9058 | 0.9059 | 0.9089 | 0.9506 | 0.9505 | 0.9500 | 0.7995 | 0.7994 | 0.8014 | 2.2592 | |
| Transformer | CSDTI [32] | 0.8630 | 0.8630 | 0.8663 | 0.8962 | 0.8958 | 0.8966 | 0.7269 | 0.7269 | 0.7306 | 9.2569 |
| TDGraphDTA [25] | 0.9411 | 0.9409 | 0.9419 | 0.9573 | 0.9571 | 0.9569 | 0.8083 | 0.8082 | 0.8116 | 0.3547 | |
| AMMVF [38] | 0.9287 | 0.9290 | 0.9314 | 0.9636 | 0.9635 | 0.9632 | 0.7814 | 0.7814 | 0.7849 | 1.5670 | |
| IIFDTI [66] | 0.9392 | 0.9392 | 0.9409 | 0.9679 | 0.9679 | 0.9675 | 0.8084 | 0.8084 | 0.8125 | 0.0110 | |
| ICAN [42] | 0.9387 | 0.9389 | 0.9410 | 0.9644 | 0.9644 | 0.9637 | 0.7731 | 0.7731 | 0.7745 | 1.4721 | |
| MolTrans [20] | 0.9453 | 0.9455 | 0.9471 | 0.9639 | 0.9638 | 0.9634 | 0.7910 | 0.7909 | 0.7922 | 0.5777 | |
| TransformerCPI [21] | 0.9311 | 0.9312 | 0.9333 | 0.9581 | 0.9580 | 0.9577 | 0.8097 | 0.8097 | 0.8125 | 0.6262 | |
| MRBDTA [55] | 0.9447 | 0.9446 | 0.9458 | 0.9713 | 0.9713 | 0.9709 | 0.8102 | 0.8102 | 0.8134 | -0.3778 | |
| FOTFCPI [41] | 0.9444 | 0.9444 | 0.9459 | 0.9673 | 0.9672 | 0.9669 | 0.7845 | 0.7845 | 0.7889 | 0.7307 | |
| Our combos | 0.9435 | 0.9435 | 0.9451 | 0.9688 | 0.9688 | 0.9684 | 0.8035 | 0.8035 | 0.8061 | 0 | |
Benchmark over Efficiency To analyze the training speed and memory usage, we empirically evaluate the peak memory and running time for various methods during the training procedure on one regression dataset and one classification task, respectively. To fairly compare various methods, we set the batch size as 32, as such maximum batch size is adopted by some methods. All results are measured on an RTX 3090 GPU. The memory and running time comparisons are illustrated in Fig. 3.
Obs. 8. In general, the memory usage of GNN-based methods is smaller than that of Transformer-based methods, which is positively proportional to run time. This difference is primarily due to the self-attention mechanism employed in Transformers, which requires significant memory resources. In contrast, model parameters, such as those in DeepGLSTM, do not exhibit a direct relationship with either runtime or performance.
Benchmark over Convergence. We select the top two methods from the GNN-based and Transformer-based frameworks, respectively, and evaluate them across six datasets on two tasks. The training losses are depicted in Fig. 4. In order to compare different methods, we only show the epochs before 300. Based on the empirical data, we summarize our primary observations as follows:
Obs. 9. GNN-based methods demonstrate quicker and more stable convergence compared to Transformer-based methods. This phenomenon arises from the fact that GNN-based methods have fewer memory usage and model parameters, leading to a larger batch size usage or faster convergence compared with Transformer-based methods.
5.1 Our Best Combo of Drug and Protein Encoders
Based on our benchmark results, we summarize the insights of protein and drug encoder usages and propose a light yet effective architecture, which could be treated as new strong baseline for the following explorations. Regarding the proteins, we observe that multi-scale CNNs associated with a mixture of model depths can generally learn the effective protein representations [59, 25, 64], which approximates the language model’s accuracy while having lower memory and computation costs.
Regarding the drug molecules, both GNN and Transformer-based methods, such as MRBDTA [55], MolTrans [20] and MGraphDTA [59] prove promising in DTI tasks. This encourages us to leverage information from hybrid perspectives, i.e., implicit structure (via attention in Transformers) and explicit structure learning (via message passing along edges in GNNs).
We are thus motivated to integrate these powerful modules and shed novel insight into the design philosophy of drug-target interaction modeling. Our model combos are illustrated in Fig. 5, where the multi-scale CNNs and hybrid networks of molecular Transformer and GNNs are adopted to learn the representations of proteins and drugs, respectively. As shown by the graph encoder part in Fig. 5, the hybrid networks augment the differential attention matrix in molecular Transformer with inter-atomic distances and graph adjacency matrix [67], which provides the 3D and 2D molecule conformations to further facilitate atom interaction learning.
Specifically, given the projections of molecular input at an attention head, i.e., , the adjacent matrix , and the inter-atomic distances matrix obtained using RDkit, the augmented attention is calculated as:
where is a row-wise softmax function, and and denote scalars weighting the self-attention, distance, and adjacency matrices, respectively. Besides the implicit and explicit structure learning, we integrate the features from drug SMILES. It is notable that simply utilizing the SMILES representation extracted from a transformer for downstream tasks does not perform as well as GNN. To align with the protein embedding paradigm, we adopt a simple CNN to unearth potential SMILES information, as suggested in [68]. Subsequently, due to the fact that cross-attention is more complex and hard to optimize, we implement a straightforward attention mechanism to integrate the representations of the drug graph and SMILES, denoted as and , respectively, using a weighting parameter , as follows:
Finally, the prediction is obtained by processing the concatenated protein and drug representations through a task-relevant head, as shown in Fig. 5
Benchmark Comparison to State-of-the-Art Frameworks. We compare the proposed combos with the SOTA frameworks in Tables 2 and 3, and Figures 3 and 4. It is observed that our model consistently achieves the best performance in the regression tasks across three datasets and nearly outperforms most methods in classification tasks. By leveraging the physical conformation information from the molecular graph, our combos converge faster than the other two Transformer-based methods, MRBDTA [55] and TDGraphDTA [25], particularly on the KIBA dataset. Moreover, our model uses three times less peak memory and fewer parameters than other Transformer-based methods, enabling faster computation and reduced storage requirements
6 Conclusion
In this work, we establish a benchmark with fair and consistent experimental configurations, aiming to push DTI research, particularly emphasizing the utilization of structural information. Our meticulous approach has entailed thorough exploration of diverse encoder strategies and featurization techniques for both drug molecules and proteins. Moreover, dozens of existing approaches across six representative datasets for both regression and classification tasks are investigated on various metrics, including DTI classification and regression accuracy, peak memory usage, and model convergence. Provided with the comprehensive benchmark results, we propose a novel approach that integrates the strengths of GNN and transformer-based methods. Our studies on benchmarking and rethinking help lay a solid, practical, and systematic foundation for the DTI community and provide researchers with broader and deeper insights into the intricate dynamics of drug-target interactions.
Appendix A Related Works
GNN-based Methods GNNs play a crucial role in mining the intricate features of drug molecules for drug-target prediction. Numerious models, including Graph Convolutional Network (GCN), Graph Isomorphism Network (GIN), and Graph Attention Network (GAT) have been utilized [18, 35, 38, 69, 69, 61] to process and enhance drug features. Additionally, MGraphDTA [59] employs a multi-scale GNN architecture, while DeepGLSTM [29] leverages parallel GNN structures for drug representation. DeepNC integrates advanced techniques from generalized aggregation networks [70] and hypergraph convolution [71] to improve feature extraction. BACPI [57] develops a bi-directional attention network to integrate the representations of drug molecules and proteins, enhancing their mutual interaction. Besides, BridgeDPI [63] innovates by incorporating bridging nodes between proteins and drugs, utilizing a three-layer GNN for graph embeddings.
Transformer-based Methods Transformers, known for their efficacy in handling sequence data, are extensively applied in drug and protein feature processing. For instance, models like MolTrans [20] and FOTFCPI [41] employ self-attention mechanisms to refine embeddings by focusing on drug and protein substructures. MRBDTA [55] uses multi-head attention and skip connection to enhance drug and protein representation. Additionally, a cross-attention mechanism [32, 42] is employed to facilitate the integration of drug and protein features, enabling effective mutual querying. TDGraphDTA [25] captures contextual relationships between molecular substructures by using a multi-head cross-attention mechanism and graph optimization. Lastly, DrugormerDTI [72] incorporates degree centrality with positional information to highlight the positional relevance of amino acids in proteins.
Input and Featurization Structural information is crucial at the input stage for models such as BridgeDPI [63]. Various libraries, such as DGLGraph [73], DGL-lifeSci [74], and RDKit [46], are employed to process input SMILES of drugs, with RDKit [46] being pivotal for converting SMILE strings into molecular graphs and extracting diverse chemical properties, including chemical bonds, hydrogen presence, electron properties, and so on. Additionally, some approaches [38, 69, 57, 58] incorporate molecular fingerprints [45] to capture local chemical information. For protein sequences, typical pre-processing involves converting amino acid sequences into N-grams [36, 43] or integers [18] sequences. To enhance the expressiveness of embeddings, some models leverage pre-trained Word2Vec [75, 35, 38, 57, 31, 69, 66] or pretrained protein language models [24].
Appendix B Model Descriptions
This section provides a comprehensive overview of 31 DTI methods, which are classified into GNN-based and Transformer-based approaches. The DTI framework can be simplified as using two encoders to process drugs and proteins separately, followed by an MLP to handle the integrated representations.
B.1 GNN-based Methods
B.1.1 GCN
GraphDTA-GCN [18]: GraphDTA-GCN uses GCN to process the molecular graph, which is derived from SMILES using the RDkit tool, and a simple CNN with integer encoding to handle protein sequences.
GraphCPI-GCN [35]: Similar to GraphDTA, GraphCPI-GCN employs 3-gram encoding with pretrained Word2Vec to process protein sequences, followed by a CNN to handle the protein embeddings.
MGraphDTA [59]: MGraphDTA utilizes a multiscale GCN, inspired by dense connections, and a multiscale CNN to process drug graphs and protein sequences, respectively.
SAGDTA [60]: Similar to GraphDTA, SAGDTA introduces global or hierarchical pooling after GCN to aggregate node representations weightedly.
EmbedDTI [61]: For protein sequences, EmbedDTI leverages GloVe for pretraining amino acid feature embeddings, which are then fed into a CNN. For drugs, it constructs both an atom graph and a substructure graph to capture structural information at different levels, processed by GCN.
DeepGLSTM [29]: DeepGLSTM processes molecular graphs using a parallel GCN module composed of three GCNs with different layers. For protein sequences, it adopts a bi-LSTM.
CPI [31]: CPI processes drug graphs using GCN. The protein sequence is handled via n-gram with integer encoding, followed by a CNN.
DeepNC [30]: DeepNC adopts advanced techniques from generalized aggregation networks and hypergraph convolution, two variants of GCN, to capture the representations of drug. For protein sequences, it uses a simple CNN.
DrugBAN [55]: DrugBAN employs GCN and CNN blocks to encode molecular graph and proteins, respectively. Then they use a bilinear attention network module to learn local interactions between the representations of drugs and proteins.
BridgeDPI [63]: BridgeDPI innovates by constructing a learnable drug–protein association network, which is processed using a three-layer GNN for graph embeddings. The learned representations for drug and protein pairs are then concatenated for further processing.
ColdDTA [64]: ColdDTA removes the subgraphs of drugs. For the model, they adopt the dense GCN and multiscale CNN from MGraphDTA as the encoders for drugs and proteins, respectively. Additionally, an attention-based method is developed to integrate representations for improved prediction.
IMAEN [65]: IMAEN employs a molecular augmentation mechanism to enhance molecular structures by fully aggregating molecular node neighborhood information. It then uses multiscale GCN and CNN for drug and protein processing, respectively.
GanDTI [58]: Inspired by residual networks, GanDTI add the input drug fingerprints to the output of three GCN layers as graph node features and use summation to get the final drug representation.
B.1.2 GAT
GraphDTA-GAT [18]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphDTA-GATGCN [18]: GraphDTA-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphCPI-GAT [35]: GraphDTA-CPI adopts a GAT as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
GraphCPI-GATGCN [35]: GraphCPI-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
BACPI [57]: BACPI adopts a GAT and a CNN for the features of the fingerprints and protein sequence, respectively. These features are then fed into a bi-directional attention neural network to obtain integrated representations.
PGraphDTA-CNN [24]: PGraphDTA-CNN is a straightforward method that utilizes GAT for drug feature extraction and CNN for protein sequences.
B.2 GIN
GraphDTA-GIN [18]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphCPI-GIN [35]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
SubMDTA [36]: SubMDTA utilizes a pretrained GIN encoder obtained through contrastive learning for the molecular graph. For protein sequences, it employs N-gram embedding with different N to extract features at various scales, which are then processed by a BiLSTM.
B.3 Transformer-based Methods
B.3.1 Self-attention
AMMVF [38]: AWMVF introduces the multi-head mechanism to GAT to learn features in different spaces, and the update function is obtained through the concatenation of different heads’ outputs.
IIFDTI [66]: IIFDTI model attains the drug matrix and protein matrix and inputs them to the bi-directional encoder-decoder block, which considers both the drug and target directions. The decoder is mainly composed of multi-head attention.
MolTrans [20]: MolTrans uses transformer encoder layers to augment the embedding of sub-structure sequences of proteins and drugs.
FOTFCPI [41]: Similar to MolTrans, FOTFCPI uses transformer encoder layers to extract the features of protein and drug fragments after the embedding layers.
TransformerCPI [21]: TransformerCPI uses the decoder module of Transformer, which takes in the atom sequence embedding processed by GCN and the protein sequence embedding processed by word2vec and 1D CNN.
MRBDTA [55]: In MRBDTA, after the embedding layer, drug sequences are directly fed into a block consisting of three Transformer encoders. The first encoder has a linear layer before it and the following two encoders are parallel. The protein sequence is also processed by a block with similar structure.
B.3.2 Cross attention
CSDTI[32]: CSDTI use cross attention to fuse the deep representations of drugs and proteins. Specifically, the different projections of protein feature are used as key and value respectively while the projection of drug feature is used as query.
TDGraphDTA[25]: TDGraphDTA use a multi-head cross-attention mechanism with two attention heads. Both drug and protein features are linearly transformed into query, key and value matrices. One cross attention layer uses a drug query matrix, a protein key matrix, and a protein value matrix, while its parallel counterparts use the rest of the matrices. The outputs of these two layers are concatenated and fed into MLP to get the final output.
Appendix C Datasets Descriptions
In this subsection, we provide a detailed description of the datasets for both the regression task and classification task. The statistical characteristics of the datasets are summarized in Table 4.
| Regression | Classification | |||||
| Davis | KIBA | BindingDB_Kd | Human | C. elegans | DrugBank | |
| Number of drugs | 68 | 2068 | 10661 | 2726 | 1767 | 6645 |
| Number of target proteins | 379 | 229 | 1413 | 2001 | 1876 | 4256 |
| Number of total samples | 25772 | 117657 | 52274 | 6728 | 7786 | 35021 |
Appendix D Evaluation Metrics
We adopt distinct sets of metrics to evaluate the classification and regression tasks. In particular, considering the classification task, we utilize the common metrics including Area Under Receiver Operating Characteristic Curve (ROC-AUC), Precision-Recall Area Under Curve (PR-AUC), LogAUC, accuracy, precision, recall, and F1 score. For the continuous binding affinity regression, we benchmark the models using metrics of mean squared error (MSE), mean absolute error (MAE), coefficient of determination (R2), Pearson correlation coefficient, Concordance Index (CI), and Spearman correlation coefficient. Each of these metrics offers unique insights into different aspects of model performance, allowing us to assess predictive accuracy, correlation with observed values, and consistency in ranking predictions.
Appendix E Greedy Search on Hyperparameter
Here we present the detailed experiment on hyperparameters with full metrics. The detailed review of the hyperparameters associated with the selected models is in Table 5. The detailed greedy hyperparameter seraching results on Davis and Human datasets are in Table 6.
| Category | Models | Batch size | Total epoch | Learning rate & Decay & Decay epoch | Weight decay | Dropout | Optimizer |
| GNN | GraphDTA-GCN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam |
| GraphDTA-GAT [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphDTA-GATGCN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphDTA-GIN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GCN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GAT [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GATGCN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GIN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| MGraphDTA [59] | 512 | 292 | 0.0005 | - | 0.1 | Adam | |
| SAGDTA [60] | 512 | 2000 | 0.001 | - | 0.1 | Adam | |
| EmbedDTI [61] | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
| DeepGLSTM [29] | 128 | 1000 | 0.0005 | - | 0.2 | Adam | |
| CPI [31] | 1 | 100 | {0.001, 0.5, 10} | 1e-6 | 0 | Adam | |
| BACPI [57] | 16 | 20 | {0.0005, 0.5, 10} | - | 0.1 | Adam | |
| DeepNC-HGC [30] | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
| DeepNC-GEN [30] | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
| DrugBAN [62] | 64 | 100 | 0.0005 | - | 0 | Adam | |
| GANDTI [58] | 1 | 30/15 | 0.001 | 1e-6 | 0.5 | Adam | |
| PGraphDTA-CNN [24] | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
| BridgeDPI [63] | 512 | 100 | 0.001 | - | 0.5 | Adam | |
| ColdDTA [64] | 128 | 292 | 0.003 | - | 0 | Adam | |
| SubMDTA [36] | 512 | 1200 | 0.0005 | - | 0.2 | Adam | |
| IMAEN [65] | 128 | 1000 | 0.0005 | - | 0.2 | Adam | |
| Transformer | CSDTI [32] | 256 | 292 | 0.0005 | - | {0.5,0.2} | Adam |
| AMMVF [38] | 32 | 40 | {0.001, 0.5, 5} | 1e-4 | 0.1 | Adam | |
| TDGraphDTA [25] | 1024 | 3000 | 0.0005 | - | 0.1 | Adam | |
| IIFDTI [66] | 64 | 100 | 0.001 | 1e-6 | 0.2 | AdamW | |
| ICAN [42] | 128 | 50 | 0.001 | - | 0.1 | Adam | |
| MolTrans [20] | 16 | 13 | 0.001 | - | 0.1 | Adam | |
| TransformerCPI [21] | 8 | 40 | 0.0001 | 0.0001 | 0.2 | Radam | |
| MRBDTA [55] | 256 | 300 | 0.001 | - | 0.1 | Adam | |
| FOTFCPI [41] | 64 | 100 | 0.0001 | - | 0.1 | Adam |
| Models | Hyper parameter | Regression | Classification | |||||||||||||
| Batch size | Learning rate | Dropout | MSE | MAE | R2 | PCC | CI | Spearman | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | |
| GraphDTA | 512 | 0.0005 | 0.2 | 0.2838 | 0.3053 | 0.5773 | 0.7619 | 0.8508 | 0.6226 | 0.9117 | 0.8807 | 0.4712 | 0.9117 | 0.9157 | 0.9125 | 0.9141 |
| 256 | 0.0005 | 0.2 | 0.2781 | 0.2924 | 0.5859 | 0.7677 | 0.8523 | 0.6250 | 0.9132 | 0.8805 | 0.4584 | 0.9135 | 0.9115 | 0.9218 | 0.9165 | |
| 128 | 0.0005 | 0.2 | 0.2634 | 0.2552 | 0.6077 | 0.7840 | 0.8700 | 0.6828 | 0.9134 | 0.8822 | 0.4716 | 0.9135 | 0.9157 | 0.9166 | 0.9160 | |
| 64 | 0.0005 | 0.2 | 0.2591 | 0.2506 | 0.6141 | 0.7885 | 0.8634 | 0.6862 | 0.9121 | 0.8793 | 0.4608 | 0.9123 | 0.9112 | 0.9195 | 0.9153 | |
| 512 | 0.0001 | 0.2 | 0.3209 | 0.3148 | 0.5221 | 0.7299 | 0.8369 | 0.6008 | 0.9165 | 0.8851 | 0.4775 | 0.9168 | 0.9159 | 0.9232 | 0.9195 | |
| 512 | 0.001 | 0.2 | 0.2909 | 0.3076 | 0.5668 | 0.7812 | 0.8425 | 0.6073 | 0.9133 | 0.8827 | 0.4765 | 0.9134 | 0.9172 | 0.9145 | 0.9157 | |
| 512 | 0.0005 | 0 | 0.2827 | 0.2923 | 0.5791 | 0.7632 | 0.8527 | 0.6267 | 0.9177 | 0.8863 | 0.4764 | 0.9180 | 0.9163 | 0.9252 | 0.9207 | |
| 512 | 0.0005 | 0.1 | 0.2771 | 0.2947 | 0.5873 | 0.7695 | 0.8521 | 0.6241 | 0.9170 | 0.8873 | 0.4830 | 0.9171 | 0.9204 | 0.9189 | 0.9194 | |
| 512 | 0.0005 | 0.5 | 0.2951 | 0.3165 | 0.5605 | 0.7494 | 0.8478 | 0.6176 | 0.9128 | 0.8819 | 0.4637 | 0.9129 | 0.9161 | 0.9151 | 0.9155 | |
| GraphCPI | 512 | 0.0005 | 0.2 | 0.3173 | 0.3309 | 0.5275 | 0.7325 | 0.8267 | 0.5825 | 0.9049 | 0.8687 | 0.4275 | 0.9053 | 0.9005 | 0.9177 | 0.9089 |
| 256 | 0.0005 | 0.2 | 0.2781 | 0.2924 | 0.5859 | 0.7677 | 0.8523 | 0.6250 | 0.9058 | 0.8712 | 0.4394 | 0.9061 | 0.9047 | 0.9140 | 0.9093 | |
| 128 | 0.0005 | 0.2 | 0.3064 | 0.2988 | 0.5436 | 0.7402 | 0.8467 | 0.6213 | 0.9085 | 0.8762 | 0.4602 | 0.9086 | 0.9115 | 0.9111 | 0.9112 | |
| 64 | 0.0005 | 0.2 | 0.3484 | 0.3569 | 0.4812 | 0.6944 | 0.8249 | 0.5850 | 0.9123 | 0.8774 | 0.4453 | 0.9128 | 0.9060 | 0.9270 | 0.9163 | |
| 512 | 0.0001 | 0.2 | 0.3461 | 0.3494 | 0.4846 | 0.7075 | 0.8317 | 0.5927 | 0.9063 | 0.8691 | 0.4217 | 0.9068 | 0.8985 | 0.9235 | 0.9107 | |
| 512 | 0.001 | 0.2 | 0.2870 | 0.2733 | 0.5726 | 0.7631 | 0.8598 | 0.6416 | 0.9064 | 0.8717 | 0.4438 | 0.9067 | 0.9048 | 0.9151 | 0.9099 | |
| 512 | 0.0005 | 0 | 0.3431 | 0.3527 | 0.4891 | 0.7096 | 0.8297 | 0.5890 | 0.9060 | 0.8709 | 0.4386 | 0.9064 | 0.9036 | 0.9160 | 0.9097 | |
| 512 | 0.0005 | 0.1 | 0.3291 | 0.3388 | 0.5100 | 0.7265 | 0.8294 | 0.5885 | 0.9060 | 0.8706 | 0.4385 | 0.9064 | 0.9029 | 0.9169 | 0.9098 | |
| 512 | 0.0005 | 0.5 | 0.3329 | 0.3264 | 0.5043 | 0.7304 | 0.8361 | 0.6211 | 0.9064 | 0.8702 | 0.4286 | 0.9068 | 0.9010 | 0.9203 | 0.9105 | |
| MRBDTA | 128 | 0.001 | 0.1 | 0.2397 | 0.2648 | 0.6430 | 0.8025 | 0.8716 | 0.6546 | 0.9340 | 0.9119 | 0.5240 | 0.9338 | 0.9427 | 0.9281 | 0.9353 |
| 64 | 0.001 | 0.1 | 0.2504 | 0.2679 | 0.6272 | 0.7935 | 0.8674 | 0.6483 | 0.9348 | 0.9113 | 0.5219 | 0.9348 | 0.9396 | 0.9336 | 0.9365 | |
| 32 | 0.001 | 0.1 | 0.2567 | 0.2721 | 0.6177 | 0.7878 | 0.8622 | 0.6400 | 0.9328 | 0.9127 | 0.5331 | 0.9324 | 0.9474 | 0.9200 | 0.9334 | |
| 128 | 0.0005 | 0.1 | 0.2350 | 0.2565 | 0.6499 | 0.8069 | 0.8775 | 0.6639 | 0.9382 | 0.9164 | 0.5304 | 0.9381 | 0.9442 | 0.9353 | 0.9397 | |
| 128 | 0.0001 | 0.1 | 0.2335 | 0.2696 | 0.6523 | 0.8084 | 0.8767 | 0.6630 | 0.9372 | 0.9171 | 0.5350 | 0.9370 | 0.9481 | 0.9284 | 0.9381 | |
| 128 | 0.001 | 0 | 0.2425 | 0.2663 | 0.6388 | 0.7997 | 0.8739 | 0.6587 | 0.9279 | 0.9010 | 0.4997 | 0.9280 | 0.9294 | 0.9316 | 0.9303 | |
| 128 | 0.001 | 0.2 | 0.2408 | 0.2620 | 0.6414 | 0.8027 | 0.8679 | 0.6478 | 0.9438 | 0.9218 | 0.5363 | 0.9438 | 0.9450 | 0.9460 | 0.9454 | |
| 128 | 0.001 | 0.5 | 0.2584 | 0.2666 | 0.6153 | 0.7894 | 0.8727 | 0.6573 | 0.9328 | 0.9123 | 0.5295 | 0.9325 | 0.9462 | 0.9215 | 0.9336 | |
| TransformerCPI | 128 | 0.001 | 0.1 | 0.3747 | 0.4111 | 0.4420 | 0.6787 | 0.8164 | 0.5668 | 0.9294 | 0.9024 | 0.5047 | 0.9296 | 0.9301 | 0.9339 | 0.9317 |
| 64 | 0.001 | 0.1 | 0.5916 | 0.5422 | 0.1191 | 0.3531 | 0.6644 | 0.3014 | 0.9318 | 0.9061 | 0.5100 | 0.9319 | 0.9337 | 0.9342 | 0.9339 | |
| 32 | 0.001 | 0.1 | 0.6219 | 0.5197 | 0.0740 | 0.2993 | 0.6360 | 0.2506 | 0.9269 | 0.8959 | 0.4860 | 0.9273 | 0.9198 | 0.9411 | 0.9302 | |
| 128 | 0.0005 | 0.1 | 0.2869 | 0.3389 | 0.5728 | 0.7750 | 0.8326 | 0.5910 | 0.9270 | 0.9008 | 0.5030 | 0.9270 | 0.9312 | 0.9270 | 0.9290 | |
| 128 | 0.0001 | 0.1 | 0.2877 | 0.3411 | 0.5716 | 0.7674 | 0.8348 | 0.5950 | 0.9208 | 0.8898 | 0.4664 | 0.9211 | 0.9181 | 0.9307 | 0.9239 | |
| 128 | 0.001 | 0 | 0.3874 | 0.3948 | 0.4232 | 0.6481 | 0.7985 | 0.5358 | 0.9297 | 0.9015 | 0.4999 | 0.9300 | 0.9272 | 0.9376 | 0.9324 | |
| 128 | 0.001 | 0.2 | 0.5181 | 0.5219 | 0.2284 | 0.5249 | 0.7476 | 0.4494 | 0.9236 | 0.8938 | 0.4780 | 0.9239 | 0.9217 | 0.9319 | 0.9265 | |
| 128 | 0.001 | 0.5 | 1.5318 | 1.1598 | -1.2809 | 0.2866 | 0.6441 | 0.2647 | 0.9218 | 0.8895 | 0.4746 | 0.9223 | 0.9153 | 0.9359 | 0.9253 | |
Appendix F Comparison of different featurization
In this section, we present the summarized featurization methods in Table 7, the detailed description of all properties is shown in Table 8. Besides, an ablation study on featurization strategies is in Table 9.
| Model Information | Atomic Properties | Hydrogen Information | Electron Properties | Stereochemistry | Structure | |||||||||||||
| Models | Graph | Atom Type | Degree | Implicit Valence | Explicit Valence | Hybridization | Aromaticity | Formal Charge | # Atom | # Hs | # Explicit Hs | # Implicit Hs | # Radical Electrons | Electron Affinity | CIP | Chirality | Ring | |
| GraphDTA | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
| GraphCPI | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
| MGraphDTA | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| SAGDTA | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
| EmbedDTI | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||
| DeepGLSTM | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
| CPI | Fingerprints | ✓ | ✓ | |||||||||||||||
| BACPI | Fingerprints | ✓ | ✓ | |||||||||||||||
| DeepNC | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
| DrugBAN | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||||
| GANDTI | Fingerprints | ✓ | ✓ | |||||||||||||||
| PGraphDTA-CNN | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||||
| BridgeDPI | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||||||||
| ColdDTA | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||
| SubMDTA | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
| IMAEN | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
| CSDTI | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| TDGraphDTA | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||||
| AMMVF | Both | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||||
| TransformerCPI | Mol. Graphs | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||||
| Name | Description |
| Atomic Properties | |
| Atom Type | Type of the atom (e.g., C, N, O, H) |
| Degree | Number of directly bonded neighbors |
| Implicit Valence | Number of implicit valence of the atom |
| Explicit Valence | Number of explicit valence of the atom |
| Hybridization | The state of hybridization (e.g., sp3, sp2) |
| Aromaticity | Whether the atom is part of an aromatic system |
| Formal Charge | The charge assigned to an atom |
| # Atom | Total number of atoms |
| Hydrogen Information | |
| # Hs | Total number of hydrogens |
| # Explicit Hs | Number of explicit hydrogens on the atom |
| # Implicit Hs | Number of implicit hydrogens on the atom |
| Electron Properties | |
| # Radical Electrons | Number of radical electrons |
| Electron Affinity | Tendency of an atom to accept electrons |
| Stereochemistry | |
| CIP | The CIP code (R or S) of the atom |
| Chirality | If an atom is a possible chiral center |
| Structure | |
| Ring | Whether the atom is part of a ring structure |
| Models | Intial Feature | Regression | Classification | |||||||||||
| MSE | MAE | R2 | PCC | CI | Spearman | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ||
| GraphDTA | Basic | 0.2771 | 0.2947 | 0.5873 | 0.7695 | 0.8521 | 0.6241 | 0.9170 | 0.8873 | 0.4830 | 0.9171 | 0.9204 | 0.9189 | 0.9194 |
| Basic+AP | 0.2772 | 0.2978 | 0.5873 | 0.7671 | 0.8496 | 0.6200 | 0.9153 | 0.8817 | 0.4517 | 0.9157 | 0.9099 | 0.9290 | 0.9191 | |
| Basic+HI | 0.2783 | 0.2983 | 0.5855 | 0.7663 | 0.8483 | 0.6185 | 0.9211 | 0.8903 | 0.4815 | 0.9214 | 0.9188 | 0.9296 | 0.9241 | |
| Basic+EP | 0.2775 | 0.3068 | 0.5868 | 0.7682 | 0.8499 | 0.6205 | 0.9165 | 0.8862 | 0.4795 | 0.9166 | 0.9185 | 0.9198 | 0.9191 | |
| Basic+Ste | 0.2838 | 0.3030 | 0.5773 | 0.7624 | 0.8523 | 0.6254 | 0.9200 | 0.8905 | 0.4869 | 0.9200 | 0.9216 | 0.9235 | 0.9224 | |
| Basic+Str | 0.2783 | 0.2991 | 0.5857 | 0.7668 | 0.8505 | 0.6228 | 0.9198 | 0.8865 | 0.4649 | 0.9202 | 0.9124 | 0.9351 | 0.9235 | |
| Basic+AP+HI | 0.2851 | 0.3029 | 0.5755 | 0.7610 | 0.8504 | 0.6222 | 0.9163 | 0.8822 | 0.4629 | 0.9168 | 0.9094 | 0.9313 | 0.9201 | |
| Basic+AP+HI+EP | 0.2845 | 0.2917 | 0.5763 | 0.7620 | 0.8510 | 0.6227 | 0.9140 | 0.8811 | 0.4580 | 0.9143 | 0.9115 | 0.9232 | 0.9173 | |
| Basic+AP+HI+EP+Ste | 0.2811 | 0.3099 | 0.5814 | 0.7640 | 0.8500 | 0.6212 | 0.9192 | 0.8899 | 0.4853 | 0.9193 | 0.9218 | 0.9215 | 0.9216 | |
| Basic+AP+HI+EP+Ste+Str | 0.2801 | 0.2916 | 0.5829 | 0.7659 | 0.8538 | 0.6278 | 0.9217 | 0.8905 | 0.4794 | 0.9220 | 0.9180 | 0.9319 | 0.9248 | |
| GraphCPI | Basic | 0.3291 | 0.3388 | 0.5100 | 0.7265 | 0.8294 | 0.5885 | 0.9060 | 0.8706 | 0.4385 | 0.9064 | 0.9029 | 0.9169 | 0.9098 |
| Basic+AP | 0.3331 | 0.3389 | 0.5040 | 0.7198 | 0.8223 | 0.5761 | 0.9038 | 0.8657 | 0.4103 | 0.9043 | 0.8955 | 0.9218 | 0.9084 | |
| Basic+HI | 0.3402 | 0.3457 | 0.4934 | 0.7157 | 0.8228 | 0.5769 | 0.9051 | 0.8713 | 0.4495 | 0.9052 | 0.9058 | 0.9094 | 0.9080 | |
| Basic+EP | 0.3408 | 0.3505 | 0.4926 | 0.7123 | 0.8211 | 0.5749 | 0.9053 | 0.8713 | 0.4442 | 0.9055 | 0.9060 | 0.9111 | 0.9085 | |
| Basic+Ste | 0.3398 | 0.3634 | 0.4940 | 0.7119 | 0.8274 | 0.5855 | 0.9061 | 0.8692 | 0.4261 | 0.9065 | 0.8992 | 0.9221 | 0.9104 | |
| Basic+Str | 0.3419 | 0.3562 | 0.4909 | 0.7113 | 0.8226 | 0.5766 | 0.9066 | 0.8683 | 0.4079 | 0.9073 | 0.8957 | 0.9281 | 0.9115 | |
| Basic+AP+HI | 0.3326 | 0.3471 | 0.5048 | 0.7212 | 0.8210 | 0.5734 | 0.9010 | 0.8659 | 0.4288 | 0.9012 | 0.9018 | 0.9071 | 0.9043 | |
| Basic+AP+HI+EP | 0.3404 | 0.3476 | 0.4931 | 0.7150 | 0.8212 | 0.5748 | 0.9015 | 0.8612 | 0.3821 | 0.9022 | 0.8890 | 0.9258 | 0.9070 | |
| Basic+AP+HI+EP+Ste | 0.3403 | 0.3445 | 0.4932 | 0.7111 | 0.8169 | 0.5671 | 0.9109 | 0.8763 | 0.4511 | 0.9113 | 0.9065 | 0.9229 | 0.9146 | |
| Basic+AP+HI+EP+Ste+Str | 0.3469 | 0.3550 | 0.4834 | 0.7073 | 0.8228 | 0.5775 | 0.9134 | 0.8772 | 0.4440 | 0.9140 | 0.9033 | 0.9328 | 0.9178 | |
Appendix G Full experiment
The complete result on the regression task is shown in Table 10, and the complete result on classification task is shown in Table 11. All experiments are run on the RTX 3090 with more than 1000 hours.
| Category | Models | DAVIS | KIBA | BindingDB | |||||||||||||||
| MSE | MAE | R2 | PCC | CI | Spearman | MSE | MAE | R2 | PCC | CI | Spearman | MSE | MAE | R2 | PCC | CI | Spearman | ||
| GNN | GraphDTA-GCN | 0.2771 | 0.2947 | 0.5873 | 0.7695 | 0.8521 | 0.6241 | 0.0005 | 0.0155 | 0.3924 | 0.6291 | 0.7502 | 0.6448 | 0.5033 | 0.4314 | 0.7259 | 0.8536 | 0.8576 | 0.7795 |
| GraphDTA-GAT | 0.2806 | 0.2928 | 0.5820 | 0.7856 | 0.8576 | 0.6331 | 0.0004 | 0.0123 | 0.6012 | 0.7720 | 0.7977 | 0.7259 | 0.5613 | 0.4460 | 0.6943 | 0.8358 | 0.8496 | 0.7662 | |
| GraphDTA-GATGCN | 0.2570 | 0.2798 | 0.6173 | 0.7867 | 0.8607 | 0.6380 | 0.0004 | 0.0126 | 0.6128 | 0.7836 | 0.8086 | 0.7357 | 0.5447 | 0.4450 | 0.7033 | 0.8415 | 0.8489 | 0.7641 | |
| GraphDTA-GIN | 0.2309 | 0.2714 | 0.6562 | 0.8105 | 0.8711 | 0.6540 | 0.0004 | 0.0131 | 0.5544 | 0.7451 | 0.7996 | 0.7400 | 0.5033 | 0.4314 | 0.7259 | 0.8537 | 0.8576 | 0.7795 | |
| GraphCPI-GCN | 0.3291 | 0.3388 | 0.5100 | 0.7265 | 0.8294 | 0.5885 | 0.0006 | 0.0156 | 0.3454 | 0.5864 | 0.7378 | 0.6154 | 0.6372 | 0.5002 | 0.6530 | 0.8133 | 0.8329 | 0.7360 | |
| GraphCPI-GAT | 0.3649 | 0.3852 | 0.4566 | 0.7609 | 0.8498 | 0.6215 | 0.0007 | 0.0179 | 0.1836 | 0.4196 | 0.6845 | 0.4896 | 0.6565 | 0.5007 | 0.6424 | 0.8163 | 0.8350 | 0.7407 | |
| GraphCPI-GATGCN | 0.3078 | 0.3193 | 0.5416 | 0.7397 | 0.8365 | 0.5995 | 0.0005 | 0.0156 | 0.4062 | 0.6421 | 0.7530 | 0.6522 | 0.6023 | 0.4858 | 0.6719 | 0.8242 | 0.8370 | 0.7442 | |
| GraphCPI-GIN | 0.2413 | 0.2848 | 0.6407 | 0.8020 | 0.8671 | 0.6478 | 0.0005 | 0.0138 | 0.4963 | 0.7093 | 0.7808 | 0.7049 | 0.5069 | 0.4247 | 0.7239 | 0.8531 | 0.8572 | 0.7784 | |
| MGraphDTA | 0.2179 | 0.2355 | 0.6755 | 0.8239 | 0.8820 | 0.6704 | 0.0003 | 0.0090 | 0.7208 | 0.8508 | 0.8649 | 0.8439 | 0.4887 | 0.3852 | 0.7338 | 0.8586 | 0.8649 | 0.7897 | |
| SAGDTA | 0.2656 | 0.2796 | 0.6045 | 0.7805 | 0.8675 | 0.6498 | 0.0039 | 0.0127 | -3.3777 | 0.5064 | 0.8096 | 0.7547 | 0.6590 | 0.4709 | 0.6410 | 0.8061 | 0.8366 | 0.7418 | |
| EmbedDTI | 0.2561 | 0.2798 | 0.6186 | 0.7874 | 0.8624 | 0.6410 | 0.0007 | 0.0175 | 0.2364 | 0.6864 | 0.6374 | 0.6897 | 0.5095 | 0.4193 | 0.7225 | 0.8516 | 0.8559 | 0.7765 | |
| DeepGLSTM | 0.2915 | 0.2941 | 0.5659 | 0.7605 | 0.8476 | 0.6176 | 0.0003 | 0.0101 | 0.6919 | 0.8325 | 0.8480 | 0.8231 | 0.5385 | 0.4246 | 0.7067 | 0.8433 | 0.8529 | 0.7713 | |
| CPI | 0.3503 | 0.3428 | 0.4784 | 0.6692 | 0.8319 | 0.5873 | 0.0003 | 0.0117 | 0.6274 | 0.7926 | 0.8190 | 0.7767 | 0.6962 | 0.5244 | 0.6208 | 0.7900 | 0.8261 | 0.7252 | |
| BACPI | 0.4036 | 0.3534 | 0.3990 | 0.6518 | 0.7982 | 0.5373 | 0.0006 | 0.0122 | 0.5320 | 0.7338 | 0.8175 | 0.7719 | 0.6468 | 0.4805 | 0.6477 | 0.8109 | 0.8297 | 0.7303 | |
| DeepNC-HGC | 0.2782 | 0.3010 | 0.5857 | 0.7669 | 0.8551 | 0.6297 | 0.0005 | 0.0141 | 0.4983 | 0.7127 | 0.7823 | 0.7052 | 0.5611 | 0.4512 | 0.6944 | 0.8367 | 0.8464 | 0.7600 | |
| DeepNC-GEN | 0.2543 | 0.2830 | 0.6213 | 0.7893 | 0.8634 | 0.6419 | 0.0004 | 0.0132 | 0.5355 | 0.7447 | 0.7981 | 0.7349 | 0.5561 | 0.4350 | 0.6971 | 0.8395 | 0.8496 | 0.7629 | |
| DrugBAN | 0.2391 | 0.2663 | 0.6440 | 0.8035 | 0.8757 | 0.6619 | 0.0004 | 0.0104 | 0.5611 | 0.7665 | 0.8388 | 0.8081 | 0.4485 | 0.3945 | 0.7557 | 0.8696 | 0.8693 | 0.7989 | |
| GANDTI | 0.3082 | 0.3305 | 0.5410 | 0.7391 | 0.8414 | 0.6079 | 0.0003 | 0.0108 | 0.6569 | 0.8146 | 0.8342 | 0.8045 | 0.6714 | 0.5221 | 0.6343 | 0.7999 | 0.8322 | 0.7342 | |
| PGraphDTA-CNN | 0.3273 | 0.3891 | 0.5126 | 0.7929 | 0.8701 | 0.6542 | 0.0005 | 0.0144 | 0.4833 | 0.6928 | 0.7473 | 0.6217 | 0.5334 | 0.4365 | 0.7095 | 0.8500 | 0.8590 | 0.7833 | |
| BridgeDPI | 0.3623 | 0.3432 | 0.6477 | 0.8062 | 0.8991 | 0.6704 | 0.0004 | 0.0133 | 0.5686 | 0.7357 | 0.7849 | 0.7239 | 0.4482 | 0.3900 | 0.7559 | 0.8702 | 0.8698 | 0.7982 | |
| ColdDTA | 0.2346 | 0.2511 | 0.6507 | 0.8083 | 0.8693 | 0.6501 | 0.0004 | 0.0129 | 0.5948 | 0.7764 | 0.8018 | 0.7451 | 0.4697 | 0.3783 | 0.7442 | 0.8645 | 0.8644 | 0.7866 | |
| SubMDTA | 0.2326 | 0.2726 | 0.6537 | 0.8091 | 0.8691 | 0.6503 | 0.0003 | 0.0098 | 0.6855 | 0.8324 | 0.8485 | 0.8243 | 0.4566 | 0.3995 | 0.7513 | 0.8677 | 0.8670 | 0.7953 | |
| IMAEN | 0.2412 | 0.2764 | 0.6409 | 0.8029 | 0.8721 | 0.6557 | 0.0004 | 0.0122 | 0.5800 | 0.7632 | 0.8061 | 0.7494 | 0.4720 | 0.3935 | 0.7429 | 0.8637 | 0.8553 | 0.7938 | |
| Transformer | CSDTI | 0.3029 | 0.3011 | 0.5490 | 0.7436 | 0.8395 | 0.6045 | 0.0007 | 0.0184 | 0.2448 | 0.4504 | 0.6475 | 0.3937 | 0.6408 | 0.4723 | 0.6510 | 0.8118 | 0.8369 | 0.7422 |
| TDGraphDTA | 0.2217 | 0.2399 | 0.6698 | 0.8201 | 0.6685 | 0.8804 | 0.0008 | 0.0292 | 0.0533 | 0.2604 | 0.3429 | 0.6243 | 0.4750 | 0.3894 | 0.7413 | 0.8631 | 0.7887 | 0.8642 | |
| AMMVF | 0.3325 | 0.3433 | 0.5048 | 0.7238 | 0.8307 | 0.5896 | 0.0006 | 0.0147 | 0.4696 | 0.6957 | 0.7711 | 0.6879 | 0.6597 | 0.4879 | 0.6407 | 0.8086 | 0.8336 | 0.7398 | |
| IIFDTI | 0.2741 | 0.3006 | 0.5918 | 0.7816 | 0.8500 | 0.6202 | 0.0005 | 0.0156 | 0.2584 | 0.7894 | 0.7952 | 0.7439 | 0.5097 | 0.4474 | 0.7170 | 0.8576 | 0.8585 | 0.7792 | |
| ICAN | 0.3481 | 0.3185 | 0.4816 | 0.7168 | 0.8211 | 0.5822 | 0.0008 | 0.0194 | 0.1159 | 0.7646 | 0.8256 | 0.7877 | 0.6582 | 0.5028 | 0.6415 | 0.8145 | 0.8277 | 0.7271 | |
| MolTrans | 0.2588 | 0.2740 | 0.6146 | 0.7906 | 0.8601 | 0.5927 | 0.0003 | 0.0114 | 0.6378 | 0.8064 | 0.8453 | 0.7865 | 0.5138 | 0.4101 | 0.7201 | 0.8517 | 0.8570 | 0.7786 | |
| TransformerCPI | 0.2869 | 0.3389 | 0.5728 | 0.7750 | 0.8326 | 0.5910 | 0.0008 | 0.0233 | 0.0728 | 0.8051 | 0.8357 | 0.7809 | 0.5704 | 0.4624 | 0.6894 | 0.8394 | 0.8426 | 0.7507 | |
| MRBDTA | 0.2350 | 0.2565 | 0.6499 | 0.8069 | 0.8775 | 0.6639 | 0.0006 | 0.0171 | 0.3210 | 0.5718 | 0.7239 | 0.5878 | 0.4977 | 0.4133 | 0.7289 | 0.8557 | 0.8629 | 0.7874 | |
| FOTFCPI | 0.2803 | 0.3004 | 0.5825 | 0.7704 | 0.8546 | 0.6286 | 0.0004 | 0.0133 | 0.5342 | 0.7552 | 0.7947 | 0.7353 | 0.5743 | 0.4472 | 0.6872 | 0.8322 | 0.8444 | 0.7569 | |
| our | 0.2063 | 0.2481 | 0.6927 | 0.8330 | 0.8901 | 0.6839 | 0.0003 | 0.0094 | 0.7168 | 0.8512 | 0.8677 | 0.8432 | 0.4651 | 0.3878 | 0.7467 | 0.8657 | 0.8683 | 0.7956 | |
| Categories | Models | Human | C.elegans | Drugbank | ||||||||||||||||||
| ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ||
| GNN | GraphDTA-GCN | 0.9222 | 0.8922 | 0.4852 | 0.9224 | 0.9210 | 0.9293 | 0.9251 | 0.9488 | 0.9174 | 0.5342 | 0.9468 | 0.9368 | 0.9564 | 0.9465 | 0.7590 | 0.6925 | 0.1490 | 0.7589 | 0.7458 | 0.7842 | 0.7645 |
| GraphDTA-GAT | 0.8935 | 0.8557 | 0.3986 | 0.8937 | 0.8925 | 0.9024 | 0.8974 | 0.9289 | 0.8909 | 0.4942 | 0.9287 | 0.9154 | 0.9423 | 0.9286 | 0.7684 | 0.7030 | 0.1595 | 0.7683 | 0.7587 | 0.7855 | 0.7718 | |
| GraphDTA-GATGCN | 0.9296 | 0.9024 | 0.5030 | 0.9297 | 0.9298 | 0.9342 | 0.9319 | 0.9487 | 0.9176 | 0.5318 | 0.9484 | 0.9335 | 0.9640 | 0.9485 | 0.7712 | 0.7046 | 0.1585 | 0.7711 | 0.7573 | 0.7968 | 0.7764 | |
| GraphDTA-GIN | 0.9019 | 0.8674 | 0.4393 | 0.9021 | 0.9037 | 0.9065 | 0.9051 | 0.9470 | 0.9174 | 0.5342 | 0.9468 | 0.9368 | 0.9564 | 0.9465 | 0.7871 | 0.7232 | 0.1797 | 0.7871 | 0.7797 | 0.7988 | 0.7892 | |
| GraphCPI-GCN | 0.9034 | 0.8649 | 0.4048 | 0.9040 | 0.8941 | 0.9229 | 0.9083 | 0.9322 | 0.8940 | 0.4967 | 0.9319 | 0.9155 | 0.9493 | 0.9320 | 0.7362 | 0.6688 | 0.1292 | 0.7361 | 0.7178 | 0.7764 | 0.7459 | |
| GraphCPI-GAT | 0.8935 | 0.8557 | 0.3986 | 0.8937 | 0.8925 | 0.9024 | 0.8974 | 0.9281 | 0.8913 | 0.4965 | 0.9279 | 0.9181 | 0.9373 | 0.9275 | 0.7515 | 0.6862 | 0.1470 | 0.7515 | 0.7433 | 0.7667 | 0.7548 | |
| GraphCPI-GATGCN | 0.9097 | 0.8763 | 0.4644 | 0.9099 | 0.9090 | 0.9169 | 0.9129 | 0.9372 | 0.9028 | 0.5109 | 0.9370 | 0.9245 | 0.9496 | 0.9369 | 0.7561 | 0.6909 | 0.1508 | 0.7561 | 0.7479 | 0.7711 | 0.7593 | |
| GraphCPI-GIN | 0.8870 | 0.8484 | 0.3791 | 0.8872 | 0.8881 | 0.8935 | 0.8908 | 0.9407 | 0.9072 | 0.5173 | 0.9405 | 0.9270 | 0.9543 | 0.9404 | 0.7866 | 0.7241 | 0.1845 | 0.7866 | 0.7840 | 0.7903 | 0.7870 | |
| MGraphDTA | 0.9408 | 0.9166 | 0.5244 | 0.9410 | 0.9393 | 0.9466 | 0.9429 | 0.9631 | 0.9407 | 0.5795 | 0.9630 | 0.9535 | 0.9723 | 0.9628 | 0.8146 | 0.7539 | 0.2154 | 0.8146 | 0.8090 | 0.8228 | 0.8157 | |
| SAGDTA | 0.9021 | 0.8728 | 0.4631 | 0.9018 | 0.9172 | 0.8897 | 0.9032 | 0.9380 | 0.9050 | 0.5165 | 0.9378 | 0.9281 | 0.9472 | 0.9375 | 0.7655 | 0.7018 | 0.1631 | 0.7655 | 0.7624 | 0.7700 | 0.7662 | |
| EmbedDTI | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.7625 | 0.6963 | 0.1526 | 0.7625 | 0.7504 | 0.7852 | 0.7674 | |
| DeepGLSTM | 0.9180 | 0.8857 | 0.4758 | 0.9183 | 0.9145 | 0.9281 | 0.9212 | 0.9414 | 0.9103 | 0.5257 | 0.9412 | 0.9325 | 0.9496 | 0.9409 | 0.7640 | 0.6988 | 0.1567 | 0.7640 | 0.7551 | 0.7801 | 0.7674 | |
| CPI | 0.9110 | 0.8789 | 0.4638 | 0.9111 | 0.9126 | 0.9152 | 0.9138 | 0.9312 | 0.8940 | 0.4983 | 0.9310 | 0.9176 | 0.9446 | 0.9309 | 0.7467 | 0.6812 | 0.1434 | 0.7466 | 0.7368 | 0.7683 | 0.7515 | |
| BACPI | 0.9249 | 0.8958 | 0.4924 | 0.9251 | 0.9239 | 0.9313 | 0.9276 | 0.9556 | 0.9314 | 0.5637 | 0.9555 | 0.9493 | 0.9610 | 0.9551 | 0.7748 | 0.7100 | 0.1664 | 0.7748 | 0.7663 | 0.7895 | 0.7776 | |
| DeepNC-HGC | 0.8796 | 0.8399 | 0.3579 | 0.8798 | 0.8823 | 0.8845 | 0.8834 | 0.9418 | 0.9106 | 0.5251 | 0.9417 | 0.9325 | 0.9504 | 0.9413 | 0.7673 | 0.7013 | 0.1571 | 0.7673 | 0.7556 | 0.7889 | 0.7718 | |
| DeepNC-GEN | 0.9178 | 0.8873 | 0.4809 | 0.9180 | 0.9186 | 0.9224 | 0.9205 | 0.9501 | 0.9203 | 0.5385 | 0.9499 | 0.9366 | 0.9637 | 0.9498 | 0.7464 | 0.6785 | 0.1357 | 0.7463 | 0.7262 | 0.7928 | 0.7571 | |
| DrugBAN | 0.9302 | 0.9048 | 0.5083 | 0.9302 | 0.9342 | 0.9299 | 0.9320 | 0.9596 | 0.9346 | 0.5645 | 0.9594 | 0.9478 | 0.9710 | 0.9593 | 0.8188 | 0.7610 | 0.2325 | 0.8188 | 0.8203 | 0.8157 | 0.8179 | |
| GANDTI | 0.9333 | 0.9090 | 0.5162 | 0.9333 | 0.9374 | 0.9327 | 0.9351 | 0.9404 | 0.9082 | 0.5213 | 0.9403 | 0.9300 | 0.9504 | 0.9400 | 0.7427 | 0.6737 | 0.1298 | 0.7425 | 0.7188 | 0.7954 | 0.7550 | |
| PGraphDTA-CNN | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.7910 | 0.7260 | 0.1787 | 0.7910 | 0.7787 | 0.8117 | 0.7948 | |
| BridgeDPI | 0.9456 | 0.9250 | 0.5435 | 0.9456 | 0.9485 | 0.9457 | 0.9471 | 0.9651 | 0.9470 | 0.6054 | 0.9651 | 0.9629 | 0.9663 | 0.9646 | 0.7825 | 0.7181 | 0.1742 | 0.7825 | 0.7739 | 0.7974 | 0.7853 | |
| ColdDTA | 0.9420 | 0.9203 | 0.5348 | 0.9421 | 0.9453 | 0.9420 | 0.9436 | 0.9623 | 0.9403 | 0.5785 | 0.9622 | 0.9543 | 0.9697 | 0.9619 | 0.8186 | 0.7578 | 0.2175 | 0.8186 | 0.8108 | 0.8300 | 0.8203 | |
| SubMDTA | 0.9326 | 0.9054 | 0.5066 | 0.9328 | 0.9304 | 0.9400 | 0.9351 | 0.9570 | 0.9313 | 0.5589 | 0.9568 | 0.9461 | 0.9674 | 0.9566 | 0.8045 | 0.7414 | 0.1968 | 0.8045 | 0.7946 | 0.8204 | 0.8072 | |
| IMAEN | 0.9058 | 0.8719 | 0.4489 | 0.9059 | 0.9067 | 0.9111 | 0.9089 | 0.9506 | 0.9255 | 0.5546 | 0.9505 | 0.9469 | 0.9530 | 0.9500 | 0.7995 | 0.7365 | 0.1931 | 0.7994 | 0.7918 | 0.8113 | 0.8014 | |
| Transformer | CSDTI | 0.8630 | 0.8216 | 0.3223 | 0.8630 | 0.8707 | 0.8620 | 0.8663 | 0.8962 | 0.8446 | 0.3710 | 0.8958 | 0.8759 | 0.9183 | 0.8966 | 0.7269 | 0.6626 | 0.1304 | 0.7269 | 0.7196 | 0.7420 | 0.7306 |
| TDGraphDTA | 0.9411 | 0.9226 | 0.5461 | 0.9409 | 0.9525 | 0.9316 | 0.9419 | 0.9573 | 0.9315 | 0.5591 | 0.9571 | 0.9460 | 0.9681 | 0.9569 | 0.8083 | 0.7448 | 0.1982 | 0.8082 | 0.7959 | 0.8281 | 0.8116 | |
| AMMVF | 0.9287 | 0.9004 | 0.4994 | 0.9290 | 0.9268 | 0.9362 | 0.9314 | 0.9636 | 0.9426 | 0.5892 | 0.9635 | 0.9567 | 0.9700 | 0.9632 | 0.7814 | 0.7142 | 0.1717 | 0.7814 | 0.7728 | 0.7973 | 0.7849 | |
| IIFDTI | 0.9392 | 0.9163 | 0.5279 | 0.9392 | 0.9420 | 0.9400 | 0.9409 | 0.9679 | 0.9500 | 0.6054 | 0.9679 | 0.9633 | 0.9718 | 0.9675 | 0.8084 | 0.7441 | 0.1950 | 0.8084 | 0.7934 | 0.8327 | 0.8125 | |
| ICAN | 0.9387 | 0.9134 | 0.5194 | 0.9389 | 0.9361 | 0.9460 | 0.9410 | 0.9644 | 0.9479 | 0.6146 | 0.9644 | 0.9662 | 0.9613 | 0.9637 | 0.7731 | 0.7092 | 0.1682 | 0.7731 | 0.7681 | 0.7812 | 0.7745 | |
| MolTrans | 0.9453 | 0.9233 | 0.5377 | 0.9455 | 0.9451 | 0.9492 | 0.9471 | 0.9639 | 0.9428 | 0.5847 | 0.9638 | 0.9565 | 0.9705 | 0.9634 | 0.7910 | 0.7283 | 0.1879 | 0.7909 | 0.7861 | 0.7988 | 0.7922 | |
| TransformerCPI | 0.9311 | 0.9045 | 0.5060 | 0.9312 | 0.9316 | 0.9351 | 0.9333 | 0.9581 | 0.9343 | 0.5696 | 0.9580 | 0.9501 | 0.9655 | 0.9577 | 0.8097 | 0.7470 | 0.2024 | 0.8097 | 0.7992 | 0.8265 | 0.8125 | |
| MRBDTA | 0.9447 | 0.9258 | 0.5521 | 0.9446 | 0.9523 | 0.9397 | 0.9458 | 0.9713 | 0.9569 | 0.6335 | 0.9713 | 0.9708 | 0.9710 | 0.9709 | 0.8102 | 0.7469 | 0.2003 | 0.8102 | 0.7978 | 0.8298 | 0.8134 | |
| FOTFCPI | 0.9444 | 0.9235 | 0.5413 | 0.9444 | 0.9477 | 0.9443 | 0.9459 | 0.9673 | 0.9486 | 0.6009 | 0.9672 | 0.9618 | 0.9721 | 0.9669 | 0.7845 | 0.7188 | 0.1713 | 0.7845 | 0.7713 | 0.8075 | 0.7889 | |
| our | 0.9435 | 0.9221 | 0.5400 | 0.9435 | 0.9462 | 0.9443 | 0.9451 | 0.9688 | 0.9518 | 0.6178 | 0.9688 | 0.9652 | 0.9718 | 0.9684 | 0.8035 | 0.7405 | 0.1961 | 0.8035 | 0.7941 | 0.8185 | 0.8061 | |
Appendix H Memory and Parameter Comparison
| Categories | Models | Regression | Classification | ||||
| Model parameter | Memory Usage (MB) | Time(s) | Model parameter | Memory Usage (MB) | Run Time (s) | ||
| Graph | GraphDTA-GCN | 7.87 | 86.45 | 8.92 | 7.87 | 86.33 | 2.43 |
| GraphDTA-GAT | 6.58 | 104.71 | 9.62 | 6.58 | 99.40 | 2.43 | |
| GraphDTA-GATGCN | 18.12 | 148.25 | 8.37 | 18.12 | 145.13 | 2.35 | |
| GraphDTA-GIN | 5.97 | 78.00 | 12.33 | 5.95 | 77.47 | 3.13 | |
| GraphCPI-GCN | 10.46 | 98.13 | 7.02 | 10.48 | 63.37 | 1.92 | |
| GraphCPI-GAT | 9.16 | 116.19 | 9.38 | 9.18 | 112.34 | 2.48 | |
| GraphCPI-GATGCN | 20.70 | 158.21 | 9.47 | 20.73 | 156.22 | 2.20 | |
| GraphCPI-GIN | 8.55 | 88.55 | 12.54 | 8.56 | 88.02 | 2.92 | |
| MGraphDTA | 11.75 | 235.97 | 69.84 | 11.43 | 217.15 | 17.59 | |
| SAGDTA | 7.45 | 88.31 | 20.87 | 7.44 | 87.54 | 4.34 | |
| EmbedDTI | 16.97 | 152.55 | 17.80 | 16.97 | - | - | |
| DeepGLSTM | 131.92 | 1287.92 | 20.69 | 131.93 | 1287.16 | 11.22 | |
| CPI | 0.37 | 14.00 | 11.29 | 0.6 | 14.82 | 2.69 | |
| BACPI | 4.05 | 1051.91 | 43.27 | 6.13 | 1058.95 | 12.38 | |
| DeepNC-HGC | 16.61 | 123.70 | 9.85 | 16.60 | 123.65 | 3.46 | |
| DeepNC-GEN | 18.84 | 174.00 | 11.35 | 18.84 | 166.55 | 3.46 | |
| DrugBAN | 4.10 | 940.22 | 30.06 | 4.10 | 940.23 | 7.84 | |
| GANDTI | 1.48 | 35.89 | 6.01 | 2.43 | 39.95 | 1.54 | |
| PGraphDTA-CNN | 9.03 | 102.85 | 13.71 | 9.03 | - | - | |
| BridgeDPI | 39.32 | 232.53 | 16.27 | 39.32 | 232.53 | 4.36 | |
| ColdDTA | 13.14 | 282.74 | 72.98 | 13.14 | 262.91 | 18.56 | |
| SubMDTA | 169.37 | 992.61 | 35.12 | 195.50 | 1095.73 | 8.49 | |
| IMAEN | 10.43 | 174.34 | 35.77 | 10.43 | 172.86 | 4.41 | |
| Transformer | CSDTI | 9.67 | 281.23 | 17.66 | 9.66 | 281.02 | 4.35 |
| TDGraphDTA | 8.62 | 247.23 | 116.38 | 8.62 | 236.02 | 28.43 | |
| AMMVF | 6.68 | 17847.62 | 216.20 | 7.49 | 17850.79 | 57.99 | |
| IIFDTI | 10.75 | 7946.92 | 141.12 | 10.75 | 11890.79 | 56.95 | |
| ICAN | 63.89 | 649.55 | 12.44 | 63.89 | 648.67 | 2.84 | |
| MolTrans | 239.73 | 10624.55 | 70.19 | 239.74 | 10624.55 | 25.06 | |
| TransformerCPI | 4.44 | 1219.58 | 28.98 | 4.45 | 1219.60 | 7.17 | |
| MRBDTA | 17.83 | 3893.76 | 66.47 | 17.84 | 3893.78 | 16.13 | |
| FOTFCPI | 189.15 | 6780.35 | 58.75 | 189.15 | 6780.35 | 14.80 | |
| Our | 19.02 | 1081.99 | 94.71 | 19.02 | 1082.68 | 13.68 | |
Appendix I Limitations
In this section, we provide limitations on the proposed benchmark. A limitation of the proposed benchmark is that it considers a limited number of datasets and solely on random splitting for dataset splitting. Additionally, the benchmark only evaluates models using simple inputs without incorporating additional information, which may not reflect the current trend where many models depend on pre-trained language models.

### v2

@@
-Benchmark on Drug Target Interaction Modeling from a Structure Perspective
+Benchmark on Drug Target Interaction Modeling from a Drug Structure Perspective
 Abstract
@@
 However, the benchmarking of these novel methods often varies significantly in terms of hyperparameter settings and datasets, which limits algorithmic progress.
-In view of these, we conduct a comprehensive survey and benchmark for drug-target interaction modeling from a structure perspective, via integrating tens of explicit (i.e., GNN-based) and implicit (i.e., Transformer-based) structure learning algorithms.
-To this end, we first unify the hyperparameter setting within each class of structure learning methods.
-Moreover, we conduct a macroscopical comparison between these two classes of encoding strategies as well as the different featurization techniques that inform molecules’ chemical and physical properties.
-We then carry out the microscopical comparison between all the integrated models across the six datasets, via comprehensively benchmarking their effectiveness and efficiency.
+In view of these, we conducted a comprehensive survey and benchmark for drug-target interaction modeling from a structural perspective via integrating tens of explicit (i.e., GNN-based) and implicit (i.e., Transformer-based) structure learning algorithms.
+We conducted a macroscopical comparison between these two classes of encoding strategies as well as the different featurization techniques that inform molecules’ chemical and physical properties.
+We then carry out the microscopical comparison between all the integrated models across the six datasets via comprehensively benchmarking their effectiveness and efficiency.
+To ensure fairness, we investigate model performance under individually optimized configuration.
 Remarkably, the summarized insights from the benchmark studies lead to the design of model combos.
 We demonstrate that our combos can achieve new state-of-the-art performance on various datasets associated with cost-effective memory and computation.
-Our code is available at https://github.com/justinwjl/GTB-DTI/tree/main.
-1 Introduction
+11footnotetext: These authors contributed equally to this work.1 Introduction
 The prediction modeling of drug-target interactions (DTI) has emerged as an irreplaceable task for efficacious therapeutic interventions.
-The binding affinity between a drug molecule and its target protein plays a significant role in the design and repurpose of drugs, where a high affinity typically indicates the desired therapeutics, target specificity, long residence, and drug resistance delay [1, 2, 3].
-The precise modeling of DTI can expedite the drug discovery process and circumvent the associated cost [4, 5].
-Deep learning based frameworks have recently revolutionized this field, enabling more accurate and efficient predictions compared with laboratory experimental methods [6, 7, 8].
-Within the deep learning frameworks [9, 10], drugs are commonly represented using the Simplified Molecular Input Line Entry System (SMILES) [11], and proteins are represented as sequences of amino acids.
-These representations are processed by separate convolutional neural networks (CNNs) [12, 13] and subsequently integrated and processed using a multi-layer perceptron (MLP) for DTI prediction.
+The binding affinity between a drug molecule and its target protein plays a significant role in the design and repurposing of drugs, where a high affinity typically indicates the desired therapeutics, target specificity, long residence, and drug resistance delay (hughes2011principles; copeland2006drug; swinney2004biochemical).
+The precise modeling of DTI can expedite the drug discovery process and circumvent the associated cost (Ashburn2004; Strittmatter2014).
+Deep learning-based frameworks have recently revolutionized this field, enabling more accurate predictions and accelerating the discovery of new compounds by guiding laboratory experiments more efficiently (wen2017deep; abbasi2021deep; huang2020deeppurpose).
+Within deep learning frameworks (ozturk2018deepdta; ozturk2019widedta), drugs are commonly represented using the Simplified Molecular Input Line Entry System (SMILES)(Smiles), and proteins are represented as sequences of amino acids.
+These representations are typically processed using various neural network architectures, such as convolutional neural networks (CNNs) (krizhevsky2017imagenet; he2016deep), recurrent neural networks (RNNs), Transformers, and so on, before being integrated and processed by a multi-layer perceptron (MLP) for DTI prediction.
 It is notorious that the reliance on sequence-based representations can result in the loss of structural information, which can potentially compromise the DTI predictive capability.
-From the drug perspective, molecular structure modeling helps identify the specific binding sites [14], contribute to predicting pharmacokinetic properties [15], and allow conformational flexibility [16].
+From the drug perspective, molecular structure modeling helps identify the specific binding sites (ma2011molecular), contributes to predicting pharmacokinetic properties (ekins2007silico), and allows conformational flexibility (karplus2005molecular).
 To address this problem, a number of drug algorithms have been proposed to promote DTI prediction, which can be categorized into explicit and implicit structure learning.
-First, graph neural networks (GNNs) [17, 18] have been widely adopted to learn the molecular structures, owing to their ability to directly operate on graph-based representations of molecules.
+First, graph neural networks (GNNs) (kipf2016semi; nguyen2021graphdta) have been widely adopted to learn the molecular structures, owing to their ability to directly operate on graph-based representations of molecules.
 By explicitly propagating information through the graph, GNNs can learn node and edge features and thereby capture the structural and functional relationships between atoms and bonds.
-Second, Transformers, originally focused on natural language processing [19], have also shown promise in biomedical applications [20, 21].
+Second, Transformers, originally focused on natural language processing (Transformer), have also shown promise in biomedical applications (huang2021moltrans; chen2020Transformercpi).
 They rely on self-attention mechanisms to implicitly weight the correlations between different parts of the input SMILES, allowing them to capture long-range dependencies and contextual information.
 While these techniques contribute to the learning of drug structures, there is still a key knob under-explored: we lack a systematic study to benchmark their effectiveness and efficiency.
-Without such a standardized benchmark, it is unachievable to offer fair comparisons and subsequently summarize designing philosophy necessary to inform DTI.
-There have been several surveys and benchmarks on computational methods for DTI prediction [9, 8, 22, 23], which leave out the recent developments of structure learning algorithms and unavoidably fail to focus on drug structure benchmarking.
-Moreover, although massive efforts [24, 25, 18] have been made to explore the effectiveness of modeling structural information, they predominantly use their proprietary training hyperparameters, datasets, and evaluation metrics.
-Due to the various settings, one cannot reach convincing answers whether a configuration of structure encoders and/or featurization methods generally performs well.
+Without such a standardized benchmark, it is unachievable to offer fair comparisons and subsequently summarize the design philosophy necessary to inform DTI.
+There have been several surveys and benchmarks on computational methods for DTI prediction(ozturk2018deepdta; huang2020deeppurpose; huang2021therapeutics; xu2022peer), which leave out the recent developments of structure learning algorithms and unavoidably fail to focus on drug structure benchmarking.
+Moreover, although massive efforts (bal2024pgraphdta; zhu2023tdgraph; nguyen2021graphdta) have been made to explore the effectiveness of modeling structural information, they predominantly use their proprietary training hyperparameters, datasets, and evaluation metrics.
+Due to the various settings, one cannot reach convincing answers as to whether a configuration of structure encoders and/or featurization methods generally performs well.
 The complex of DTI classification and regression tasks and datasets complicates the benchmark comparison.
 In this study, we introduce GTB-DTI, a comprehensive benchmark customized for GNN and Transformer-based methodologies for DTI prediction.
-i) We thoroughly examine the implementation details for each category of drug structure learning methods and integrate three widely-used datasets for classification and regression tasks, respectively.
-Then, we harmonize the sensitive hyperparameters across different methods using a greedy search to identify an optimal sweet spot configuration.
-The unified setting lays the foundation for a fair and reproducible benchmark.
+I) We thoroughly examine the implementation details for each category of drug structure learning methods and integrate three widely used datasets for classification and regression tasks, respectively.
+We use the optimal hyperparameters reported in their papers to lay the foundation for a fair and reproducible benchmark.
 II) To gain macroscopical insights into the structure encoders and featurization methods, we fix the drug encoder to be either GNN or Transformer-based approaches and benchmark these two strategies in the various settings.
 We also integrate tens of drug features given their importance to inform molecules’ chemistry and physical properties and evaluate them on the representative datasets.
-iii) To gain macroscopical insights into nuance between 31 concerned models, we conduct the benchmark studies of their effectiveness on the six datasets with the unified setting.
+III) To gain macroscopical insights into nuance between 31 concerned models, we conduct the benchmark studies of their effectiveness on the six datasets.
 Moreover, we assess the efficiency of each method by measuring peak GPU memory usage, running time, and convergences.
-iv) The comprehensive study finally provides a number of surprising observations: ❶ The CNN-encoder accompanied with integer features has the close protein embedding performance compared to the Transformer or larger language models, but they are more efficient.
-❷ The explicit and implicit structure encoders for drugs exhibit unequal performances across the different datasets, which suggests their hybrid usage for generalization purpose.
-❸ Inspired from these insights, we conclude with a model combos that leads us to attaining state-of-the-art (SOTA) regression results and performing similarly to SOTA in the DTI classifications.
-Our combos further deliver cost-effective memory usage and running time as well as faster convergence, which can serve as new baseline for the following explorations.
+IV) The comprehensive study finally provides a number of surprising observations: i) The CNN encoder accompanied by integer features has close protein embedding performance compared to the Transformer or larger language models, but they are more efficient.
+ii) The explicit and implicit structure encoders for drugs exhibit unequal performances across the different datasets, which suggests their hybrid usage for generalization purposes.
+iii) Inspired by these insights, we conclude with a model combo that leads us to attaining state-of-the-art (SOTA) regression results and performing similarly to SOTA in the DTI classifications.
+Our combos further deliver cost-effective memory usage and running time as well as faster convergence, which can serve as a new baseline for the following explorations.
 2 Formulations for Drug-target Interaction Modeling
-In this research, we focus on the formulations of recently-emerging structure modeling approaches for drug molecules, which could be categorized into explicit methods based on graph neural networks and implicit methods based on transformer.
+In this research, we focus on the formulations of recently emerging structure modeling approaches for drug molecules, which could be categorized into explicit methods based on graph neural networks and implicit methods based on Transformers.
 The target proteins are learned by the sophisticated tools of convolutional/recurrent neural networks (CNNs/RNNs) or Transformers, after which both the molecules’ and proteins’ embeddings are integrated to facilitate interaction prediction.
-We will also summarize and benchmark the various widely-adopted molecule features.
+We will also summarize and benchmark the various widely adopted molecule features.
 2.1 Graph Neural Networks based Methods
-A drug molecule is typically represented as a graph , where and denotes the sets of atoms and chemical bonds, respectively.
-The classical GNN frameworks involve key processes of aggregating and updating node features, collectively referred to as message passing, which can be mathematically represented as [26, 27]:
+A drug molecule is typically represented as a graph , where and denote the sets of atoms and chemical bonds, respectively.
+The classical GNN frameworks involve key processes of aggregating and updating node features, collectively referred to as message passing, which can be mathematically represented as follows(scarselli2008graph; duan2022comprehensive):
 where is the feature representation of node at layer , is the feature representation of edge between nodes and , refers to the set of neighboring nodes next to node .
-Functions and aim to aggregate the neighborhood representations and integrate them together with the nodes features, respectively.
+Functions , aim to aggregate the neighborhood representations and integrate them together with the node features, respectively.
 Additionally, and are feature mapping functions, parameterized by and , respectively.
-The molecule’s representation can be derived using READOUT function, which processes on the set of vertex features at the last layer.
+The molecule’s representation can be derived using READOUT a function that processes the set of vertex features at the last layer.
 Graph Convolutional Networks (GCN).
@@
 Considering the self-connection of atoms, we have .
-Let denote the initial atom feature matrix.
-GCN [28] models the message passing as follows:
+Let’s denote the initial atom feature matrix.
+GCN (gcn) models the message passing as follows:
 where is the node feature matrix at layer , starting with .
 Matrix represents the learnable weights for layer , denotes a non-linear activation function, e.g., ReLU, and is a diagonal degree matrix of .
-A couple of pioneering works have leveraged GCN to facilitate drug-protein interaction prediction [29, 30, 31, 32].
-For example, DeepGLSTM [29] uses mixture-of-depths GCNs to capture drug representations from different scales.
-CPI [31] considers cross-atom distance and introduces the concept of r-radius subgraphs [33], using r-radius vertices and edges to redefine the structure of graphs.
+A couple of pioneering works have leveraged GCN to facilitate drug-protein interaction prediction (mukherjee2022deepglstm; tran2022deepnc; tsubaki2019compound; pan2023csdti).
+For example, DeepGLSTM (mukherjee2022deepglstm) uses mixture-of-depths GCNs to capture drug representations from different scales.
+CPI (tsubaki2019compound) considers cross-atom distance and introduces the concept of r-radius subgraphs (r-radius), using r-radius vertices and edges to redefine the structure of graphs.
 Graph Isomorphism Networks (GIN).
-GIN excels in learning distinct graph features by approximating the Weisfeiler-Lehman test, enabling it to distinguish a wide range of graph structures [34].
-The message passing process at the -th layer is of the following form:
+GIN excels in learning distinct graph features by approximating the Weisfeiler-Lehman test, enabling it to distinguish a wide range of graph structures (xu2018powerful).
+The message-passing process at the -th layer is of the following form:
 where is a multi-layer perceptron that parameterizes the update function, and is a learnable parameter.
 We benchmark several GIN-based drug-target interaction modeling methods.
-GraphCPI [35] and GraphDTA [18] adopt GIN-based models with batch normalization to obtain the drug representation.
-SubMDTA [36] uses subgraph’s generation task and contrastive learning to pretrain a molecular graph encoder with multiple GIN layers for further prediction.
-Graph Attention Networks (GAT). Unlike fixed-weight aggregation, GAT [37] employs an attention mechanism to determine neighborhood importance and learn the node embeddings as:
+GraphCPI (quan2019graphcpi) and GraphDTA (nguyen2021graphdta) adopt GIN-based models with batch normalization to obtain the drug representation.
+SubMDTA (pan2023submdta) uses a subgraph’s generation task and contrastive learning to pretrain a molecular graph encoder with multiple GIN layers for further prediction.
+Graph Attention Networks (GAT). Unlike fixed-weight aggregation, GAT (gat) employs an attention mechanism to determine neighborhood importance and learn the node embeddings as follows:
+denotes attention weights, and is a concatenating operation.
+GraphDTA (nguyen2021graphdta) and AMMVF (wang2023ammvf) leverage the multi-head GAT layers to optimize the atom messaging.
+They integrate GAT with other architectural modules, such as GCN, facilitating a more comprehensive representation of drugs.
+Graph Transformers.
+Graph Transformers (graphTransformer1; maziarka2020molecule) have emerged as powerful alternatives to traditional graph neural networks (GNNs) for molecular representation learning.
+Unlike conventional GNNs, which rely on message-passing mechanisms to propagate local node information, Graph Transformers leverage self-attention mechanisms to capture both local and global dependencies more effectively.
+By integrating Message Passing Networks into Transformer-style architectures, these models enhance expressiveness, enabling more comprehensive encoding of molecular structures.
+This hybrid approach allows Graph Transformers to preserve structural information while benefiting from the flexibility of attention-based learning.
 2.2 Transformer-based Methods
-Besides the graph representation, drugs could also be decorated as SMILES strings [39] and encoded similarly to natural language processing.
-Specifically, after tokenizing SMILES strings, Transformer model utilizes multi-head attention to model the interactions between different segments of the input and obtain the molecular representations.
+Besides the graph representation, drugs could also be decorated as SMILES strings (weininger1988smiles) and encoded similarly to natural language processing.
+Specifically, after tokenizing SMILES strings, the Transformer model utilizes multi-head attention to model the interactions between different segments of the input and obtain the molecular representations.
 Positional encodings are also integrated to preserve the sequence order, enhancing the model’s ability to process sequential information effectively.
 We review and benchmark two typical types of attention mechanisms used for molecular representations.
-Self-Attention [20, 40, 41].
+Self-Attention(huang2021moltrans; qian2023mcl; yin2024fotf).
 Self-attention computes a weighted sum of all input values based on their relevance to each other.
-Considering an embedding of SMILES sequence at a specific transformer layer, where and are token length and dimension, respectively, the attention is calculated by .
+Considering an embedding of a SMILES sequence at a specific Transformer layer, where and are token length and dimension, respectively, the attention is calculated by .
 and are projections of the input matrix .
 Multi-head attention combines these projections across different subspaces for a more detailed analysis.
-Following by normalization and feed-forward neural networks, the SMILES embedding is updated as and the output from the last layer is treated as molecular representations.
-Transformer encoders like MolTrans [20] and FOTFCPI [41] are adopted to enhance sub-structure embeddings in proteins and drugs.
-Cross Attention [42, 40].
+Followed by normalization and feed-forward neural networks, the SMILES embedding is updated, and the output from the last layer is treated as molecular representations.
+Transformer encoders like MolTrans (huang2021moltrans) and FOTFCPI (yin2024fotf) are adopted to enhance substructure embeddings in proteins and drugs.
+Cross-Attention(kurata2022ican; qian2023mcl).
 Cross-attention is designed to capture the interaction between the drug and protein sequences, with the query matrix derived from one sequence and the key and value matrices from another.
-This mechanism is particularly useful in integrating hybrid representations such as drug graphs and SMILES [38], as well as drugs and proteins [32, 42].
+This mechanism is particularly useful in integrating hybrid representations such as drug graphs and SMILES (wang2023ammvf), as well as drugs and proteins (pan2023csdti; kurata2022ican).
 2.3 Feature Processing Methods
@@
 Both drugs and proteins are input as strings of ASCII characters, whose features can be extracted using statistical solutions.
-Integer encoding [18] simply converts the string to a sequence of integers, which assigns an integer to each character.
-The N-gram [43] captures the statistical dependencies between characters in an input string.
+Integer encoding (nguyen2021graphdta) simply converts the string to a sequence of integers, which assigns an integer to each character.
+The N-gram (10.1093/bioinformatics/bti801) captures the statistical dependencies between characters in an input string.
 Specifically, a 3-gram model breaks down a sequence into , analyzing the relationship between adjacent characters.
@@
 The additional chemical properties and structural details of SMILES strings are often considered to gain a more comprehensive understanding.
-Extended-Connectivity Fingerprints (ECFP) [44, 45], involves generating unique identifiers for atoms based on their local chemical environment and iteratively updating these through a hash function to capture a broader molecular context, ultimately producing a set of fingerprints that represent the molecule’s overall structure.
-Another approach, RDKit, is used to convert SMILES into molecular graphs [46, 18], where nodes represent the physical and chemical properties of molecules, and bonds are represented by an adjacency matrix.
+Extended-Connectivity Fingerprints (ECFP) (doi:10.1021/c160017a018Morgan; doi:10.1021/ci100050tecfp), involves generating unique identifiers for atoms based on their local chemical environment and iteratively updating these through a hash function to capture a broader molecular context, ultimately producing a set of fingerprints that represent the molecule’s overall structure.
+Another approach, RDKit, is used to convert SMILES into molecular graphs (landrum2006rdkit; nguyen2021graphdta), where nodes represent the physical and chemical properties of molecules, and bonds are represented by an adjacency matrix.
 For example, atomic properties such as atom type, degree, and hydrogen information (like the number of explicit hydrogens) are all crucial for constructing a graph.
-More detailed properties can be found in Appendix F.
+More detailed properties can be found in Appendix G.
 Embedding Featurization Methods.
 Embedding methods are used to translate these discrete sequences into continuous embedding spaces.
-Notably, Smi2Vec [47] and Prot2Vec [48] convert discrete tokens of drug SMILES and protein sequences into vectors that encapsulate semantic and syntactic similarities, effectively grouping similar tokens together in vector space.
-Additionally, pretrained language models [24, 49] are increasingly utilized to leverage large-scale learned patterns, fine-tuned to analyze complex protein data representations effectively.
+Notably, Smi2Vec (8621313Smi2Vec) and Prot2Vec (10.1371/journal.pone.0141287Prot2Vec) convert discrete tokens of drug SMILES and protein sequences into vectors that encapsulate semantic and syntactic similarities, effectively grouping similar tokens together in vector space.
+Additionally, pretrained language models (bal2024pgraphdta; lin2022language) are increasingly utilized to leverage large-scale learned patterns, fine-tuned to analyze complex protein data representations effectively.
 3 A Fair Benchmark Platform Setup
@@
 These models are derived from papers spanning the years 2018 to 2024.
-We run these models on 6 frequently evaluated datasets including both binary interaction classification and continuous affinity regression.
-For the classification aspect, we utilize datasets including Human [50], Caenorhabditis elegans (C.
-elegans) [31], and DrugBank [51].
-For regression, we employ the Davis [52], KIBA [53], and BindingDB [54] with dissociation constant (Kd) measures datasets, as processed in [22].
-The statistical details of these models and datasets are presented in Section B and Table 4 of Appendix, respectively.
-Unifying Hyperparameter Configuration.
-Given the critical role of hyperparameters in achieving optimal performance, we perform a detailed review of the hyperparameters associated with the selected models in Section E of the Appendix.
-There is significant variability in the hyperparameters across different models, making it unfair to conduct comparisons directly.
-To achieve equitable comparisons between varied models, we select two representative approaches from both the GNN-based and Transformer-based categories, i.e., GraphDTA [18], GraphCPI [35], MRBDTA [55], and TransformerCPI [21], to perform a greedy hyperparameter search to find their sweet spot for classification and regression tasks, respectively.
-For the search space of hyperparameters, we mainly focus on the influence of batch size (BS), learning rate (LR), and dropout rate (DR), as these are the common hyperparameters utilized by all models.
-Additionally, we standardize the hyperparameters for epochs, weight decay, and the choice of optimizer, setting a consistent 1000 epochs for GNN-based methods and 300 epochs for Transformer-based methods, with a weight decay of 0 and the Adam optimizer for all models according to Table 5.
-We illustrate the selected results for the metrics MSE and CI for the regression task, along with AUC-ROC and accuracy for the classification task, in Fig.
-1 and the results of all metrics in Table 6.
-In all experiments, we employ the five-fold cross-validation method with a random split to evaluate all different methods and report the averaged results.
-We observe that different models exhibit distinct preferences for hyperparameters.
-Taking into account various models and metrics, we recommend the configuration as the sweet point hyperparameter configuration for the GNN-based model.
-Similarly, for the Transformer-based model, we suggest .
-We strictly follow it in the following experiments.
-4 A Macroscopical Benchmark on Encoder and Featurization Strategies
+We run these models on 6 frequently evaluated datasets, including both binary interaction classification and continuous affinity regression.
+For the classification aspect, we utilize datasets including Human (liu2015improving), Caenorhabditis elegans (C.
+elegans) (tsubaki2019compound), and DrugBank (wishart2008drugbank).
+For regression, we employ the Davis (davis2011comprehensive), KIBA (tang2014making), and BindingDB datasets (liu2007bindingdb) with dissociation constant (Kd) measures, as processed in huang2021therapeutics.
+The statistical details of these models and datasets are presented in Appendix B and Table 3, respectively.
+Hyperparameter Configuration.
+Given the critical role of hyperparameters in achieving optimal performance, we perform a systematic review of the hyperparameters for all selected models in Appendix E.
+To ensure that comparisons across models are equitable, we consider comparing each model using its optimal hyperparameters, as reported in the corresponding perspective papers.
+Data Split.
+We treat each dataset independently to prevent any information leakage that could arise from training a single model on multiple datasets.
+For duplicate drug–protein pairs in the regression dataset, only the entry with the maximum affinity score is retained.
+For duplicates in the classification dataset, all entries are removed if conflicting labels are present; otherwise, a single instance is kept.
+Following data cleaning, each dataset is split into a training set and a test set.
+We apply k-fold cross-validation on each training set, where each fold consists of a unique training–validation split, and models trained on different folds are completely independent, thereby eliminating any possibility of cross-fold leakage.
+The training sets are used to fit the models, the validation sets are used to select the best model during training, and the test sets are for final evaluation.
+Other Training Details.
+We train on the training set and see its performance on the validation set at the end of every epoch, and the model that achieves the best validation performance will be saved.
+After training, we evaluate the saved model on the test set and save their results.
+We average the final performance metrics across all folds as the final results.
+Considering the original training epochs, we use 1000 as maximum epochs limitation.
+To avoid overfitting, we consider an early stop mechanism in training.
+Given the complexity in the dataset, we use 50 patience for all datasets.
+We use MSE as an early stopping evaluation metric for regression and F1 for classification.
+The detailed results are provided in Appendix F.
+4 A Macroscopic Benchmark on Encoder and Featurization Strategies
 Encoder Exploration for Drugs and Proteins.
-To investigate the influence of different encoding strategies for extracting the structural information of drugs, we employ GCN [17] and vanilla Transformer [56] as the encoders for drugs.
+To investigate the influence of different encoding strategies for extracting the structural information of drugs, we employ GIN (gin) and vanilla Transformer (vaswani2017attention) as the encoders for drugs.
 Meanwhile, integer encoding with CNN, n-gram encoding with CNN, and the vanilla Transformer are considered to capture protein’s representations, which are frequently adopted.
-To leverage the advantages of the pretrained protein information, we include a language model, i.e., Evolutionary Scale Modeling (ESM2) [49].
-The results of various combinations of drug and protein encoders are listed in Table 1.
+To leverage the advantages of the pretrained protein information, we include a language model, i.e., Evolutionary Scale Modeling (ESM2) (lin2022language).
+The results of various combinations of drug and protein encoders are shown in Fig.
+1.All results are averaged by five-fold cross-validation with an early stop mechanism.
 Obs.
@@
 GNN and Transformer-based drug encoders exhibit unequal performance depending on DTI tasks.
-When the encoder for the protein sequence is fixed, drug features extracted by the Transformer generally perform better than those by GNNs in regression tasks, but the opposite is true in classification tasks.
-Notably, in the classification tasks on Human dataset, the combination of GNN and Transformer used, respectively, for drugs and proteins yields excellent performance but falls short in the regression task.
-This disparity may be due to the smaller size of the Human dataset compared to the Davis dataset, which allows for faster convergence in classification tasks than in regression tasks under a fixed epoch.
+When the encoder for the protein sequence is fixed, drug features extracted by the GNN structures GIN generally perform better than those by Transformers in regression tasks, but the opposite is true in classification tasks.
+This disparity may be due to the smaller size of the Human dataset compared to the Davis dataset, which allows for faster convergence in classification tasks than in regression tasks.
 Obs.
 2.
-Transformer models are better in extracting features from protein.
-Although we only consider the simplest pretrained protein language model of ESM2, it still significantly outperforms other encoders.
+Transformer models are better but sensitive in extracting features from protein.
+Although we only consider the simplest pretrained protein language model of ESM2, it still significantly outperforms other encoders in both tasks.
 This improvement can likely be attributed to the robust and generalizable representations learned from extensive data by the pretrained model.
-In the classification task, transformers achieve the best performance, underscoring their effectiveness in extracting protein sequence features.
+In addition, the Transformer encoder for the protein achieves the best performance on the classification task but shows unstable performance in the regression task.
+This is likely due to the smaller size and simpler classification dataset compared to the regression dataset, making the training stop for a fixed early stop threshold.
 Obs 3.
@@
 This implies that the simple relationships in amino acids’ immediate neighbors, as modeled by Word2Vec, do not capture much useful information compared with simple integer encoding.
-| Drug Encoder | Protein Encoder | Regression | Classification | |||||||||||
-| MSE | MAE | R2 | PCC | CI | Spearman | ROC-AUC | PR-AUC | Log-AUC | Acc. | Precision | Recall | F1 | ||
-| GCN | Int encoding with CNN | 0.3181 | 0.3314 | 0.5263 | 0.7315 | 0.8311 | 0.5906 | 0.9122 | 0.8761 | 0.4452 | 0.9127 | 0.9033 | 0.9301 | 0.9165 |
-| GCN | 3-gram pretrained with CNN | 0.3361 | 0.3613 | 0.4994 | 0.7133 | 0.8332 | 0.5956 | 0.9016 | 0.8634 | 0.4045 | 0.9021 | 0.8944 | 0.9183 | 0.9061 |
-| GCN | ESM2 | 0.2611 | 0.3112 | 0.6111 | 0.7843 | 0.8665 | 0.6480 | 0.9331 | 0.9046 | 0.5034 | 0.9334 | 0.9275 | 0.9446 | 0.9371 |
-| GCN | Trans | 0.2911 | 0.3326 | 0.5664 | 0.7600 | 0.8571 | 0.6329 | 0.9435 | 0.9192 | 0.5282 | 0.9438 | 0.9395 | 0.9524 | 0.9458 |
-| Trans | Int Encoding with CNN | 0.2553 | 0.2853 | 0.6197 | 0.7890 | 0.8620 | 0.6397 | 0.9096 | 0.8783 | 0.4623 | 0.9096 | 0.9142 | 0.9099 | 0.9121 |
-| Trans | 3-gram pretrained with CNN | 0.2618 | 0.3032 | 0.6101 | 0.7820 | 0.8581 | 0.6338 | 0.9011 | 0.8674 | 0.4327 | 0.9012 | 0.9054 | 0.9030 | 0.9040 |
-| Trans | ESM2 | 0.2609 | 0.3052 | 0.6115 | 0.7869 | 0.8617 | 0.6391 | 0.9203 | 0.8924 | 0.4902 | 0.9203 | 0.9257 | 0.9192 | 0.9224 |
-| Trans | Trans | 0.2828 | 0.3222 | 0.5788 | 0.7721 | 0.8554 | 0.6297 | 0.9306 | 0.9073 | 0.5158 | 0.9304 | 0.9391 | 0.9250 | 0.9319 |
 Featurization Exploration.
 Despite the efficacy of GNNs in learning drug structures, the featurization of nodes plays a critical role in capturing both the intrinsic properties of atoms and their contextual relevance.
-We conduct a detailed analysis of various methods (summarized in Section F of Appendix) for constructing graph features within the DTI context.
+We conduct a detailed analysis of various methods (summarized in Section G of the Appendix) for constructing graph features within the DTI context.
 The node feature is constructed via various characteristics, such as chemical and physical properties.
-We categorize each feature into five main classes, e.g., atomic properties (AP), hydrogen information (HI), electron properties (EP), stereochemistry (Ste) and structural information (Str).
+We categorize each feature into five main classes, e.g., atomic properties (AP), hydrogen information (HI), electron properties (EP), stereochemistry (Ste), and structural information (Str).
 To better determine which types of features are more effective in capturing the structural information, we conduct an ablation study on the different featurization strategies.
-Here we choose GprahDTA [18] and GraphCPI [35] as our backbone models.
+We choose GraphDTA (nguyen2021graphdta) and GraphCPI (quan2019graphcpi) with GIN as our backbone models.
 The results of feature combinations are reported in Fig.
@@
 4.
-More complex featurization does not necessarily bring positive effect.
-Despite employing different protein representations (different colors in Fig.
-2), GraphCPI and GraphDTA perform stably on the benchmark dataset Davis.
-Moreover, the basic feature configuration has the lowest MSE compared to other featurization strategies, suggesting that additional features may complicate the model’s ability to discern critical information for regression task.
-Additionally, the increased CI with more complex features, such as stereochemistry and structure, suggests that while they might not improve the prediction accuracy, they do contribute positively to the model’s ranking capability.
+More complex featurization does not necessarily bring a positive effect, and its effectiveness is highly task-dependent.
+As shown in Fig.
+2, adding features like atomic properties (AP), hydrogen information (HI), and stereochemistry information (Ste) improves performance in the regression task by reducing the MSE loss, suggesting that these features provide valuable information.
+However, features like electron properties (EP) and structural information (Str) may introduce noise rather than useful information, especially when combined with other features, leading to inconsistent results.
+Furthermore, in classification tasks, the trend differs, with additional features sometimes negatively impacting performance (blue line) rather than providing benefits.
+This highlights the importance of careful feature selection, as indiscriminate inclusion of complex features may lead to increased noise, affecting model generalization and robustness.
 Obs.
 5.
-The role of atomic and electron properties in modeling drug features may not be inherently detrimental, but their contribution appears to be context-dependent.
-In Fig.
-2, it suggests that these properties, when not combined with other informative features, don’t significantly enhance two models’ performance.
-In the classification task, as reflected by the F1 scores and ROC-AUC, it is evident that stereochemistry and structural information substantially improve the model’s accuracy.
-Thus, while atomic and electron properties are fundamental, their full potential is unlocked when integrated with stereochemical and structural information, underscoring the importance of a multifaceted approach in node featurization.
+Protein representation plays a crucial role in the effectiveness of different drug featurizations.
+As shown in Fig.
+2, even when using the same drug featurization strategies, the trends vary depending on the protein representation.
+This suggests that the way proteins are encoded directly influences how drug features interact with the model.
+For instance, in GraphDTA, features like atomic properties (AP) and hydrogen information (HI) help improve performance, but in GraphCPI, those same features don’t always provide the same benefits.
+In some cases, adding more drug features introduces noise rather than useful information.
+This highlights the fact that drug featurization and protein representation are deeply interconnected, and optimizing one without considering the other may not yield the best results.
+To build more effective drug-protein interaction models, both components should be considered holistically rather than in isolation.
 5 A Microscopic Benchmark on DTI Models
 Benchmark over Effectiveness.
-As shown in Table.
-2 and Table.
-3, we conduct experiments on 31 different models across two tasks and three datasets, respectively (see more comprehensive model comparisons in Section G of Appendix).
-All results are averaged by five-fold cross-validation.
+As shown in Table 1 and Table 2, we conduct experiments on models with their optimal hyperparameters across two tasks and three datasets, respectively (see more comprehensive model comparisons in Appendix F).
+All results are averaged by five-fold cross-validation with an early stop mechanism.
 Obs.
 6.
-Molecular graphs are better than fingerprints to capture the graph features of drug.
-In reference to Table.
-7, it is evident that GNN-based approaches utilizing the molecular graph generally yield superior performance compared with fingerprints (CPI [31], BACPI [57], GANDTI [58]).
-This reinforces the idea that the rich structural and atom property information inherent to molecular graphs is pivotal for representation extraction, leading to enhanced model performance.
+Molecular graphs are better than fingerprints for capturing the graph features of a drug.
+In reference to Table 7, it is evident that GNN-based approaches utilizing the molecular graph generally yield superior performance compared with fingerprints (CPI (tsubaki2019compound), BACPI (li2022bacpi), GANDTI (wang2021gandti)).
+This reinforces the idea that the rich structural and atomic property information inherent to molecular graphs is pivotal for representation extraction, leading to enhanced model performance.
 Obs.
 7.
-Graph structure is a crucial part in extracting drug’s features.
-Different GNNs have the distinct performances in both tasks when the protein representation is fixed.
+Graph structure is a crucial part of extracting a drug’s features.
+Different GNNs have distinct performances in both tasks when the protein representation is fixed.
 Specifically, GIN, with its unique ability to distinguish non-isomorphic graphs, consistently outperforms other models across different protein encoders in regression tasks.
 Although Transformer-based methods such as MRBDTA are proficient in handling sequential information from SMILES and proteins, the depth of information they capture appears to be marginally less comprehensive than that provided by molecular graph-based approaches.
-This is substantiated by the superior performance of GNN-based methods, including MGraphDTA, ColdDTA, and SubMDTA, which suggest that GNN captures intricate structural details more effectively.
-| Category | Models | DAVIS | KIBA | BindingDB_Kd | Avg. Reduce | ||||||
+This is substantiated by the superior performance of GNN-based methods, including MGraphDTA, ColdDTA, and SubMDTA, which suggests that GNN captures intricate structural details more effectively.
+| Category | Models | DAVIS | KIBA | BindingDB_Kd | Avg. Reduction | ||||||
 | MSE | R2 | CI | MSE | R2 | CI | MSE | R2 | CI | of MSE (%) | ||
-| GNN | GraphDTA-GIN [18] | 0.2309 | 0.6562 | 0.8711 | 0.0004 | 0.5544 | 0.7996 | 0.5033 | 0.7259 | 0.8576 | 8.5625 |
-| GraphCPI-GIN [35] | 0.2413 | 0.6407 | 0.8671 | 0.0005 | 0.4963 | 0.7808 | 0.5069 | 0.7239 | 0.8572 | 10.2845 | |
-| MGraphDTA [59] | 0.2179 | 0.6755 | 0.8820 | 0.0003 | 0.7208 | 0.8649 | 0.4887 | 0.7338 | 0.8649 | 4.9795 | |
-| SAGDTA [60] | 0.2656 | 0.6045 | 0.8675 | 0.0039 | -3.3777 | 0.8096 | 0.6590 | 0.6410 | 0.8366 | 27.6575 | |
-| EmbedDTI [61] | 0.2561 | 0.6186 | 0.8624 | 0.0007 | 0.2364 | 0.6374 | 0.5095 | 0.7225 | 0.8559 | 12.3450 | |
-| DeepGLSTM [29] | 0.2915 | 0.5659 | 0.8476 | 0.0003 | 0.6919 | 0.8480 | 0.5385 | 0.7067 | 0.8529 | 19.1015 | |
-| CPI [31] | 0.3503 | 0.4784 | 0.8319 | 0.0003 | 0.6274 | 0.8190 | 0.6962 | 0.6208 | 0.8261 | 35.8330 | |
-| BACPI [57] | 0.4036 | 0.3990 | 0.7982 | 0.0006 | 0.5320 | 0.8175 | 0.6468 | 0.6477 | 0.8297 | 36.0894 | |
-| DeepNC-HGC [30] | 0.2782 | 0.5857 | 0.8551 | 0.0005 | 0.4983 | 0.7823 | 0.5611 | 0.6944 | 0.8464 | 20.0167 | |
-| DeepNC-GEN [30] | 0.2543 | 0.6213 | 0.8634 | 0.0004 | 0.5355 | 0.7981 | 0.5561 | 0.6971 | 0.8496 | 17.1559 | |
-| DrugBAN [62] | 0.2391 | 0.6440 | 0.8757 | 0.0004 | 0.5611 | 0.8388 | 0.4485 | 0.7557 | 0.8693 | 2.3692 | |
-| GANDTI [58] | 0.3082 | 0.5410 | 0.8414 | 0.0003 | 0.6569 | 0.8342 | 0.6714 | 0.6343 | 0.8322 | 31.4522 | |
-| PGraphDTA-CNN [24] | 0.3273 | 0.5126 | 0.8701 | 0.0005 | 0.4833 | 0.7473 | 0.5334 | 0.7095 | 0.8590 | 22.0042 | |
-| BridgeDPI [63] | 0.3623 | 0.6477 | 0.8991 | 0.0004 | 0.5686 | 0.7849 | 0.4482 | 0.7559 | 0.8698 | 17.1661 | |
-| ColdDTA [64] | 0.2346 | 0.6507 | 0.8693 | 0.0004 | 0.5948 | 0.8018 | 0.4697 | 0.7442 | 0.8644 | 4.6828 | |
-| SubMDTA [36] | 0.2326 | 0.6537 | 0.8691 | 0.0003 | 0.6855 | 0.8485 | 0.4566 | 0.7513 | 0.8670 | 2.5816 | |
-| IMAEN [65] | 0.2412 | 0.6409 | 0.8721 | 0.0004 | 0.5800 | 0.8061 | 0.4720 | 0.7429 | 0.8553 | 5.8716 | |
-| Transformer | CSDTI [32] | 0.3029 | 0.5490 | 0.8395 | 0.0007 | 0.2448 | 0.6475 | 0.6408 | 0.6510 | 0.8369 | 28.8755 |
-| TDGraphDTA [25] | 0.2217 | 0.6698 | 0.6685 | 0.0008 | 0.0533 | 0.3429 | 0.4750 | 0.7413 | 0.7887 | 3.6989 | |
-| AMMVF [38] | 0.3325 | 0.5048 | 0.8307 | 0.0006 | 0.4696 | 0.7711 | 0.6597 | 0.6407 | 0.8336 | 32.3429 | |
-| IIFDTI [66] | 0.2741 | 0.5918 | 0.8500 | 0.0005 | 0.2584 | 0.7952 | 0.5097 | 0.7170 | 0.8585 | 14.3568 | |
-| ICAN [42] | 0.3481 | 0.4816 | 0.8211 | 0.0008 | 0.1159 | 0.8256 | 0.6582 | 0.6415 | 0.8277 | 33.3035 | |
-| MolTrans [20] | 0.2588 | 0.6146 | 0.8601 | 0.0003 | 0.6378 | 0.8453 | 0.5138 | 0.7201 | 0.8570 | 13.0935 | |
-| TransformerCPI [21] | 0.2869 | 0.5728 | 0.8326 | 0.0008 | 0.0728 | 0.8357 | 0.5704 | 0.6894 | 0.8426 | 21.7224 | |
-| MRBDTA [55] | 0.2350 | 0.6499 | 0.8775 | 0.0006 | 0.3210 | 0.7239 | 0.4977 | 0.7289 | 0.8629 | 8.4004 | |
-| FOTFCPI [41] | 0.2803 | 0.5825 | 0.8546 | 0.0004 | 0.5342 | 0.7947 | 0.5743 | 0.6872 | 0.8444 | 21.4386 | |
-| Our combos | 0.2063 | 0.6927 | 0.8901 | 0.0003 | 0.7168 | 0.8677 | 0.4651 | 0.7467 | 0.8683 | 0 | |
-| Category | Models | Human | C.elegans | Drugbank | Avg. Improve | ||||||
+| GNN | GraphDTA-GIN | ||||||||||
+| GraphCPI-GIN | |||||||||||
+| MGraphDTA | |||||||||||
+| SAGDTA | |||||||||||
+| EmbedDTI | |||||||||||
+| DeepGLSTM | |||||||||||
+| CPI | |||||||||||
+| BACPI | |||||||||||
+| DeepNC-HGC | |||||||||||
+| DeepNC-GEN | |||||||||||
+| DrugBAN | |||||||||||
+| GANDTI | |||||||||||
+| BridgeDPI | |||||||||||
+| ColdDTA | |||||||||||
+| SubMDTA | |||||||||||
+| IMAEN | |||||||||||
+| Transformer | CSDTI | ||||||||||
+| TDGraphDTA | |||||||||||
+| AMMVF | |||||||||||
+| IIFDTI | |||||||||||
+| ICAN | |||||||||||
+| MolTrans | |||||||||||
+| TransformerCPI | |||||||||||
+| MRBDTA | |||||||||||
+| FOTFCPI | |||||||||||
+| Our combos | |||||||||||
+| Category | Models | Human | C.elegans | Drugbank | Avg. Improvement | ||||||
 | ROC-AUC | Accuracy | F1 | ROC-AUC | Accuracy | F1 | ROC-AUC | Accuracy | F1 | of Accuracy (%) | ||
-| GNN | GraphDTA-GATGCN[18] | 0.9296 | 0.9297 | 0.9319 | 0.9487 | 0.9484 | 0.9485 | 0.7712 | 0.7711 | 0.7764 | 2.5140 |
-| GraphCPI-GATGCN[35] | 0.9097 | 0.9099 | 0.9129 | 0.9372 | 0.9370 | 0.9369 | 0.7561 | 0.7561 | 0.7593 | 4.3335 | |
-| MGraphDTA[59] | 0.9408 | 0.9410 | 0.9429 | 0.9631 | 0.9630 | 0.9628 | 0.8146 | 0.8146 | 0.8157 | -0.1030 | |
-| SAGDTA[60] | 0.9021 | 0.9018 | 0.9032 | 0.9380 | 0.9378 | 0.9375 | 0.7655 | 0.7655 | 0.7662 | 4.2494 | |
-| EmbedDTI[61] | - | - | - | - | - | - | 0.7625 | 0.7625 | 0.7674 | 5.3770 | |
-| DeepGLSTM[29] | 0.9180 | 0.9183 | 0.9212 | 0.9414 | 0.9412 | 0.9409 | 0.7640 | 0.7640 | 0.7674 | 3.5182 | |
-| CPI[31] | 0.9110 | 0.9111 | 0.9138 | 0.9312 | 0.9310 | 0.9309 | 0.7467 | 0.7466 | 0.7515 | 4.9098 | |
-| BACPI[57] | 0.9249 | 0.9251 | 0.9276 | 0.9556 | 0.9555 | 0.9551 | 0.7748 | 0.7748 | 0.7776 | 2.2746 | |
-| DeepNC-HGC[30] | 0.8796 | 0.8798 | 0.8834 | 0.9418 | 0.9417 | 0.9413 | 0.7673 | 0.7673 | 0.7718 | 4.9057 | |
-| DeepNC-GEN[30] | 0.9178 | 0.9180 | 0.9205 | 0.9501 | 0.9499 | 0.9498 | 0.7464 | 0.7463 | 0.7571 | 3.8865 | |
-| DrugBAN[62] | 0.9302 | 0.9302 | 0.9320 | 0.9596 | 0.9594 | 0.9593 | 0.8188 | 0.8188 | 0.8179 | 0.2732 | |
-| GANDTI [58] | 0.9333 | 0.9333 | 0.9351 | 0.9404 | 0.9403 | 0.9400 | 0.7427 | 0.7425 | 0.7550 | 3.8110 | |
-| PGraphDTA-CNN [24] | - | - | - | - | - | - | 0.7910 | 0.7910 | 0.7948 | 1.5803 | |
-| BridgeDPI [63] | 0.9456 | 0.9456 | 0.9471 | 0.9651 | 0.9651 | 0.9646 | 0.7825 | 0.7825 | 0.7853 | 0.8392 | |
-| ColdDTA [64] | 0.9420 | 0.9421 | 0.9436 | 0.9623 | 0.9622 | 0.9619 | 0.8186 | 0.8186 | 0.8203 | -0.2608 | |
-| SubMDTA [36] | 0.9326 | 0.9328 | 0.9351 | 0.9570 | 0.9568 | 0.9566 | 0.8045 | 0.8045 | 0.8072 | 0.8055 | |
-| IMAEN [65] | 0.9058 | 0.9059 | 0.9089 | 0.9506 | 0.9505 | 0.9500 | 0.7995 | 0.7994 | 0.8014 | 2.2592 | |
-| Transformer | CSDTI [32] | 0.8630 | 0.8630 | 0.8663 | 0.8962 | 0.8958 | 0.8966 | 0.7269 | 0.7269 | 0.7306 | 9.2569 |
-| TDGraphDTA [25] | 0.9411 | 0.9409 | 0.9419 | 0.9573 | 0.9571 | 0.9569 | 0.8083 | 0.8082 | 0.8116 | 0.3547 | |
-| AMMVF [38] | 0.9287 | 0.9290 | 0.9314 | 0.9636 | 0.9635 | 0.9632 | 0.7814 | 0.7814 | 0.7849 | 1.5670 | |
-| IIFDTI [66] | 0.9392 | 0.9392 | 0.9409 | 0.9679 | 0.9679 | 0.9675 | 0.8084 | 0.8084 | 0.8125 | 0.0110 | |
-| ICAN [42] | 0.9387 | 0.9389 | 0.9410 | 0.9644 | 0.9644 | 0.9637 | 0.7731 | 0.7731 | 0.7745 | 1.4721 | |
-| MolTrans [20] | 0.9453 | 0.9455 | 0.9471 | 0.9639 | 0.9638 | 0.9634 | 0.7910 | 0.7909 | 0.7922 | 0.5777 | |
-| TransformerCPI [21] | 0.9311 | 0.9312 | 0.9333 | 0.9581 | 0.9580 | 0.9577 | 0.8097 | 0.8097 | 0.8125 | 0.6262 | |
-| MRBDTA [55] | 0.9447 | 0.9446 | 0.9458 | 0.9713 | 0.9713 | 0.9709 | 0.8102 | 0.8102 | 0.8134 | -0.3778 | |
-| FOTFCPI [41] | 0.9444 | 0.9444 | 0.9459 | 0.9673 | 0.9672 | 0.9669 | 0.7845 | 0.7845 | 0.7889 | 0.7307 | |
-| Our combos | 0.9435 | 0.9435 | 0.9451 | 0.9688 | 0.9688 | 0.9684 | 0.8035 | 0.8035 | 0.8061 | 0 | |
+| GNN | GraphDTA-GIN | 3.312% | |||||||||
+| GraphCPI-GIN | 4.275% | ||||||||||
+| MGraphDTA | 1.322% | ||||||||||
+| SAGDTA | 4.600% | ||||||||||
+| EmbedDTI | 3.871% | ||||||||||
+| DeepGLSTM | 4.356% | ||||||||||
+| CPI | 7.708% | ||||||||||
+| BACPI | 2.522% | ||||||||||
+| DeepNC-HGC | 6.006% | ||||||||||
+| DeepNC-GEN | 4.194% | ||||||||||
+| DrugBAN | 0.675% | ||||||||||
+| GANDTI | 3.671% | ||||||||||
+| BridgeDPI | 11.189% | ||||||||||
+| ColdDTA | 0.037% | ||||||||||
+| SubMDTA | 1.055% | ||||||||||
+| IMAEN | 4.560% | ||||||||||
+| Transformer | CSDTI | 11.467% | |||||||||
+| TDGraphDTA | 0.299% | ||||||||||
+| AMMVF | 6.595% | ||||||||||
+| IIFDTI | 1.437% | ||||||||||
+| ICAN | 1.899% | ||||||||||
+| MolTrans | 0.525% | ||||||||||
+| TransformerCPI | 0.979% | ||||||||||
+| MRBDTA | 0.789% | ||||||||||
+| FOTFCPI | 0.112% | ||||||||||
+| Our combos | 0.000% | ||||||||||
 Benchmark over Efficiency To analyze the training speed and memory usage, we empirically evaluate the peak memory and running time for various methods during the training procedure on one regression dataset and one classification task, respectively.
@@
 The memory and running time comparisons are illustrated in Fig.
-3.
+4.
 Obs.
@@
 Benchmark over Convergence.
-We select the top two methods from the GNN-based and Transformer-based frameworks, respectively, and evaluate them across six datasets on two tasks.
+We select the two representative methods from the GNN-based and Transformer-based frameworks, respectively, and evaluate them across six datasets on two tasks.
 The training losses are depicted in Fig.
 4.
-In order to compare different methods, we only show the epochs before 300.
-Based on the empirical data, we summarize our primary observations as follows:
+To ensure a fair comparison of convergence behavior, we use the previous early stopping setting.
+Based on the empirical results, we summarize our key observations as follows:
 Obs.
 9.
-GNN-based methods demonstrate quicker and more stable convergence compared to Transformer-based methods.
-This phenomenon arises from the fact that GNN-based methods have fewer memory usage and model parameters, leading to a larger batch size usage or faster convergence compared with Transformer-based methods.
+GNN-based methods demonstrate quicker convergence compared to Transformer-based methods.
+This phenomenon arises from the fact that GNN-based methods have less memory usage and fewer model parameters, leading to larger batch size usage or faster convergence compared with Transformer-based methods.
 5.1 Our Best Combo of Drug and Protein Encoders
-Based on our benchmark results, we summarize the insights of protein and drug encoder usages and propose a light yet effective architecture, which could be treated as new strong baseline for the following explorations.
-Regarding the proteins, we observe that multi-scale CNNs associated with a mixture of model depths can generally learn the effective protein representations [59, 25, 64], which approximates the language model’s accuracy while having lower memory and computation costs.
-Regarding the drug molecules, both GNN and Transformer-based methods, such as MRBDTA [55], MolTrans [20] and MGraphDTA [59] prove promising in DTI tasks.
+Deriving Combo from Benchmark Insights.
+Based on our benchmark results, we summarize the insights of protein and drug encoder usages and propose a light yet effective architecture, which could be treated as a new strong baseline for future explorations.
+Regarding the proteins, we observe that multi-scale CNNs associated with a mixture of model depths can generally learn the effective protein representations (yang2022MGraphDTA; zhu2023tdgraph; fang2023colddta), which approximate the language model’s accuracy while having lower memory and computation costs.
+Regarding the drug molecules, both GNN- and Transformer-based methods, such as MRBDTA (zhang2022predicting), MolTrans (huang2021moltrans), and MGraphDTA (yang2022MGraphDTA) prove promising in DTI tasks.
 This encourages us to leverage information from hybrid perspectives, i.e., implicit structure (via attention in Transformers) and explicit structure learning (via message passing along edges in GNNs).
-We are thus motivated to integrate these powerful modules and shed novel insight into the design philosophy of drug-target interaction modeling.
-Our model combos are illustrated in Fig.
-5, where the multi-scale CNNs and hybrid networks of molecular Transformer and GNNs are adopted to learn the representations of proteins and drugs, respectively.
-As shown by the graph encoder part in Fig.
-5, the hybrid networks augment the differential attention matrix in molecular Transformer with inter-atomic distances and graph adjacency matrix [67], which provides the 3D and 2D molecule conformations to further facilitate atom interaction learning.
-Specifically, given the projections of molecular input at an attention head, i.e., , the adjacent matrix , and the inter-atomic distances matrix obtained using RDkit, the augmented attention is calculated as:
+Our model design, illustrated in Fig.
+5, integrates these components.
+Specifically, for drug graphs, we adopt a hybrid network that augments the self-attention mechanism with inter-atomic distances and graph adjacency matrices (maziarka2020molecule), incorporating both 2D and 3D molecular structural information.
+Given the projections of molecular input at an attention head, i.e., the adjacent matrix , and the inter-atomic distances matrix obtained using RDkit, the augmented attention is calculated as follows:
 where is a row-wise softmax function, and and denote scalars weighting the self-attention, distance, and adjacency matrices, respectively.
@@
 It is notable that simply utilizing the SMILES representation extracted from a Transformer for downstream tasks does not perform as well as GNN.
-To align with the protein embedding paradigm, we adopt a simple CNN to unearth potential SMILES information, as suggested in [68].
+To align with the protein embedding paradigm, we adopt a simple CNN to unearth potential SMILES information, as suggested in zhao2022hyperattentiondti.
 Subsequently, due to the fact that cross-attention is more complex and hard to optimize, we implement a straightforward attention mechanism to integrate the representations of the drug graph and SMILES, denoted as and , respectively, using a weighting parameter , as follows:
 Finally, the prediction is obtained by processing the concatenated protein and drug representations through a task-relevant head, as shown in Fig. 5
+Novelty.
+As opposed to the previous strategy, which heuristically stacked a large amount of modules of different types, our model design is driven by systematic benchmarking and empirical insights.
+Through extensive experiments under fair and controlled conditions, we identify key encoder and featurization strategies that consistently outperform others.
+Notably, we disentangle and quantify the distinct molecular features.
+For instance, atomic properties and hydrogen information significantly enhance predictive performance, while adding electron properties may introduce noise.
+Thus, our combo is not merely an ad hoc combination but a carefully validated design that effectively obtains a superior balance between accuracy and computational efficiency on both classification and regression tasks.
+The clear empirical guidance to model design offered through this study helps to establish a more principled framework for future work in drug-target interaction modeling and provides a robust, reproducible new baseline for the community.
 Benchmark Comparison to State-of-the-Art Frameworks.
-We compare the proposed combos with the SOTA frameworks in Tables 2 and 3, and Figures 3 and 4.
+We compare the proposed combos with the SOTA frameworks in Tables˜1 and 2, and Figures˜4 and 4.
 It is observed that our model consistently achieves the best performance in the regression tasks across three datasets and nearly outperforms most methods in classification tasks.
-By leveraging the physical conformation information from the molecular graph, our combos converge faster than the other two Transformer-based methods, MRBDTA [55] and TDGraphDTA [25], particularly on the KIBA dataset.
-Moreover, our model uses three times less peak memory and fewer parameters than other Transformer-based methods, enabling faster computation and reduced storage requirements
+By leveraging the physical conformation information from the molecular graph, our combos converge faster than the other two Transformer-based methods, MRBDTA (zhang2022predicting) and TDGraphDTA (zhu2023tdgraph), particularly on the KIBA dataset.
+Moreover, our model uses three times less peak memory and fewer parameters than other Transformer-based methods, enabling faster computation and reduced storage requirements.
 6 Conclusion
@@
 GNN-based Methods GNNs play a crucial role in mining the intricate features of drug molecules for drug-target prediction.
-Numerious models, including Graph Convolutional Network (GCN), Graph Isomorphism Network (GIN), and Graph Attention Network (GAT) have been utilized [18, 35, 38, 69, 69, 61] to process and enhance drug features.
-Additionally, MGraphDTA [59] employs a multi-scale GNN architecture, while DeepGLSTM [29] leverages parallel GNN structures for drug representation.
-DeepNC integrates advanced techniques from generalized aggregation networks [70] and hypergraph convolution [71] to improve feature extraction.
-BACPI [57] develops a bi-directional attention network to integrate the representations of drug molecules and proteins, enhancing their mutual interaction.
-Besides, BridgeDPI [63] innovates by incorporating bridging nodes between proteins and drugs, utilizing a three-layer GNN for graph embeddings.
+Numerous models, including Graph Convolutional Network (GCN), Graph Isomorphism Network (GIN), and Graph Attention Network (GAT), have been utilized (nguyen2021graphdta; quan2019graphcpi; wang2023ammvf; lin2020deepgs; jin2021embeddti) to process and enhance drug features.
+Additionally, MGraphDTA (yang2022MGraphDTA) employs a multi-scale GNN architecture, while DeepGLSTM (mukherjee2022deepglstm) leverages parallel GNN structures for drug representation.
+DeepNC integrates advanced techniques from generalized aggregation networks (li2020deepergcn) and hypergraph convolution (BAI2021107637) to improve feature extraction.
+BACPI (li2022bacpi) develops a bi-directional attention network to integrate the representations of drug molecules and proteins, enhancing their mutual interaction.
+Besides, BridgeDPI (wu2022bridgedpi) innovates by incorporating bridging nodes between proteins and drugs, utilizing a three-layer GNN for graph embeddings.
 Transformer-based Methods Transformers, known for their efficacy in handling sequence data, are extensively applied in drug and protein feature processing.
-For instance, models like MolTrans [20] and FOTFCPI [41] employ self-attention mechanisms to refine embeddings by focusing on drug and protein substructures.
-MRBDTA [55] uses multi-head attention and skip connection to enhance drug and protein representation.
-Additionally, a cross-attention mechanism [32, 42] is employed to facilitate the integration of drug and protein features, enabling effective mutual querying.
-TDGraphDTA [25] captures contextual relationships between molecular substructures by using a multi-head cross-attention mechanism and graph optimization.
-Lastly, DrugormerDTI [72] incorporates degree centrality with positional information to highlight the positional relevance of amino acids in proteins.
-Input and Featurization Structural information is crucial at the input stage for models such as BridgeDPI [63].
-Various libraries, such as DGLGraph [73], DGL-lifeSci [74], and RDKit [46], are employed to process input SMILES of drugs, with RDKit [46] being pivotal for converting SMILE strings into molecular graphs and extracting diverse chemical properties, including chemical bonds, hydrogen presence, electron properties, and so on.
-Additionally, some approaches [38, 69, 57, 58] incorporate molecular fingerprints [45] to capture local chemical information.
-For protein sequences, typical pre-processing involves converting amino acid sequences into N-grams [36, 43] or integers [18] sequences.
-To enhance the expressiveness of embeddings, some models leverage pre-trained Word2Vec [75, 35, 38, 57, 31, 69, 66] or pretrained protein language models [24].
+For instance, models like MolTrans (huang2021moltrans) and FOTFCPI (yin2024fotf) employ self-attention mechanisms to refine embeddings by focusing on drug and protein substructures.
+MRBDTA (zhang2022predicting) uses multi-head attention and skip connection to enhance drug and protein representation.
+Additionally, a cross-attention mechanism (pan2023csdti; kurata2022ican) is employed to facilitate the integration of drug and protein features, enabling effective mutual querying.
+TDGraphDTA (zhu2023tdgraph) captures contextual relationships between molecular substructures by using a multi-head cross-attention mechanism and graph optimization.
+Lastly, DrugormerDTI (hu2023drugormerdti) incorporates degree centrality with positional information to highlight the positional relevance of amino acids in proteins.
+Input and Featurization Structural information is crucial at the input stage for models such as BridgeDPI (wu2022bridgedpi).
+Various libraries, such as DGLGraph (wang2019DGL), DGL-lifeSci (li2021dgl-lifesci), and RDKit (landrum2006rdkit), are employed to process input SMILES of drugs, with RDKit (landrum2006rdkit) being pivotal for converting SMILE strings into molecular graphs and extracting diverse chemical properties, including chemical bonds, hydrogen presence, electron properties, and so on.
+Additionally, some approaches (wang2023ammvf; lin2020deepgs; li2022bacpi; wang2021gandti) incorporate molecular fingerprints (doi:10.1021/ci100050tecfp) to capture local chemical information.
+For protein sequences, typical preprocessing involves converting amino acid sequences into N-grams (pan2023submdta; 10.1093/bioinformatics/bti801) or integer (nguyen2021graphdta) sequences.
+To enhance the expressiveness of embeddings, some models leverage pre-trained Word2Vec (10.5555/2999792.2999959word2vec; quan2019graphcpi; wang2023ammvf; li2022bacpi; tsubaki2019compound; lin2020deepgs; cheng2022iifdti) or pre-trained protein language models (bal2024pgraphdta).
 Appendix B Model Descriptions
@@
 B.1.1 GCN
-GraphDTA-GCN [18]: GraphDTA-GCN uses GCN to process the molecular graph, which is derived from SMILES using the RDkit tool, and a simple CNN with integer encoding to handle protein sequences.
-GraphCPI-GCN [35]: Similar to GraphDTA, GraphCPI-GCN employs 3-gram encoding with pretrained Word2Vec to process protein sequences, followed by a CNN to handle the protein embeddings.
-MGraphDTA [59]: MGraphDTA utilizes a multiscale GCN, inspired by dense connections, and a multiscale CNN to process drug graphs and protein sequences, respectively.
-SAGDTA [60]: Similar to GraphDTA, SAGDTA introduces global or hierarchical pooling after GCN to aggregate node representations weightedly.
-EmbedDTI [61]: For protein sequences, EmbedDTI leverages GloVe for pretraining amino acid feature embeddings, which are then fed into a CNN.
+GraphDTA-GCN (nguyen2021graphdta): GraphDTA-GCN uses GCN to process the molecular graph, which is derived from SMILES using the RDkit tool, and a simple CNN with integer encoding to handle protein sequences.
+GraphCPI-GCN (quan2019graphcpi): Similar to GraphDTA, GraphCPI-GCN employs 3-gram encoding with pretrained Word2Vec to process protein sequences, followed by a CNN to handle the protein embeddings.
+MGraphDTA (yang2022MGraphDTA): MGraphDTA utilizes a multiscale GCN, inspired by dense connections, and a multiscale CNN to process drug graphs and protein sequences, respectively.
+SAGDTA (zhang2021sag): Similar to GraphDTA, SAGDTA introduces global or hierarchical pooling after GCN to aggregate node representations weightedly.
+EmbedDTI (jin2021embeddti): For protein sequences, EmbedDTI leverages GloVe for pretraining amino acid feature embeddings, which are then fed into a CNN.
 For drugs, it constructs both an atom graph and a substructure graph to capture structural information at different levels, processed by GCN.
-DeepGLSTM [29]: DeepGLSTM processes molecular graphs using a parallel GCN module composed of three GCNs with different layers. For protein sequences, it adopts a bi-LSTM.
-CPI [31]: CPI processes drug graphs using GCN. The protein sequence is handled via n-gram with integer encoding, followed by a CNN.
-DeepNC [30]: DeepNC adopts advanced techniques from generalized aggregation networks and hypergraph convolution, two variants of GCN, to capture the representations of drug.
+DeepGLSTM (mukherjee2022deepglstm): DeepGLSTM processes molecular graphs using a parallel GCN module composed of three GCNs with different layers. For protein sequences, it adopts a bi-LSTM.
+CPI (tsubaki2019compound): CPI processes drug graphs using GCN. The protein sequence is handled via n-gram with integer encoding, followed by a CNN.
+DeepNC (tran2022deepnc): DeepNC adopts advanced techniques from generalized aggregation networks and hypergraph convolution, two variants of GCN, to capture the representations of drugs.
 For protein sequences, it uses a simple CNN.
-DrugBAN [55]: DrugBAN employs GCN and CNN blocks to encode molecular graph and proteins, respectively.
+DrugBAN (zhang2022predicting): DrugBAN employs GCN and CNN blocks to encode molecular graphs and proteins, respectively.
 Then they use a bilinear attention network module to learn local interactions between the representations of drugs and proteins.
-BridgeDPI [63]: BridgeDPI innovates by constructing a learnable drug–protein association network, which is processed using a three-layer GNN for graph embeddings.
+BridgeDPI (wu2022bridgedpi): BridgeDPI innovates by constructing a learnable drug–protein association network, which is processed using a three-layer GNN for graph embeddings.
 The learned representations for drug and protein pairs are then concatenated for further processing.
-ColdDTA [64]: ColdDTA removes the subgraphs of drugs.
+ColdDTA (fang2023colddta): ColdDTA removes the subgraphs of drugs.
 For the model, they adopt the dense GCN and multiscale CNN from MGraphDTA as the encoders for drugs and proteins, respectively.
 Additionally, an attention-based method is developed to integrate representations for improved prediction.
-IMAEN [65]: IMAEN employs a molecular augmentation mechanism to enhance molecular structures by fully aggregating molecular node neighborhood information.
+IMAEN (zhang2024imaen): IMAEN employs a molecular augmentation mechanism to enhance molecular structures by fully aggregating molecular node neighborhood information.
 It then uses multiscale GCN and CNN for drug and protein processing, respectively.
-GanDTI [58]: Inspired by residual networks, GanDTI add the input drug fingerprints to the output of three GCN layers as graph node features and use summation to get the final drug representation.
+GanDTI (wang2021gandti): Inspired by residual networks, GanDTI adds the input drug fingerprints to the output of three GCN layers as graph node features and uses summation to get the final drug representation.
 B.1.2 GAT
-GraphDTA-GAT [18]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
-GraphDTA-GATGCN [18]: GraphDTA-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
-GraphCPI-GAT [35]: GraphDTA-CPI adopts a GAT as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
-GraphCPI-GATGCN [35]: GraphCPI-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
-BACPI [57]: BACPI adopts a GAT and a CNN for the features of the fingerprints and protein sequence, respectively.
-These features are then fed into a bi-directional attention neural network to obtain integrated representations.
-PGraphDTA-CNN [24]: PGraphDTA-CNN is a straightforward method that utilizes GAT for drug feature extraction and CNN for protein sequences.
+GraphDTA-GAT (nguyen2021graphdta): GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
+GraphDTA-GATGCN (nguyen2021graphdta): GraphDTA-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
+GraphCPI-GAT (quan2019graphcpi): GraphDTA-CPI adopts a GAT as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
+GraphCPI-GATGCN (quan2019graphcpi): GraphCPI-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
+BACPI (li2022bacpi): BACPI adopts a GAT and a CNN for the features of the fingerprints and protein sequence, respectively.
+These features are then fed into a bidirectional attention neural network to obtain integrated representations.
+PGraphDTA-CNN (bal2024pgraphdta): PGraphDTA-CNN is a straightforward method that utilizes GAT for drug feature extraction and CNN for protein sequences.
 B.2 GIN
-GraphDTA-GIN [18]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
-GraphCPI-GIN [35]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
-SubMDTA [36]: SubMDTA utilizes a pretrained GIN encoder obtained through contrastive learning for the molecular graph.
+GraphDTA-GIN (nguyen2021graphdta): GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
+GraphCPI-GIN (quan2019graphcpi): GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
+SubMDTA (pan2023submdta): SubMDTA utilizes a pretrained GIN encoder obtained through contrastive learning for the molecular graph.
 For protein sequences, it employs N-gram embedding with different N to extract features at various scales, which are then processed by a BiLSTM.
@@
 B.3.1 Self-attention
-AMMVF [38]: AWMVF introduces the multi-head mechanism to GAT to learn features in different spaces, and the update function is obtained through the concatenation of different heads’ outputs.
-IIFDTI [66]: IIFDTI model attains the drug matrix and protein matrix and inputs them to the bi-directional encoder-decoder block, which considers both the drug and target directions.
+AMMVF (wang2023ammvf): AWMVF introduces the multi-head mechanism to GAT to learn features in different spaces, and the update function is obtained through the concatenation of different heads’ outputs.
+IIFDTI (cheng2022iifdti): IIFDTI model attains the drug matrix and protein matrix and inputs them to the bi-directional encoder-decoder block, which considers both the drug and target directions.
 The decoder is mainly composed of multi-head attention.
-MolTrans [20]: MolTrans uses transformer encoder layers to augment the embedding of sub-structure sequences of proteins and drugs.
-FOTFCPI [41]: Similar to MolTrans, FOTFCPI uses transformer encoder layers to extract the features of protein and drug fragments after the embedding layers.
-TransformerCPI [21]: TransformerCPI uses the decoder module of Transformer, which takes in the atom sequence embedding processed by GCN and the protein sequence embedding processed by word2vec and 1D CNN.
-MRBDTA [55]: In MRBDTA, after the embedding layer, drug sequences are directly fed into a block consisting of three Transformer encoders.
+MolTrans (huang2021moltrans): MolTrans uses Transformer encoder layers to augment the embedding of substructure sequences of proteins and drugs.
+FOTFCPI (yin2024fotf): Similar to MolTrans, FOTFCPI uses Transformer encoder layers to extract the features of protein and drug fragments after the embedding layers.
+TransformerCPI (chen2020Transformercpi): TransformerCPI uses the decoder module of Transformer, which takes in the atom sequence embedding processed by GCN and the protein sequence embedding processed by word2vec and 1D CNN.
+MRBDTA (zhang2022predicting): In MRBDTA, after the embedding layer, drug sequences are directly fed into a block consisting of three Transformer encoders.
 The first encoder has a linear layer before it and the following two encoders are parallel.
-The protein sequence is also processed by a block with similar structure.
+The protein sequence is also processed by a block with a similar structure.
 B.3.2 Cross attention
-CSDTI[32]: CSDTI use cross attention to fuse the deep representations of drugs and proteins.
-Specifically, the different projections of protein feature are used as key and value respectively while the projection of drug feature is used as query.
-TDGraphDTA[25]: TDGraphDTA use a multi-head cross-attention mechanism with two attention heads.
-Both drug and protein features are linearly transformed into query, key and value matrices.
-One cross attention layer uses a drug query matrix, a protein key matrix, and a protein value matrix, while its parallel counterparts use the rest of the matrices.
+CSDTI(pan2023csdti): CSDTI uses cross-attention to fuse the deep representations of drugs and proteins.
+Specifically, the different projections of protein features are used as keys and values, respectively, while the projection of drug features is used as a query.
+TDGraphDTA(zhu2023tdgraph): TDGraphDTA uses a multi-head cross-attention mechanism with two attention heads.
+Both drug and protein features are linearly transformed into query, key, and value matrices.
+One cross-attention layer uses a drug query matrix, a protein key matrix, and a protein value matrix, while its parallel counterparts use the rest of the matrices.
 The outputs of these two layers are concatenated and fed into MLP to get the final output.
 Appendix C Datasets Descriptions
-In this subsection, we provide a detailed description of the datasets for both the regression task and classification task. The statistical characteristics of the datasets are summarized in Table 4.
+In this subsection, we provide a detailed description of the datasets for both the regression task and classification task.
+The statistical characteristics of the datasets are summarized in Table 3.
+Here we present the statistics after we cleaned the data as described in Section 3.
 | Regression | Classification | |||||
@@
 | Number of target proteins | 379 | 229 | 1413 | 2001 | 1876 | 4256 |
-| Number of total samples | 25772 | 117657 | 52274 | 6728 | 7786 | 35021 |
+| Number of total samples | 25772 | 117657 | 52284 | 5997 | 6552 | 34740 |
 Appendix D Evaluation Metrics
 We adopt distinct sets of metrics to evaluate the classification and regression tasks.
-In particular, considering the classification task, we utilize the common metrics including Area Under Receiver Operating Characteristic Curve (ROC-AUC), Precision-Recall Area Under Curve (PR-AUC), LogAUC, accuracy, precision, recall, and F1 score.
+In particular, considering the classification task, we utilize the common metrics, including Area Under Receiver Operating Characteristic Curve (ROC-AUC), Precision-Recall Area Under Curve (PR-AUC), LogAUC, accuracy, precision, recall, and F1 score.
 For the continuous binding affinity regression, we benchmark the models using metrics of mean squared error (MSE), mean absolute error (MAE), coefficient of determination (R2), Pearson correlation coefficient, concordance index (CI), and Spearman correlation coefficient.
 Each of these metrics offers unique insights into different aspects of model performance, allowing us to assess predictive accuracy, correlation with observed values, and consistency in ranking predictions.
-Appendix E Greedy Search on Hyperparameter
-Here we present the detailed experiment on hyperparameters with full metrics.
-The detailed review of the hyperparameters associated with the selected models is in Table 5.
-The detailed greedy hyperparameter seraching results on Davis and Human datasets are in Table 6.
+Appendix E Original Hyperparameter
+To have a basic understanding of hyperparameters before greedy search and to find the optimized setting for each model, we summarize the hyperparameters reported in the corresponding paper or codes in Table 4.
 | Category | Models | Batch size | Total epoch | Learning rate & Decay & Decay epoch | Weight decay | Dropout | Optimizer |
-| GNN | GraphDTA-GCN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam |
-| GraphDTA-GAT [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| GraphDTA-GATGCN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| GraphDTA-GIN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| GraphCPI-GCN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| GraphCPI-GAT [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| GraphCPI-GATGCN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| GraphCPI-GIN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| MGraphDTA [59] | 512 | 292 | 0.0005 | - | 0.1 | Adam | |
-| SAGDTA [60] | 512 | 2000 | 0.001 | - | 0.1 | Adam | |
-| EmbedDTI [61] | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
-| DeepGLSTM [29] | 128 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| CPI [31] | 1 | 100 | {0.001, 0.5, 10} | 1e-6 | 0 | Adam | |
-| BACPI [57] | 16 | 20 | {0.0005, 0.5, 10} | - | 0.1 | Adam | |
-| DeepNC-HGC [30] | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| DeepNC-GEN [30] | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| DrugBAN [62] | 64 | 100 | 0.0005 | - | 0 | Adam | |
-| GANDTI [58] | 1 | 30/15 | 0.001 | 1e-6 | 0.5 | Adam | |
-| PGraphDTA-CNN [24] | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
-| BridgeDPI [63] | 512 | 100 | 0.001 | - | 0.5 | Adam | |
-| ColdDTA [64] | 128 | 292 | 0.003 | - | 0 | Adam | |
-| SubMDTA [36] | 512 | 1200 | 0.0005 | - | 0.2 | Adam | |
-| IMAEN [65] | 128 | 1000 | 0.0005 | - | 0.2 | Adam | |
-| Transformer | CSDTI [32] | 256 | 292 | 0.0005 | - | {0.5,0.2} | Adam |
-| AMMVF [38] | 32 | 40 | {0.001, 0.5, 5} | 1e-4 | 0.1 | Adam | |
-| TDGraphDTA [25] | 1024 | 3000 | 0.0005 | - | 0.1 | Adam | |
-| IIFDTI [66] | 64 | 100 | 0.001 | 1e-6 | 0.2 | AdamW | |
-| ICAN [42] | 128 | 50 | 0.001 | - | 0.1 | Adam | |
-| MolTrans [20] | 16 | 13 | 0.001 | - | 0.1 | Adam | |
-| TransformerCPI [21] | 8 | 40 | 0.0001 | 0.0001 | 0.2 | Radam | |
-| MRBDTA [55] | 256 | 300 | 0.001 | - | 0.1 | Adam | |
-| FOTFCPI [41] | 64 | 100 | 0.0001 | - | 0.1 | Adam |
-| Models | Hyper parameter | Regression | Classification | |||||||||||||
-| Batch size | Learning rate | Dropout | MSE | MAE | R2 | PCC | CI | Spearman | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | |
-| GraphDTA | 512 | 0.0005 | 0.2 | 0.2838 | 0.3053 | 0.5773 | 0.7619 | 0.8508 | 0.6226 | 0.9117 | 0.8807 | 0.4712 | 0.9117 | 0.9157 | 0.9125 | 0.9141 |
-| 256 | 0.0005 | 0.2 | 0.2781 | 0.2924 | 0.5859 | 0.7677 | 0.8523 | 0.6250 | 0.9132 | 0.8805 | 0.4584 | 0.9135 | 0.9115 | 0.9218 | 0.9165 | |
-| 128 | 0.0005 | 0.2 | 0.2634 | 0.2552 | 0.6077 | 0.7840 | 0.8700 | 0.6828 | 0.9134 | 0.8822 | 0.4716 | 0.9135 | 0.9157 | 0.9166 | 0.9160 | |
-| 64 | 0.0005 | 0.2 | 0.2591 | 0.2506 | 0.6141 | 0.7885 | 0.8634 | 0.6862 | 0.9121 | 0.8793 | 0.4608 | 0.9123 | 0.9112 | 0.9195 | 0.9153 | |
-| 512 | 0.0001 | 0.2 | 0.3209 | 0.3148 | 0.5221 | 0.7299 | 0.8369 | 0.6008 | 0.9165 | 0.8851 | 0.4775 | 0.9168 | 0.9159 | 0.9232 | 0.9195 | |
-| 512 | 0.001 | 0.2 | 0.2909 | 0.3076 | 0.5668 | 0.7812 | 0.8425 | 0.6073 | 0.9133 | 0.8827 | 0.4765 | 0.9134 | 0.9172 | 0.9145 | 0.9157 | |
-| 512 | 0.0005 | 0 | 0.2827 | 0.2923 | 0.5791 | 0.7632 | 0.8527 | 0.6267 | 0.9177 | 0.8863 | 0.4764 | 0.9180 | 0.9163 | 0.9252 | 0.9207 | |
-| 512 | 0.0005 | 0.1 | 0.2771 | 0.2947 | 0.5873 | 0.7695 | 0.8521 | 0.6241 | 0.9170 | 0.8873 | 0.4830 | 0.9171 | 0.9204 | 0.9189 | 0.9194 | |
-| 512 | 0.0005 | 0.5 | 0.2951 | 0.3165 | 0.5605 | 0.7494 | 0.8478 | 0.6176 | 0.9128 | 0.8819 | 0.4637 | 0.9129 | 0.9161 | 0.9151 | 0.9155 | |
-| GraphCPI | 512 | 0.0005 | 0.2 | 0.3173 | 0.3309 | 0.5275 | 0.7325 | 0.8267 | 0.5825 | 0.9049 | 0.8687 | 0.4275 | 0.9053 | 0.9005 | 0.9177 | 0.9089 |
-| 256 | 0.0005 | 0.2 | 0.2781 | 0.2924 | 0.5859 | 0.7677 | 0.8523 | 0.6250 | 0.9058 | 0.8712 | 0.4394 | 0.9061 | 0.9047 | 0.9140 | 0.9093 | |
-| 128 | 0.0005 | 0.2 | 0.3064 | 0.2988 | 0.5436 | 0.7402 | 0.8467 | 0.6213 | 0.9085 | 0.8762 | 0.4602 | 0.9086 | 0.9115 | 0.9111 | 0.9112 | |
-| 64 | 0.0005 | 0.2 | 0.3484 | 0.3569 | 0.4812 | 0.6944 | 0.8249 | 0.5850 | 0.9123 | 0.8774 | 0.4453 | 0.9128 | 0.9060 | 0.9270 | 0.9163 | |
-| 512 | 0.0001 | 0.2 | 0.3461 | 0.3494 | 0.4846 | 0.7075 | 0.8317 | 0.5927 | 0.9063 | 0.8691 | 0.4217 | 0.9068 | 0.8985 | 0.9235 | 0.9107 | |
-| 512 | 0.001 | 0.2 | 0.2870 | 0.2733 | 0.5726 | 0.7631 | 0.8598 | 0.6416 | 0.9064 | 0.8717 | 0.4438 | 0.9067 | 0.9048 | 0.9151 | 0.9099 | |
-| 512 | 0.0005 | 0 | 0.3431 | 0.3527 | 0.4891 | 0.7096 | 0.8297 | 0.5890 | 0.9060 | 0.8709 | 0.4386 | 0.9064 | 0.9036 | 0.9160 | 0.9097 | |
-| 512 | 0.0005 | 0.1 | 0.3291 | 0.3388 | 0.5100 | 0.7265 | 0.8294 | 0.5885 | 0.9060 | 0.8706 | 0.4385 | 0.9064 | 0.9029 | 0.9169 | 0.9098 | |
-| 512 | 0.0005 | 0.5 | 0.3329 | 0.3264 | 0.5043 | 0.7304 | 0.8361 | 0.6211 | 0.9064 | 0.8702 | 0.4286 | 0.9068 | 0.9010 | 0.9203 | 0.9105 | |
-| MRBDTA | 128 | 0.001 | 0.1 | 0.2397 | 0.2648 | 0.6430 | 0.8025 | 0.8716 | 0.6546 | 0.9340 | 0.9119 | 0.5240 | 0.9338 | 0.9427 | 0.9281 | 0.9353 |
-| 64 | 0.001 | 0.1 | 0.2504 | 0.2679 | 0.6272 | 0.7935 | 0.8674 | 0.6483 | 0.9348 | 0.9113 | 0.5219 | 0.9348 | 0.9396 | 0.9336 | 0.9365 | |
-| 32 | 0.001 | 0.1 | 0.2567 | 0.2721 | 0.6177 | 0.7878 | 0.8622 | 0.6400 | 0.9328 | 0.9127 | 0.5331 | 0.9324 | 0.9474 | 0.9200 | 0.9334 | |
-| 128 | 0.0005 | 0.1 | 0.2350 | 0.2565 | 0.6499 | 0.8069 | 0.8775 | 0.6639 | 0.9382 | 0.9164 | 0.5304 | 0.9381 | 0.9442 | 0.9353 | 0.9397 | |
-| 128 | 0.0001 | 0.1 | 0.2335 | 0.2696 | 0.6523 | 0.8084 | 0.8767 | 0.6630 | 0.9372 | 0.9171 | 0.5350 | 0.9370 | 0.9481 | 0.9284 | 0.9381 | |
-| 128 | 0.001 | 0 | 0.2425 | 0.2663 | 0.6388 | 0.7997 | 0.8739 | 0.6587 | 0.9279 | 0.9010 | 0.4997 | 0.9280 | 0.9294 | 0.9316 | 0.9303 | |
-| 128 | 0.001 | 0.2 | 0.2408 | 0.2620 | 0.6414 | 0.8027 | 0.8679 | 0.6478 | 0.9438 | 0.9218 | 0.5363 | 0.9438 | 0.9450 | 0.9460 | 0.9454 | |
-| 128 | 0.001 | 0.5 | 0.2584 | 0.2666 | 0.6153 | 0.7894 | 0.8727 | 0.6573 | 0.9328 | 0.9123 | 0.5295 | 0.9325 | 0.9462 | 0.9215 | 0.9336 | |
-| TransformerCPI | 128 | 0.001 | 0.1 | 0.3747 | 0.4111 | 0.4420 | 0.6787 | 0.8164 | 0.5668 | 0.9294 | 0.9024 | 0.5047 | 0.9296 | 0.9301 | 0.9339 | 0.9317 |
-| 64 | 0.001 | 0.1 | 0.5916 | 0.5422 | 0.1191 | 0.3531 | 0.6644 | 0.3014 | 0.9318 | 0.9061 | 0.5100 | 0.9319 | 0.9337 | 0.9342 | 0.9339 | |
-| 32 | 0.001 | 0.1 | 0.6219 | 0.5197 | 0.0740 | 0.2993 | 0.6360 | 0.2506 | 0.9269 | 0.8959 | 0.4860 | 0.9273 | 0.9198 | 0.9411 | 0.9302 | |
-| 128 | 0.0005 | 0.1 | 0.2869 | 0.3389 | 0.5728 | 0.7750 | 0.8326 | 0.5910 | 0.9270 | 0.9008 | 0.5030 | 0.9270 | 0.9312 | 0.9270 | 0.9290 | |
-| 128 | 0.0001 | 0.1 | 0.2877 | 0.3411 | 0.5716 | 0.7674 | 0.8348 | 0.5950 | 0.9208 | 0.8898 | 0.4664 | 0.9211 | 0.9181 | 0.9307 | 0.9239 | |
-| 128 | 0.001 | 0 | 0.3874 | 0.3948 | 0.4232 | 0.6481 | 0.7985 | 0.5358 | 0.9297 | 0.9015 | 0.4999 | 0.9300 | 0.9272 | 0.9376 | 0.9324 | |
-| 128 | 0.001 | 0.2 | 0.5181 | 0.5219 | 0.2284 | 0.5249 | 0.7476 | 0.4494 | 0.9236 | 0.8938 | 0.4780 | 0.9239 | 0.9217 | 0.9319 | 0.9265 | |
-| 128 | 0.001 | 0.5 | 1.5318 | 1.1598 | -1.2809 | 0.2866 | 0.6441 | 0.2647 | 0.9218 | 0.8895 | 0.4746 | 0.9223 | 0.9153 | 0.9359 | 0.9253 | |
-Appendix F Comparison of different featurization
+| GNN | GraphDTA-GCN (nguyen2021graphdta) | 512 | 1000 | 0.0005 | - | 0.2 | Adam |
+| GraphDTA-GAT (nguyen2021graphdta) | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
+| GraphDTA-GATGCN (nguyen2021graphdta) | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
+| GraphDTA-GIN (nguyen2021graphdta) | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
+| GraphCPI-GCN (quan2019graphcpi) | 512 | 1000 | 0.0005 | - | 0.5 | Adam | |
+| GraphCPI-GAT (quan2019graphcpi) | 512 | 1000 | 0.0005 | - | 0.6 | Adam | |
+| GraphCPI-GATGCN (quan2019graphcpi) | 512 | 1000 | 0.0005 | - | - | Adam | |
+| GraphCPI-GIN (quan2019graphcpi) | 512 | 1000 | 0.0005 | - | 0.6 | Adam | |
+| MGraphDTA (yang2022MGraphDTA) | 512 | 3000 | 0.0005 | - | 0.1 | Adam | |
+| SAGDTA (zhang2021sag) | 512 | 2000 | 0.001 | - | 0.1 | Adam | |
+| EmbedDTI (jin2021embeddti) | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
+| DeepGLSTM (mukherjee2022deepglstm) | 512/128 | 1000 | 0.0005 | - | 0.2 | Adam | |
+| CPI (tsubaki2019compound) | 1 | 100 | 0.001, 0.5, 10 | 1e-6 | 0 | Adam | |
+| BACPI (li2022bacpi) | 16 | 20 | 0.0005, 0.5, 10 | - | 0.1 | Adam | |
+| DeepNC-HGC (tran2022deepnc) | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
+| DeepNC-GEN (tran2022deepnc) | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
+| DrugBAN (bai2023interpretable-drugban) | 64 | 100 | 0.00005 | - | 0 | Adam | |
+| GANDTI (wang2021gandti) | 1 | 30/15 | 0.001 | 1e-6 | 0.5 | Adam | |
+| PGraphDTA-CNN (bal2024pgraphdta) | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
+| BridgeDPI (wu2022bridgedpi) | 512 | 100 | 0.001 | - | 0.5 | Adam | |
+| ColdDTA (fang2023colddta) | 128 | 700/300 | 0.0003 | - | 0 | Adam | |
+| SubMDTA (pan2023submdta) | 512 | 1200 | 0.0005 | - | 0.2 | Adam | |
+| IMAEN (zhang2024imaen) | 128 | 1000 | 0.0005 | - | 0.2 | Adam | |
+| Transformer | CSDTI (pan2023csdti) | 256 | 3000 | 0.0005 | - | 0.2 | Adam |
+| AMMVF (wang2023ammvf) | 32 | 40 | 0.001, 0.5, 5 | 1e-4 | 0.1 | Adam | |
+| TDGraphDTA (zhu2023tdgraph) | 1024 | 3000 | 0.0005 | - | 0.1 | Adam | |
+| IIFDTI (cheng2022iifdti) | 64 | 200 | 0.001 | 1e-6 | 0.2 | AdamW | |
+| ICAN (kurata2022ican) | 128 | 50 | 0.001 | - | 0.1 | Adam | |
+| MolTrans (huang2021moltrans) | 64 | 30 | 0.00001 | - | 0.1 | Adam | |
+| TransformerCPI (chen2020Transformercpi) | 8 | 40 | 0.0001, 0.5, 5 | 1e-4 | 0.2 | RAdam | |
+| MRBDTA (zhang2022predicting) | 1024/256 | 600/300 | 0.001 | - | 0.1 | Adam | |
+| FOTFCPI (yin2024fotf) | 64 | 100 | 0.0001 | - | 0.1 | Adam |
+Appendix F Full experiment
+To evaluate the model’s best performance, based on the hyperparameters given in its paper or codes, we found the optimized hyperparameters for each model.
+On top of the mean value, we also provide the standard deviation across five-fold.
+The complete result on the regression task is shown in Table 5, and the complete result on the classification task is shown in Table 6.
+| Category | Models | DAVIS | KIBA | BindingDB | |||||||||||||||
+| MSE | MAE | R2 | PCC | CI | Spearman | MSE () | MAE | R2 | PCC | CI | Spearman | MSE | MAE | R2 | PCC | CI | Spearman | ||
+| GNN | GraphDTA-GCN | ||||||||||||||||||
+| GraphDTA-GAT | |||||||||||||||||||
+| GraphDTA-GATGCN | |||||||||||||||||||
+| GraphDTA-GIN | |||||||||||||||||||
+| GraphCPI-GCN | |||||||||||||||||||
+| GraphCPI-GAT | |||||||||||||||||||
+| GraphCPI-GATGCN | |||||||||||||||||||
+| GraphCPI-GIN | |||||||||||||||||||
+| MGraphDTA | |||||||||||||||||||
+| SAGDTA | |||||||||||||||||||
+| EmbedDTI | |||||||||||||||||||
+| DeepGLSTM | |||||||||||||||||||
+| CPI | |||||||||||||||||||
+| BACPI | |||||||||||||||||||
+| DeepNC-HGC | |||||||||||||||||||
+| DeepNC-GEN | |||||||||||||||||||
+| DrugBAN | |||||||||||||||||||
+| GANDTI | |||||||||||||||||||
+| BridgeDPI | |||||||||||||||||||
+| ColdDTA | |||||||||||||||||||
+| SubMDTA | |||||||||||||||||||
+| IMAEN | |||||||||||||||||||
+| Transformer | CSDTI | ||||||||||||||||||
+| TDGraphDTA | |||||||||||||||||||
+| AMMVF | |||||||||||||||||||
+| IIFDTI | |||||||||||||||||||
+| ICAN | - | - | |||||||||||||||||
+| MolTrans | |||||||||||||||||||
+| TransformerCPI | |||||||||||||||||||
+| MRBDTA | |||||||||||||||||||
+| FOTFCPI | |||||||||||||||||||
+| Our combos | |||||||||||||||||||
+| Categories | Models | Human | C.elegans | Drugbank | ||||||||||||||||||
+| ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ||
+| GNN | GraphDTA-GCN | |||||||||||||||||||||
+| GraphDTA-GAT | ||||||||||||||||||||||
+| GraphDTA-GATGCN | ||||||||||||||||||||||
+| GraphDTA-GIN | ||||||||||||||||||||||
+| GraphCPI-GCN | ||||||||||||||||||||||
+| GraphCPI-GAT | ||||||||||||||||||||||
+| GraphCPI-GATGCN | ||||||||||||||||||||||
+| GraphCPI-GIN | ||||||||||||||||||||||
+| MGraphDTA | ||||||||||||||||||||||
+| SAGDTA | ||||||||||||||||||||||
+| EmbedDTI | ||||||||||||||||||||||
+| DeepGLSTM | ||||||||||||||||||||||
+| CPI | ||||||||||||||||||||||
+| BACPI | ||||||||||||||||||||||
+| DeepNC-HGC | ||||||||||||||||||||||
+| DeepNC-GEN | ||||||||||||||||||||||
+| DrugBAN | ||||||||||||||||||||||
+| GANDTI | ||||||||||||||||||||||
+| BridgeDPI | ||||||||||||||||||||||
+| ColdDTA | ||||||||||||||||||||||
+| SubMDTA | ||||||||||||||||||||||
+| IMAEN | ||||||||||||||||||||||
+| Transformer | CSDTI | |||||||||||||||||||||
+| TDGraphDTA | ||||||||||||||||||||||
+| AMMVF | ||||||||||||||||||||||
+| IIFDTI | ||||||||||||||||||||||
+| ICAN | ||||||||||||||||||||||
+| MolTrans | ||||||||||||||||||||||
+| TransformerCPI | ||||||||||||||||||||||
+| MRBDTA | ||||||||||||||||||||||
+| FOTFCPI | ||||||||||||||||||||||
+| Our combos | ||||||||||||||||||||||
+Appendix G Comparison of different featurization
 In this section, we present the summarized featurization methods in Table 7, the detailed description of all properties is shown in Table 8.
@@
 | Basic+AP+HI+EP+Ste+Str | 0.3469 | 0.3550 | 0.4834 | 0.7073 | 0.8228 | 0.5775 | 0.9134 | 0.8772 | 0.4440 | 0.9140 | 0.9033 | 0.9328 | 0.9178 | |
-Appendix G Full experiment
-The complete result on the regression task is shown in Table 10, and the complete result on classification task is shown in Table 11. All experiments are run on the RTX 3090 with more than 1000 hours.
-| Category | Models | DAVIS | KIBA | BindingDB | |||||||||||||||
-| MSE | MAE | R2 | PCC | CI | Spearman | MSE | MAE | R2 | PCC | CI | Spearman | MSE | MAE | R2 | PCC | CI | Spearman | ||
-| GNN | GraphDTA-GCN | 0.2771 | 0.2947 | 0.5873 | 0.7695 | 0.8521 | 0.6241 | 0.0005 | 0.0155 | 0.3924 | 0.6291 | 0.7502 | 0.6448 | 0.5033 | 0.4314 | 0.7259 | 0.8536 | 0.8576 | 0.7795 |
-| GraphDTA-GAT | 0.2806 | 0.2928 | 0.5820 | 0.7856 | 0.8576 | 0.6331 | 0.0004 | 0.0123 | 0.6012 | 0.7720 | 0.7977 | 0.7259 | 0.5613 | 0.4460 | 0.6943 | 0.8358 | 0.8496 | 0.7662 | |
-| GraphDTA-GATGCN | 0.2570 | 0.2798 | 0.6173 | 0.7867 | 0.8607 | 0.6380 | 0.0004 | 0.0126 | 0.6128 | 0.7836 | 0.8086 | 0.7357 | 0.5447 | 0.4450 | 0.7033 | 0.8415 | 0.8489 | 0.7641 | |
-| GraphDTA-GIN | 0.2309 | 0.2714 | 0.6562 | 0.8105 | 0.8711 | 0.6540 | 0.0004 | 0.0131 | 0.5544 | 0.7451 | 0.7996 | 0.7400 | 0.5033 | 0.4314 | 0.7259 | 0.8537 | 0.8576 | 0.7795 | |
-| GraphCPI-GCN | 0.3291 | 0.3388 | 0.5100 | 0.7265 | 0.8294 | 0.5885 | 0.0006 | 0.0156 | 0.3454 | 0.5864 | 0.7378 | 0.6154 | 0.6372 | 0.5002 | 0.6530 | 0.8133 | 0.8329 | 0.7360 | |
-| GraphCPI-GAT | 0.3649 | 0.3852 | 0.4566 | 0.7609 | 0.8498 | 0.6215 | 0.0007 | 0.0179 | 0.1836 | 0.4196 | 0.6845 | 0.4896 | 0.6565 | 0.5007 | 0.6424 | 0.8163 | 0.8350 | 0.7407 | |
-| GraphCPI-GATGCN | 0.3078 | 0.3193 | 0.5416 | 0.7397 | 0.8365 | 0.5995 | 0.0005 | 0.0156 | 0.4062 | 0.6421 | 0.7530 | 0.6522 | 0.6023 | 0.4858 | 0.6719 | 0.8242 | 0.8370 | 0.7442 | |
-| GraphCPI-GIN | 0.2413 | 0.2848 | 0.6407 | 0.8020 | 0.8671 | 0.6478 | 0.0005 | 0.0138 | 0.4963 | 0.7093 | 0.7808 | 0.7049 | 0.5069 | 0.4247 | 0.7239 | 0.8531 | 0.8572 | 0.7784 | |
-| MGraphDTA | 0.2179 | 0.2355 | 0.6755 | 0.8239 | 0.8820 | 0.6704 | 0.0003 | 0.0090 | 0.7208 | 0.8508 | 0.8649 | 0.8439 | 0.4887 | 0.3852 | 0.7338 | 0.8586 | 0.8649 | 0.7897 | |
-| SAGDTA | 0.2656 | 0.2796 | 0.6045 | 0.7805 | 0.8675 | 0.6498 | 0.0039 | 0.0127 | -3.3777 | 0.5064 | 0.8096 | 0.7547 | 0.6590 | 0.4709 | 0.6410 | 0.8061 | 0.8366 | 0.7418 | |
-| EmbedDTI | 0.2561 | 0.2798 | 0.6186 | 0.7874 | 0.8624 | 0.6410 | 0.0007 | 0.0175 | 0.2364 | 0.6864 | 0.6374 | 0.6897 | 0.5095 | 0.4193 | 0.7225 | 0.8516 | 0.8559 | 0.7765 | |
-| DeepGLSTM | 0.2915 | 0.2941 | 0.5659 | 0.7605 | 0.8476 | 0.6176 | 0.0003 | 0.0101 | 0.6919 | 0.8325 | 0.8480 | 0.8231 | 0.5385 | 0.4246 | 0.7067 | 0.8433 | 0.8529 | 0.7713 | |
-| CPI | 0.3503 | 0.3428 | 0.4784 | 0.6692 | 0.8319 | 0.5873 | 0.0003 | 0.0117 | 0.6274 | 0.7926 | 0.8190 | 0.7767 | 0.6962 | 0.5244 | 0.6208 | 0.7900 | 0.8261 | 0.7252 | |
-| BACPI | 0.4036 | 0.3534 | 0.3990 | 0.6518 | 0.7982 | 0.5373 | 0.0006 | 0.0122 | 0.5320 | 0.7338 | 0.8175 | 0.7719 | 0.6468 | 0.4805 | 0.6477 | 0.8109 | 0.8297 | 0.7303 | |
-| DeepNC-HGC | 0.2782 | 0.3010 | 0.5857 | 0.7669 | 0.8551 | 0.6297 | 0.0005 | 0.0141 | 0.4983 | 0.7127 | 0.7823 | 0.7052 | 0.5611 | 0.4512 | 0.6944 | 0.8367 | 0.8464 | 0.7600 | |
-| DeepNC-GEN | 0.2543 | 0.2830 | 0.6213 | 0.7893 | 0.8634 | 0.6419 | 0.0004 | 0.0132 | 0.5355 | 0.7447 | 0.7981 | 0.7349 | 0.5561 | 0.4350 | 0.6971 | 0.8395 | 0.8496 | 0.7629 | |
-| DrugBAN | 0.2391 | 0.2663 | 0.6440 | 0.8035 | 0.8757 | 0.6619 | 0.0004 | 0.0104 | 0.5611 | 0.7665 | 0.8388 | 0.8081 | 0.4485 | 0.3945 | 0.7557 | 0.8696 | 0.8693 | 0.7989 | |
-| GANDTI | 0.3082 | 0.3305 | 0.5410 | 0.7391 | 0.8414 | 0.6079 | 0.0003 | 0.0108 | 0.6569 | 0.8146 | 0.8342 | 0.8045 | 0.6714 | 0.5221 | 0.6343 | 0.7999 | 0.8322 | 0.7342 | |
-| PGraphDTA-CNN | 0.3273 | 0.3891 | 0.5126 | 0.7929 | 0.8701 | 0.6542 | 0.0005 | 0.0144 | 0.4833 | 0.6928 | 0.7473 | 0.6217 | 0.5334 | 0.4365 | 0.7095 | 0.8500 | 0.8590 | 0.7833 | |
-| BridgeDPI | 0.3623 | 0.3432 | 0.6477 | 0.8062 | 0.8991 | 0.6704 | 0.0004 | 0.0133 | 0.5686 | 0.7357 | 0.7849 | 0.7239 | 0.4482 | 0.3900 | 0.7559 | 0.8702 | 0.8698 | 0.7982 | |
-| ColdDTA | 0.2346 | 0.2511 | 0.6507 | 0.8083 | 0.8693 | 0.6501 | 0.0004 | 0.0129 | 0.5948 | 0.7764 | 0.8018 | 0.7451 | 0.4697 | 0.3783 | 0.7442 | 0.8645 | 0.8644 | 0.7866 | |
-| SubMDTA | 0.2326 | 0.2726 | 0.6537 | 0.8091 | 0.8691 | 0.6503 | 0.0003 | 0.0098 | 0.6855 | 0.8324 | 0.8485 | 0.8243 | 0.4566 | 0.3995 | 0.7513 | 0.8677 | 0.8670 | 0.7953 | |
-| IMAEN | 0.2412 | 0.2764 | 0.6409 | 0.8029 | 0.8721 | 0.6557 | 0.0004 | 0.0122 | 0.5800 | 0.7632 | 0.8061 | 0.7494 | 0.4720 | 0.3935 | 0.7429 | 0.8637 | 0.8553 | 0.7938 | |
-| Transformer | CSDTI | 0.3029 | 0.3011 | 0.5490 | 0.7436 | 0.8395 | 0.6045 | 0.0007 | 0.0184 | 0.2448 | 0.4504 | 0.6475 | 0.3937 | 0.6408 | 0.4723 | 0.6510 | 0.8118 | 0.8369 | 0.7422 |
-| TDGraphDTA | 0.2217 | 0.2399 | 0.6698 | 0.8201 | 0.6685 | 0.8804 | 0.0008 | 0.0292 | 0.0533 | 0.2604 | 0.3429 | 0.6243 | 0.4750 | 0.3894 | 0.7413 | 0.8631 | 0.7887 | 0.8642 | |
-| AMMVF | 0.3325 | 0.3433 | 0.5048 | 0.7238 | 0.8307 | 0.5896 | 0.0006 | 0.0147 | 0.4696 | 0.6957 | 0.7711 | 0.6879 | 0.6597 | 0.4879 | 0.6407 | 0.8086 | 0.8336 | 0.7398 | |
-| IIFDTI | 0.2741 | 0.3006 | 0.5918 | 0.7816 | 0.8500 | 0.6202 | 0.0005 | 0.0156 | 0.2584 | 0.7894 | 0.7952 | 0.7439 | 0.5097 | 0.4474 | 0.7170 | 0.8576 | 0.8585 | 0.7792 | |
-| ICAN | 0.3481 | 0.3185 | 0.4816 | 0.7168 | 0.8211 | 0.5822 | 0.0008 | 0.0194 | 0.1159 | 0.7646 | 0.8256 | 0.7877 | 0.6582 | 0.5028 | 0.6415 | 0.8145 | 0.8277 | 0.7271 | |
-| MolTrans | 0.2588 | 0.2740 | 0.6146 | 0.7906 | 0.8601 | 0.5927 | 0.0003 | 0.0114 | 0.6378 | 0.8064 | 0.8453 | 0.7865 | 0.5138 | 0.4101 | 0.7201 | 0.8517 | 0.8570 | 0.7786 | |
-| TransformerCPI | 0.2869 | 0.3389 | 0.5728 | 0.7750 | 0.8326 | 0.5910 | 0.0008 | 0.0233 | 0.0728 | 0.8051 | 0.8357 | 0.7809 | 0.5704 | 0.4624 | 0.6894 | 0.8394 | 0.8426 | 0.7507 | |
-| MRBDTA | 0.2350 | 0.2565 | 0.6499 | 0.8069 | 0.8775 | 0.6639 | 0.0006 | 0.0171 | 0.3210 | 0.5718 | 0.7239 | 0.5878 | 0.4977 | 0.4133 | 0.7289 | 0.8557 | 0.8629 | 0.7874 | |
-| FOTFCPI | 0.2803 | 0.3004 | 0.5825 | 0.7704 | 0.8546 | 0.6286 | 0.0004 | 0.0133 | 0.5342 | 0.7552 | 0.7947 | 0.7353 | 0.5743 | 0.4472 | 0.6872 | 0.8322 | 0.8444 | 0.7569 | |
-| our | 0.2063 | 0.2481 | 0.6927 | 0.8330 | 0.8901 | 0.6839 | 0.0003 | 0.0094 | 0.7168 | 0.8512 | 0.8677 | 0.8432 | 0.4651 | 0.3878 | 0.7467 | 0.8657 | 0.8683 | 0.7956 | |
-| Categories | Models | Human | C.elegans | Drugbank | ||||||||||||||||||
-| ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | ||
-| GNN | GraphDTA-GCN | 0.9222 | 0.8922 | 0.4852 | 0.9224 | 0.9210 | 0.9293 | 0.9251 | 0.9488 | 0.9174 | 0.5342 | 0.9468 | 0.9368 | 0.9564 | 0.9465 | 0.7590 | 0.6925 | 0.1490 | 0.7589 | 0.7458 | 0.7842 | 0.7645 |
-| GraphDTA-GAT | 0.8935 | 0.8557 | 0.3986 | 0.8937 | 0.8925 | 0.9024 | 0.8974 | 0.9289 | 0.8909 | 0.4942 | 0.9287 | 0.9154 | 0.9423 | 0.9286 | 0.7684 | 0.7030 | 0.1595 | 0.7683 | 0.7587 | 0.7855 | 0.7718 | |
-| GraphDTA-GATGCN | 0.9296 | 0.9024 | 0.5030 | 0.9297 | 0.9298 | 0.9342 | 0.9319 | 0.9487 | 0.9176 | 0.5318 | 0.9484 | 0.9335 | 0.9640 | 0.9485 | 0.7712 | 0.7046 | 0.1585 | 0.7711 | 0.7573 | 0.7968 | 0.7764 | |
-| GraphDTA-GIN | 0.9019 | 0.8674 | 0.4393 | 0.9021 | 0.9037 | 0.9065 | 0.9051 | 0.9470 | 0.9174 | 0.5342 | 0.9468 | 0.9368 | 0.9564 | 0.9465 | 0.7871 | 0.7232 | 0.1797 | 0.7871 | 0.7797 | 0.7988 | 0.7892 | |
-| GraphCPI-GCN | 0.9034 | 0.8649 | 0.4048 | 0.9040 | 0.8941 | 0.9229 | 0.9083 | 0.9322 | 0.8940 | 0.4967 | 0.9319 | 0.9155 | 0.9493 | 0.9320 | 0.7362 | 0.6688 | 0.1292 | 0.7361 | 0.7178 | 0.7764 | 0.7459 | |
-| GraphCPI-GAT | 0.8935 | 0.8557 | 0.3986 | 0.8937 | 0.8925 | 0.9024 | 0.8974 | 0.9281 | 0.8913 | 0.4965 | 0.9279 | 0.9181 | 0.9373 | 0.9275 | 0.7515 | 0.6862 | 0.1470 | 0.7515 | 0.7433 | 0.7667 | 0.7548 | |
-| GraphCPI-GATGCN | 0.9097 | 0.8763 | 0.4644 | 0.9099 | 0.9090 | 0.9169 | 0.9129 | 0.9372 | 0.9028 | 0.5109 | 0.9370 | 0.9245 | 0.9496 | 0.9369 | 0.7561 | 0.6909 | 0.1508 | 0.7561 | 0.7479 | 0.7711 | 0.7593 | |
-| GraphCPI-GIN | 0.8870 | 0.8484 | 0.3791 | 0.8872 | 0.8881 | 0.8935 | 0.8908 | 0.9407 | 0.9072 | 0.5173 | 0.9405 | 0.9270 | 0.9543 | 0.9404 | 0.7866 | 0.7241 | 0.1845 | 0.7866 | 0.7840 | 0.7903 | 0.7870 | |
-| MGraphDTA | 0.9408 | 0.9166 | 0.5244 | 0.9410 | 0.9393 | 0.9466 | 0.9429 | 0.9631 | 0.9407 | 0.5795 | 0.9630 | 0.9535 | 0.9723 | 0.9628 | 0.8146 | 0.7539 | 0.2154 | 0.8146 | 0.8090 | 0.8228 | 0.8157 | |
-| SAGDTA | 0.9021 | 0.8728 | 0.4631 | 0.9018 | 0.9172 | 0.8897 | 0.9032 | 0.9380 | 0.9050 | 0.5165 | 0.9378 | 0.9281 | 0.9472 | 0.9375 | 0.7655 | 0.7018 | 0.1631 | 0.7655 | 0.7624 | 0.7700 | 0.7662 | |
-| EmbedDTI | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.7625 | 0.6963 | 0.1526 | 0.7625 | 0.7504 | 0.7852 | 0.7674 | |
-| DeepGLSTM | 0.9180 | 0.8857 | 0.4758 | 0.9183 | 0.9145 | 0.9281 | 0.9212 | 0.9414 | 0.9103 | 0.5257 | 0.9412 | 0.9325 | 0.9496 | 0.9409 | 0.7640 | 0.6988 | 0.1567 | 0.7640 | 0.7551 | 0.7801 | 0.7674 | |
-| CPI | 0.9110 | 0.8789 | 0.4638 | 0.9111 | 0.9126 | 0.9152 | 0.9138 | 0.9312 | 0.8940 | 0.4983 | 0.9310 | 0.9176 | 0.9446 | 0.9309 | 0.7467 | 0.6812 | 0.1434 | 0.7466 | 0.7368 | 0.7683 | 0.7515 | |
-| BACPI | 0.9249 | 0.8958 | 0.4924 | 0.9251 | 0.9239 | 0.9313 | 0.9276 | 0.9556 | 0.9314 | 0.5637 | 0.9555 | 0.9493 | 0.9610 | 0.9551 | 0.7748 | 0.7100 | 0.1664 | 0.7748 | 0.7663 | 0.7895 | 0.7776 | |
-| DeepNC-HGC | 0.8796 | 0.8399 | 0.3579 | 0.8798 | 0.8823 | 0.8845 | 0.8834 | 0.9418 | 0.9106 | 0.5251 | 0.9417 | 0.9325 | 0.9504 | 0.9413 | 0.7673 | 0.7013 | 0.1571 | 0.7673 | 0.7556 | 0.7889 | 0.7718 | |
-| DeepNC-GEN | 0.9178 | 0.8873 | 0.4809 | 0.9180 | 0.9186 | 0.9224 | 0.9205 | 0.9501 | 0.9203 | 0.5385 | 0.9499 | 0.9366 | 0.9637 | 0.9498 | 0.7464 | 0.6785 | 0.1357 | 0.7463 | 0.7262 | 0.7928 | 0.7571 | |
-| DrugBAN | 0.9302 | 0.9048 | 0.5083 | 0.9302 | 0.9342 | 0.9299 | 0.9320 | 0.9596 | 0.9346 | 0.5645 | 0.9594 | 0.9478 | 0.9710 | 0.9593 | 0.8188 | 0.7610 | 0.2325 | 0.8188 | 0.8203 | 0.8157 | 0.8179 | |
-| GANDTI | 0.9333 | 0.9090 | 0.5162 | 0.9333 | 0.9374 | 0.9327 | 0.9351 | 0.9404 | 0.9082 | 0.5213 | 0.9403 | 0.9300 | 0.9504 | 0.9400 | 0.7427 | 0.6737 | 0.1298 | 0.7425 | 0.7188 | 0.7954 | 0.7550 | |
-| PGraphDTA-CNN | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.7910 | 0.7260 | 0.1787 | 0.7910 | 0.7787 | 0.8117 | 0.7948 | |
-| BridgeDPI | 0.9456 | 0.9250 | 0.5435 | 0.9456 | 0.9485 | 0.9457 | 0.9471 | 0.9651 | 0.9470 | 0.6054 | 0.9651 | 0.9629 | 0.9663 | 0.9646 | 0.7825 | 0.7181 | 0.1742 | 0.7825 | 0.7739 | 0.7974 | 0.7853 | |
-| ColdDTA | 0.9420 | 0.9203 | 0.5348 | 0.9421 | 0.9453 | 0.9420 | 0.9436 | 0.9623 | 0.9403 | 0.5785 | 0.9622 | 0.9543 | 0.9697 | 0.9619 | 0.8186 | 0.7578 | 0.2175 | 0.8186 | 0.8108 | 0.8300 | 0.8203 | |
-| SubMDTA | 0.9326 | 0.9054 | 0.5066 | 0.9328 | 0.9304 | 0.9400 | 0.9351 | 0.9570 | 0.9313 | 0.5589 | 0.9568 | 0.9461 | 0.9674 | 0.9566 | 0.8045 | 0.7414 | 0.1968 | 0.8045 | 0.7946 | 0.8204 | 0.8072 | |
-| IMAEN | 0.9058 | 0.8719 | 0.4489 | 0.9059 | 0.9067 | 0.9111 | 0.9089 | 0.9506 | 0.9255 | 0.5546 | 0.9505 | 0.9469 | 0.9530 | 0.9500 | 0.7995 | 0.7365 | 0.1931 | 0.7994 | 0.7918 | 0.8113 | 0.8014 | |
-| Transformer | CSDTI | 0.8630 | 0.8216 | 0.3223 | 0.8630 | 0.8707 | 0.8620 | 0.8663 | 0.8962 | 0.8446 | 0.3710 | 0.8958 | 0.8759 | 0.9183 | 0.8966 | 0.7269 | 0.6626 | 0.1304 | 0.7269 | 0.7196 | 0.7420 | 0.7306 |
-| TDGraphDTA | 0.9411 | 0.9226 | 0.5461 | 0.9409 | 0.9525 | 0.9316 | 0.9419 | 0.9573 | 0.9315 | 0.5591 | 0.9571 | 0.9460 | 0.9681 | 0.9569 | 0.8083 | 0.7448 | 0.1982 | 0.8082 | 0.7959 | 0.8281 | 0.8116 | |
-| AMMVF | 0.9287 | 0.9004 | 0.4994 | 0.9290 | 0.9268 | 0.9362 | 0.9314 | 0.9636 | 0.9426 | 0.5892 | 0.9635 | 0.9567 | 0.9700 | 0.9632 | 0.7814 | 0.7142 | 0.1717 | 0.7814 | 0.7728 | 0.7973 | 0.7849 | |
-| IIFDTI | 0.9392 | 0.9163 | 0.5279 | 0.9392 | 0.9420 | 0.9400 | 0.9409 | 0.9679 | 0.9500 | 0.6054 | 0.9679 | 0.9633 | 0.9718 | 0.9675 | 0.8084 | 0.7441 | 0.1950 | 0.8084 | 0.7934 | 0.8327 | 0.8125 | |
-| ICAN | 0.9387 | 0.9134 | 0.5194 | 0.9389 | 0.9361 | 0.9460 | 0.9410 | 0.9644 | 0.9479 | 0.6146 | 0.9644 | 0.9662 | 0.9613 | 0.9637 | 0.7731 | 0.7092 | 0.1682 | 0.7731 | 0.7681 | 0.7812 | 0.7745 | |
-| MolTrans | 0.9453 | 0.9233 | 0.5377 | 0.9455 | 0.9451 | 0.9492 | 0.9471 | 0.9639 | 0.9428 | 0.5847 | 0.9638 | 0.9565 | 0.9705 | 0.9634 | 0.7910 | 0.7283 | 0.1879 | 0.7909 | 0.7861 | 0.7988 | 0.7922 | |
-| TransformerCPI | 0.9311 | 0.9045 | 0.5060 | 0.9312 | 0.9316 | 0.9351 | 0.9333 | 0.9581 | 0.9343 | 0.5696 | 0.9580 | 0.9501 | 0.9655 | 0.9577 | 0.8097 | 0.7470 | 0.2024 | 0.8097 | 0.7992 | 0.8265 | 0.8125 | |
-| MRBDTA | 0.9447 | 0.9258 | 0.5521 | 0.9446 | 0.9523 | 0.9397 | 0.9458 | 0.9713 | 0.9569 | 0.6335 | 0.9713 | 0.9708 | 0.9710 | 0.9709 | 0.8102 | 0.7469 | 0.2003 | 0.8102 | 0.7978 | 0.8298 | 0.8134 | |
-| FOTFCPI | 0.9444 | 0.9235 | 0.5413 | 0.9444 | 0.9477 | 0.9443 | 0.9459 | 0.9673 | 0.9486 | 0.6009 | 0.9672 | 0.9618 | 0.9721 | 0.9669 | 0.7845 | 0.7188 | 0.1713 | 0.7845 | 0.7713 | 0.8075 | 0.7889 | |
-| our | 0.9435 | 0.9221 | 0.5400 | 0.9435 | 0.9462 | 0.9443 | 0.9451 | 0.9688 | 0.9518 | 0.6178 | 0.9688 | 0.9652 | 0.9718 | 0.9684 | 0.8035 | 0.7405 | 0.1961 | 0.8035 | 0.7941 | 0.8185 | 0.8061 | |
 Appendix H Memory and Parameter Comparison
@@
 | Our | 19.02 | 1081.99 | 94.71 | 19.02 | 1082.68 | 13.68 | |
-Appendix I Limitations
-In this section, we provide limitations on the proposed benchmark.
-A limitation of the proposed benchmark is that it considers a limited number of datasets and solely on random splitting for dataset splitting.
-Additionally, the benchmark only evaluates models using simple inputs without incorporating additional information, which may not reflect the current trend where many models depend on pre-trained language models.